### v1

# Proving the existence of localized patterns and saddle node bifurcations in 1D activator-inhibitor type models

Abstract
In this paper, we present a general framework for constructively proving the existence and stability of stationary localized 1D solutions and saddle-node bifurcations in activator–inhibitor systems using computer-assisted proofs. Specifically, we develop the necessary analysis to compute explicit upper bounds required in a Newton–Kantorovich approach. Given an approximate solution $\bar{\mathbf{u}}$, this approach relies on establishing that a well-chosen fixed point map is contracting on a neighborhood $\bar{\mathbf{u}}$. For this matter, we construct an approximate inverse of the linearization around $\bar{\mathbf{u}}$, and establish sufficient conditions under which the contraction is achieved. This provides a framework for which computer-assisted analysis can be applied to verify the existence and local uniqueness of solutions in a vicinity of $\bar{\mathbf{u}}$, and control the linearization around $\bar{\mathbf{u}}$. Furthermore, we extend the method to rigorously establish saddle-node bifurcations of localized solutions for the same type of models, by considering a well–chosen zero–finding problem. This depends on the rigorous control of the spectrum of the linearization around the bifurcation point. Finally, we demonstrate the effectiveness of the framework by proving the existence and stability of multiple steady-state patterns in various activator–inhibitor systems, as well as a saddle–node bifurcation in the Glycolysis model.
Key words. Localized stationary patterns, Activator-Inhibitor models, saddle node bifurcations, stability of solutions, Computer-Assisted Proofs
1 Introduction
In this paper, we constructively prove the existence of stationary localized solutions and saddle-node bifurcations (for branches of localized solutions) in the following class of Partial Differential Equations (PDEs):
$\begin{split}\partial_{t}u&=\delta^{2}\Delta u-cu+dv+u^{2}v+a,\\ \partial_{t}v&=\Delta v+hu-jv-u^{2}v+b\\ \end{split}\quad\penalty 10000\ \penalty 10000\ (u,v)=\left(u(x,t),v(x,t)\right),\penalty 10000\ \penalty 10000\ x\in\mathbb{R},$ (1)
where $\delta$, $c$, $d$, $h$, $j$, $a$, and $b$ are positive parameters, and $\Delta$ is the Laplacian operator. These equations arise naturally in various fields such as biology and chemistry, where they model a chemical reaction between two substances $U$ and $V$. The first species is the activator ($U$) with concentration $u$ and the other one is an inhibitor ($V$) with concentration $v$ GrayScott1985 . This system is set in the regime $0<\delta\ll 1$, where the inhibitor diffuses more than the activator. This class of PDEs encompasses many reaction-diffusion systems that have been studied extensively in the literature, such as the Schnakenberg model schankenberg_og ($c=1,d=h=j=0$), the Sel’kov Schnakenberg model selkov_schankenberg_og ($c=1,d=j,h=0$), the Brusselator model brusselator_og (h=c1>0,d=j,h=0h=c-1>0,d=j,h=0), the Root hair model root_hair_og ($a=0,d=j,h>c>0$), and the Glycolysis model glycolysis_og ($c=1,d=j,a=h=0$).
1.1 Stationary localized solutions
The study of pattern formation originates in the seminal work of Alan Turing turing_morphogenesis . In this paper, Turing demonstrated how patterns can emerge from perturbations of a steady-state solution through what is now referred to as a Turing instability, or diffusion-driven instability. In one spatial dimension, this mechanism typically manifests as a supercritical pitchfork bifurcation from the steady-state solution Meron2015 .
Reaction–Diffusion (RD) systems have attracted significant interest due to their physical relevance, with applications spanning biology root_hair_og ; selkov_schankenberg_og , chemistry GrayScott1985 ; schankenberg_og ; brusselator_og , and ecology SITEUR201481 ; ZELNIK201727 . These systems are particularly compelling because they often exhibit rich and complex behaviors while maintaining a relatively simple mathematical structure. Their nonlinearities are typically polynomial, as in (1), which includes a simple auto-catalytic term of the form $\pm u^{2}v$ in each equation. RD systems have been studied extensively using a variety of analytical approaches. For example, geometric perturbation theory and spatial dynamics have been employed to establish the existence of localized and periodic solutions on long one-dimensional domains in the Gray–Scott model Doelman1997 , later extended to higher dimensions MuratovOsipov2000 and to more general RD systems doelman_geometric . The stability of localized patterns can be investigated using asymptotic expansions, wherein the solution is analyzed through an inner expansion near the pattern and a far-field expansion in the tail. Matching these expansions provides approximations of the eigenvalues of the linearized PDE operator, yielding stability criteria. This method has been successfully applied in the one-dimensional Gierer–Meinhardt model iron_semistrong , the Schnakenberg model Iron2004 , and the Gray–Scott model muratov_semistrong . Another central phenomenon in RD systems is the connection between localized pattern formation and homoclinic snaking arising from a Hamiltonian–Hopf bifurcation, which has been investigated in knobloch_dissipative ; VERSCHUEREN2021132858 .
In recent work, Champneys et al.Champneys_Bistability outlined several mechanisms for pattern formation in RD systems. They investigated the bifurcation structure of the Brusselator model, constructing approximate solutions via matched asymptotic expansions and numerically continuing them to reveal homoclinic snaking. A similar analysis was carried out for the Schnackenberg model in alsaadi_spikes , where the same approach was employed to build approximate solutions and to numerically observe the system’s snaking behavior, alongside a stability analysis of spike solutions. It is important to note that the validity of matched asymptotic expansions is restricted to specific parameter regimes: for the Brusselator, this requires $\delta\ll 1$, while for the Schnackenberg model the appropriate scaling is $a=\epsilon\alpha$, $b=\epsilon^{2}\beta$, with $\epsilon\ll 1$. Building on these results, AlSaadi_unifying_framework extended the analysis to demonstrate the ubiquity of such behavior in systems of the form (1). Using weakly nonlinear analysis, they characterized the bifurcation structure of (1), identifying parameter regions where homoclinic snaking is expected to occur for both large and small parameter values. They then applied matched asymptotic expansions to numerically explore the snaking region and further examined the stability of solutions within this region as the diffusion coefficient $\delta$ increases. It should be emphasized, however, that while their theoretical analysis is posed on $\mathbb{R}$, the corresponding numerical results were obtained on long one-dimensional domains.
In their conclusion, the authors of AlSaadi_unifying_framework raise the critical open question of providing a rigorous proof for their results. Addressing this challenge is essential, as it bridges the gap between formal asymptotics, numerical exploration, and mathematical rigor. In the next subsection, we present a computer-assisted approach that delivers such a rigorous justification, thereby validating and strengthening their numerical observations. Constructive existence proofs of patterns, similar to Figure 1, as well as a constructive proof of a saddle node will be provided. In particular, our method is based on the computer-assisted analysis presented in gs_cadiot_blanco . As these models arise from physical context, the question of the stability of these solutions is also studied by rigorously controlling the spectrum of the jacobian of the linearization around an approximate solution using the approach of cadiot2025stabilityanalysislocalizedsolutions
(a) Approximation of a “multi-pulse” solution in the brusselator equation
(b) Approximation of a saddle node bifurcation “pulse” in the Glycolysis equation
Figure 1: Approximation of a steady-state and saddle node bifurcation of the activator-inhibitor models for different parameter regimes.
1.2 A Computer-assisted approach
One way to approach solving semilinear PDEs on $\mathbb{R}^{m}$ is to use computer-assisted proofs (CAPs) techniques. CAPs can help derive a constructive methodology that relies on a Newton-Kantorovich argument. It involves the computation of specific bounds, which can be evaluated rigorously on the computer by the use of interval arithmetic for instance julia_interval . For the ODE case, one might use the parameterization method MR1976079 ; MR1976080 ; MR2177465 . This method relies on the rigorous parametrization of invariant manifolds at the zero solution. Then, this allows one to formulate a projected boundary value problem which one can treat using Chebyshev series or splines. Still for ODEs, the analysis derived in breden2025solutionsdifferentialequationsfreudweighted ; breden2024constructiveproofssemilinearpdes allows to produce constructive existence proofs in a well-chosen class of equations. Indeed, such equations possess a confining potential, allowing the resolvent of the differential operator to be compact on a well-chosen space of functions. Note that the method derived in breden2024constructiveproofssemilinearpdes also applies to PDEs possessing a linear part of the form $\Delta u+\frac{x}{2}\cdot\nabla u$. Moving to PDEs, but in the weak formulation, the authors of plum_numerical_verif presented a method for proving weak solutions to second and fourth-order PDEs. Their approach relies on the rigorous control of the spectrum of the linearization around an approximate solution. Then, they use a homotopy argument and the Temple-Lehmann-Goerisch method (see Section 10.2 in plum_numerical_verif ). This approach was then applied by Wunderlich in plum_thesis_navierstokes . Wunderlich proved the existence of a weak solution to the Navier-Stokes equations defined on an infinite strip with an obstacle. As far as strong solutions are concerned, olivier_radial provides a methodology for proving the existence of radially symmetric solutions. Using the radially symmetric ansatz, the PDE is transformed into an ordinary differential equation. The approach then relies on a rigorous enclosure of the center stable manifold by using a Lyapunov-Perron operator. This allows one to solve a boundary value problem on $(0,\infty)$. Removing the radially symmetric assumption, the authors of unbounded_domain_cadiot provide a general method for proving the existence and local uniqueness of localized solutions to semilinear autonomous PDEs on $\mathbb{R}^{m}$. The approach is based on Fourier series and describes the necessary tools to construct an approximate solution, $u_{0}$, and an approximate inverse $\mathbb{A}$ of the linearization $D\mathbb{F}(u_{0})$ about $u_{0}$. An application to the 1D Kawahara equation was provided in the same paper. In sh_cadiot , the method of unbounded_domain_cadiot was applied to the planar Swift Hohenberg PDE. The method was then extended further by the authors of gs_cadiot_blanco . In unbounded_domain_cadiot , the method was presented for scalar PDEs. In gs_cadiot_blanco , the authors generalized the approach to systems of PDEs. This allows for rigorous proofs of localized solutions to systems of PDEs. The approach was demonstrated on the 2D Gray-Scott system of equations. Then, in symmetry_blanco_cadiot , the authors extended further this methodology to treat existence proofs of symmetric solutions, where the symmetry might be given by a general group of symmetries. The 2D planar Swift Hohenberg equation was again used as an illustration, allowing for the authors to obtain existence proofs of solutions with various dihedral symmetries.
In this paper, we base our analysis on the method developed in gs_cadiot_blanco . Our goal is to provide rigorous validations to some of the observations made by the authors of AlSaadi_unifying_framework . In fact, we aim at applying similar ideas as in gs_cadiot_blanco to a well-chosen augmented system in order to prove the existence of saddle node bifurcations. This provides a novel computer-assisted approach for proving the existence of saddle-node bifurcations for localized solutions.
Going back to the study of (1), we assume in the rest of the paper that $j=d$ and that $c>h$. Then, we use the following change of variables and re-normalisations
$u_{1}(x,t)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}u\left(x,t\right)-\lambda_{1}\text{ and }u_{2}(x,t)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}v(x,t)-\lambda_{2},\text{ where }$
$\lambda_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{a+b}{c-h}\text{ and }\lambda_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{c(c-h)(a+b)-a(c-h)^{2}}{(a+b)^{2}+d(c-h)^{2}}.$
Note that $(\lambda_{1},\lambda_{2})$ is a steady state of (1). Consequently, using the above definitions, we can look for $(u_{1},u_{2})(x,t)\to 0$ as $|x|\to\infty$. In fact, we obtain that $(u_{1},u_{2})$ equivalently solve
$\begin{split}\partial_{t}u_{1}&=\nu\Delta u_{1}+\nu_{1}u_{1}+\nu_{2}u_{2}+u_{1}^{2}u_{2}+\nu_{4}u_{1}^{2}+\nu_{5}u_{1}u_{2},\\ \partial_{t}u_{2}&=\Delta u_{2}-\nu_{2}u_{2}+\nu_{3}u_{1}-u_{1}^{2}u_{2}-\nu_{4}u_{1}^{2}-\nu_{5}u_{1}u_{2}\end{split}$ (2)
where
$\nu\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\delta^{2},\penalty 10000\ \nu_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\lambda_{1}\lambda_{2}-c,\penalty 10000\ \nu_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}d+\lambda_{1}^{2},\penalty 10000\ \nu_{3}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}h-2\lambda_{1}\lambda_{2},\penalty 10000\ \nu_{4}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\lambda_{2},\penalty 10000\ \nu_{5}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\lambda_{1}.$ (3)
Using the above, $(u,v)$ solves (1) if and only if $(u_{1},u_{2})$ solves (2). Let us denote $\mathbf{u}=(u_{1},u_{2})$ and define $\mathbb{f}$ as follows
$\mathbb{f}(\mathbf{u})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{l}\mathbf{u}+\mathbb{g}(\mathbf{u}),$ (4)
where
$\mathbb{l}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}\nu\Delta+\nu_{1}I&\nu_{2}I\\ \nu_{3}I&\Delta-\nu_{2}\end{bmatrix}=\begin{bmatrix}\mathbb{l}_{11}&\mathbb{l}_{12}\\ \mathbb{l}_{21}&\mathbb{l}_{22}\end{bmatrix},\penalty 10000\ \penalty 10000\ \mathbb{g}(\mathbf{u})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}u_{1}^{2}u_{2}+\nu_{4}u_{1}^{2}+\nu_{5}u_{1}u_{2}\\ -u_{1}^{2}u_{2}-\nu_{4}u_{1}^{2}-\nu_{5}u_{1}u_{2}\end{bmatrix}$. (5)
In fact, (2) has the condensed parabolic form
$\partial_{t}\mathbf{u}=\mathbb{f}(\mathbf{u}).$ (6)
In this paper, we will focus our attention on stationary localized solutions, that is $\mathbf{u}:\mathbb{R}\to\mathbb{R}^{2}$ such that $\mathbb{f}(\mathbf{u})=0$ and satisfying $\mathbf{u}(x)\to 0$ as $|x|\to\infty.$ Given a well-chosen Hilbert space $\mathcal{H}_{e}$ (cf. Section 2), containing even localized functions, we look for zeros of $\mathbb{f}$ in $\mathcal{H}_{e}$. In fact, we demonstrate that $\mathbb{f}:\mathcal{H}_{e}\to L^{2}_{e}$ is a smooth operator, where $L^{2}_{e}$ is the restriction of even functions of $L^{2}(\mathbb{R})\times L^{2}(\mathbb{R})$. Then, following the analysis derived in unbounded_domain_cadiot ; gs_cadiot_blanco , we prove the existence of zeros of $\mathbb{f}$ using a Newton-Kantorovich approach (cf. Section 3). In fact, our analysis relies on the construction of an approximate solution $\bar{\mathbf{u}}\in\mathcal{H}_{e}$ and of an approximate inverse $\mathbb{A}:{L}^{2}_{e}\to\mathcal{H}_{e}$ for the Fréchet derivative $D\mathbb{f}(\bar{\mathbf{u}})$. Indeed, we can then define a Newton-like fixed point operator $\mathbb{T}$ given as
$\mathbb{T}(\mathbf{u})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbf{u}-\mathbb{A}\mathbb{f}(\mathbf{u})$
and we prove that there exists $r>0$ such that $\mathbb{T}$ is contracting from $\overline{B_{r}(\bar{\mathbf{u}})}$ to itself, where $\overline{B_{r}(\bar{\mathbf{u}})}$ is the closed ball in $\mathcal{H}_{e}$ of radius $r$ centered at $\bar{\mathbf{u}}$. Using the Banach-fixed point theorem, this implies the existence of a unique zero $\tilde{\mathbf{u}}$ of $\mathbb{f}$ in $\overline{B_{r}(\bar{\mathbf{u}})}$. The verification of the contractivity of $\mathbb{T}$ on $\overline{B_{r}(\bar{\mathbf{u}})}$ involves the rigorous computation of different quantities, which we expose in Section 4. This allows us to provide novel constructive existence proofs of localized solutions in (2).
[…]
$\mathbb{f}(\tilde{\nu},\tilde{\mathbf{u}})=0,\penalty 10000\ D_{u}\mathbb{f}(\tilde{\nu},\tilde{\mathbf{u}})\mathbf{w}=0$
and some extra non-degeneracy conditions (given in jp_saddle_node for instance). In particular, jp_saddle_node provides an equivalent definition of a saddle-node bifurcation, thanks to isolated zeros of a well-chosen map. In this paper, we consider such a definition, which we explicit below. We introduce the Hilbert spaces $H_{1}=\mathbb{R}\times\mathcal{H}_{e}\times\mathcal{H}_{e}$ and $H_{2}=\mathbb{R}\times L_{e}^{2}\times L^{2}_{2}$ and we define the so-called Saddle-Node Map $\mathbb{F}:H_{1}\to H_{2}$ as follows
$\mathbb{F}(\mathbf{x})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}\phi(\mathbf{w})-\alpha\\ \mathbb{f}(\nu,\mathbf{u})\\ D_{\mathbf{u}}\mathbb{f}(\nu,\mathbf{u})\mathbf{w}\end{bmatrix},\text{ with }\mathbf{x}=(\nu,\mathbf{u},\mathbf{w})\in H_{1},$ (7)
and where $\alpha\in\mathbb{R}$, $\phi:\mathcal{H}_{e}\to\mathbb{R}$ is a non-trivial linear functional. Then, using jp_saddle_node , we have the following definition for a saddle-node bifurcation.
Definition 1.1.
[…]
Remark 1.1.
[…]
This manuscript is organized as follows. Section 2 introduces the notation, assumptions, and restrictions used throughout the paper. In Section 3, we introduce a Newton-Kantorovich approach associated to the construction of an approximate solution $\bar{\mathbf{u}}$ and of an approximate inverse $\mathbb{A}$ for $D\mathbb{f}(\bar{\mathbf{u}})$. The specific construction of our approximate objects allow us to derive explicit formulas in Section 4 for the application of the Newton-Kantorovich approach. In the same section, we present applications of our analysis and derive multiple constructive existence proofs of localized solutions. In Section 5, we conduct a similar analysis for the constructive proofs of solutions to the augmented problem (7). Finally, we derive in Section 6 a computer-assisted approach for controlling the spectrum of the Jacobian at a given localized solution. This allows us to establish the existence of a saddle-node in the Glycolysis model, as well as stability of some localized solutions.
2 Setup of the Problem
Recall that we wish to prove the existence of localized solutions to (2) for the spatial domain $\mathbb{R}$. For this matter, we introduce usual notations of functional analysis on $\mathbb{R}$. We first define the Lebesgue notation on a product space as $L^{2}=L^{2}(\mathbb{R})\times L^{2}(\mathbb{R})$ and $L^{2}({\Omega_{0}})$ on a bounded domain $\Omega_{0}$} in $\mathbb{R}$. More generally, $L^{p}$ denotes the usual $p$ Lebesgue product space with two components on $\mathbb{R}$ associated to its norm $\|\cdot\|_{p}$. Moreover, given $s\in\mathbb{R}$, denote by $H^{s}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}H^{s}(\mathbb{R})\times H^{s}(\mathbb{R})$ the usual Sobolev space on $\mathbb{R}$. For a bounded linear operator $\mathbb{K}:L^{2}\to L^{2}$, denote by $\mathbb{K}^{*}$ the adjoint of $\mathbb{K}$ in $L^{2}$. Moreover, if $\mathbf{u}\in L^{2}$ where $\mathbf{u}=(u_{1},u_{2})$, denote by $\hat{\mathbf{u}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{F}(\mathbf{u})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(\mathcal{F}(u_{1}),\mathcal{F}(u_{2}))$ the Fourier transform of $\mathbf{u}$. More specifically, $\hat{u_{j}}(\xi)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\int_{\mathbb{R}}u_{j}(x)e^{-i2\pi x\cdot\xi}dx$ for all $\xi\in\mathbb{R}$ and $j\in\{1,2\}$.
Adopting the notation of unbounded_domain_cadiot , we look for $\mathbf{u}=(u_{1},u_{2})$ such that
$\mathbb{f}(\mathbf{u})=0$ (8)
where $\mathbb{f}$ is given in (5). Note that $\mathbb{l}$ possesses a symbol (Fourier transform) $\mathscr{l}:\mathbb{R}\to\mathbb{R}\times\mathbb{R}$ given as
$\mathscr{l}(\xi)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}\mathscr{l}_{11}(\xi)&\mathscr{l}_{12}(\xi)\\ \mathscr{l}_{21}(\xi)&\mathscr{l}_{22}(\xi)\end{bmatrix}=\begin{bmatrix}-\nu|2\pi\xi|^{2}+\nu_{1}&\nu_{2}\\ \nu_{3}&-|2\pi\xi|^{2}-\nu_{2}\end{bmatrix}$ (9)
for all $\xi\in\mathbb{R}$. We now recall a necessary assumption from gs_cadiot_blanco that our framework requires. We will later verify it for a class of Activator-Inhibitor models (2).
Assumption 1.
Given $\mathscr{l}$ as in (9), assume there exists $\sigma_{0}>0$ such that
$|\det(\mathscr{l}(\xi))|\geq\sigma_{0}\text{ for all }\xi\in\mathbb{R}.$ (10)
That, is, $\det(\mathscr{l}(\xi))$ is bounded away uniformly from 0.
We now provide the values of the parameters $\nu_{j}$ ($j\in\{1,\dots,5\})$ such that $\mathscr{l}$ satisfies Assumption 1.
Lemma 2.1.
$\mathscr{l}$ is invertible if
$(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})<0,$ (11)
or if
$\begin{cases}(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})\geq 0\penalty 10000\ \penalty 10000\ \text{ and }\\ \nu\nu_{2}-\nu_{1}>\sqrt{(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})}\end{cases}$ (12)
Proof.
$\mathscr{l}$ is invertible if and only if $\mathscr{l}_{11}(\xi)\mathscr{l}_{22}(\xi)-\mathscr{l}_{12}(\xi)\mathscr{l}_{21}(\xi)\neq 0,\forall\xi\in\mathbb{R}$. This is equivalent to studying the roots of the second order polynomial $x\mapsto\nu x^{2}+(\nu\nu_{2}-\nu_{1})x-\nu_{2}(\nu_{1}+\nu_{3})$. The proof is then obtained using basic properties of second order polynomials. ∎
From now on, we assume that the parameters $\nu$ and $\nu_{j}$ ($j\in\{1,\dots,5\})$ satisfy the conditions of Lemma 2.1, yielding that $\mathscr{l}$ is invertible. This allows to define the following norm and inner product
$\|\mathbf{u}\|_{\mathcal{H}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\|\mathbb{l}\mathbf{u}\|_{2}\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ (\mathbf{u},\mathbf{v})_{\mathcal{H}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(\mathbb{l}\mathbf{u},\mathbb{l}\mathbf{v})_{2}$ (13)
for all $\mathbf{u},\mathbf{v}\in\mathcal{H}$, where $\mathcal{H}$ is the Hilbert space
$\mathcal{H}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{\mathbf{u}\in L^{2},\|\mathbf{u}\|_{\mathcal{H}}<\infty\}.$ (14)
Using Plancherel’s identity, we have
$(\mathbf{u},\mathbf{v})_{\mathcal{H}}=(\mathscr{l}\hat{\mathbf{u}},\mathscr{l}\hat{\mathbf{v}})_{2}$
for all $\mathbf{u},\mathbf{v}\in\mathcal{H}.$ In particular, note that $\mathbb{l}:\mathcal{H}\to L^{2}$ is a well-defined bounded linear operator, which is actually an isometric isomorphism. We now want to prove that the operator $\mathbb{g}:\mathcal{H}\to L^{2}$ is a smooth operator. For this matter, we recall two spaces from gs_cadiot_blanco : $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ given as
$\mathcal{M}_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{M=\left(M_{i,j}\right)_{i,j\in\{1,2\}},\text{ where }M_{i,j}\in L^{\infty}(\mathbb{R})\text{ and }\penalty 10000\ \|M\|_{\mathcal{M}_{1}}<\infty\right\}$ (15)
$\mathcal{M}_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{M=\left(M_{i,j}\right)_{i,j\in\{1,2\}},\text{ where }M_{i,j}\in L^{2}(\mathbb{R})\text{ and }\penalty 10000\ \|M\|_{\mathcal{M}_{2}}<\infty\right\}$ (16)
with their associated norms
$\|M\|_{\mathcal{M}_{1}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sup_{\xi\in\mathbb{R}}\sup\limits_{\begin{subarray}{c}x\in\mathbb{R}^{2}\\ |x|_{2}=1\end{subarray}}|M(\xi)x|_{2}$ (17) $\vskip 4.26773pt\|M\|_{\mathcal{M}_{2}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\max_{i\in\{1,2\}}\left\{\left(\sum_{j=1}^{2}\|M_{i,j}\|_{L^{2}(\mathbb{R})}^{2}\right)^{\frac{1}{2}}\right\}$ (18)
In fact, using $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, gs_cadiot_blanco provides sufficient conditions for which products on $\mathcal{H}\times\mathcal{H}\to L^{2}$ are well-defined. Using this result, we obtain the following lemma.
Lemma 2.2.
Let $\mathbf{u}=(u_{1},u_{2}),\mathbf{v}=(v_{1},v_{2}),$ and $\mathbf{w}=(w_{1},w_{2})$. Suppose Assumption 1 is verified and let $\kappa\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\|\mathscr{l}^{-1}\|_{\mathcal{M}_{1}}\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}$. Then
$\|u_{i}v_{j}\|_{2}\leq\kappa\|\mathbf{u}\|_{\mathcal{H}}\|\mathbf{v}\|_{\mathcal{H}}\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ \|u_{i}v_{j}w_{k}\|_{2}\leq\kappa\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}\|\mathbf{u}\|_{\mathcal{H}}\|\mathbf{v}\|_{\mathcal{H}}\|\mathbf{w}\|_{\mathcal{H}}$ (19)
for any $i,j,k\in\{1,\dots,2\}$.
Proof.
To begin, observe that
$\|u_{i}v_{j}\|_{2}$ $\leq\|u_{i}\|_{\infty}\|v_{j}\|_{2}\leq\|\mathbf{u}\|_{\infty}\|\mathbf{v}\|_{2}\leq\|\hat{\mathbf{u}}\|_{1}\|\hat{\mathbf{v}}\|_{2}$
where we used Plancherel’s theorem. But now, using the proof of Lemma 2.1 in gs_cadiot_blanco , we get
$\|\hat{\mathbf{v}}\|_{2}$ $=\|\mathscr{l}^{-1}\mathscr{l}\hat{\mathbf{v}}\|_{2}\leq\|\mathscr{l}^{-1}\|_{\mathcal{M}_{1}}\|\mathscr{l}\hat{\mathbf{v}}\|_{2}=\|\mathscr{l}^{-1}\|_{\mathcal{M}_{1}}\|\mathbf{v}\|_{\mathcal{H}}.$ (20)
Finally, let us analyze $\|\hat{\mathbf{u}}\|_{1}$. Using the proof of Lemma 2.1 in gs_cadiot_blanco again, we get
$\|\hat{\mathbf{u}}\|_{1}$ $=\|\mathscr{l}^{-1}\mathscr{l}\hat{\mathbf{u}}\|_{1}\leq\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}\|\mathscr{l}\hat{\mathbf{u}}\|_{2}\leq\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}\|\mathbf{u}\|_{\mathcal{H}}$
where we again used Plancheral’s theorem on the final step. All together, we have
$\|u_{i}v_{j}\|_{2}$ $\leq\|\mathscr{l}^{-1}\|_{\mathcal{M}_{1}}\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}\|\mathbf{u}\|_{\mathcal{H}}\|\mathbf{v}\|_{\mathcal{H}}.$
The case $\|u_{i}v_{j}w_{k}\|_{2}$ is treated similarly observing that
$\|u_{i}v_{j}w_{k}\|_{2}\leq\|u_{i}\|_{\infty}\|w_{k}\|_{\infty}\|v_{j}\|_{2}.\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \qed$
Using Lemma 2.2, we obtain that $\mathbb{g}:\mathcal{H}\to L^{2}$ is a smooth operator. In particular, this yields that $\mathbb{f}:\mathcal{H}\to L^{2}$ is smooth as well. In fact, Proposition 2.1 in gs_cadiot_blanco yields that solutions to (2) in $\mathcal{H}$ are actually infinitely differentiable, meaning that our analysis on $\mathcal{H}$ will yield strong solutions to (2).
Supposing that $\mathbf{u}=(u_{1},u_{2})$ is a solution of (2), then any translation in space provides a new solution. Therefore, in order to isolate a localized solution in the set of solutions, we choose to look for even solutions, that is we enforce that $\mathbf{u}(x)=\mathbf{u}(-x)$ for all $x\in\mathbb{R}$. With this in mind, we introduce the following even restriction $\mathcal{H}_{e}\subset\mathcal{H}$
$\mathcal{H}_{e}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{\mathbf{u}\in\mathcal{H}\ |\ \mathbf{u}(x)=\mathbf{u}(-x)\}.$ (21)
Similarly, denote $L_{e}^{2}$ the Hilbert subspace of $L^{2}$ satisfying the even symmetry. In particular we notice that if $\mathbf{u}\in\mathcal{H}_{e}$, then $\mathbb{l}\mathbf{u}\in L^{2}_{e}$ and $\mathbb{g}(\mathbf{u})\in L^{2}_{e}$ ; hence, $\mathbb{l}$ and $\mathbb{g}$ are well-defined as operators from $\mathcal{H}_{e}$ to $L^{2}_{e}$. Finally, we look for solutions of the following problem
$\mathbb{f}(\mathbf{u})=0\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ \mathbf{u}\in\mathcal{H}_{e}.$ (22)
2.1 Periodic Spaces
In this section, we recall some notations introduced in Section 2.4 of unbounded_domain_cadiot . In particular, the objects defined in the above sections have a corresponding representation in Fourier series when restricted to a bounded domain. Specifically, we define our bounded domain of interest $\Omega_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(-d,d)$ where $1\leq d<\infty$. Moreover, denote $\tilde{n}=\frac{n}{2d}\in\mathbb{R}$ for all $n\in\mathbb{Z}$. As for the continuous case, we want to restrict to Fourier series representing even functions. Let $\mathbf{u}_{n}=((u_{1})_{n},(u_{2})_{n})$ be the $n$ Fourier coefficient of $\mathbf{u}$. In terms of Fourier coefficients, the even restriction reads
$\mathbf{u}_{n}=\mathbf{u}_{-n}\text{ for all }n\in\mathbb{Z}.$
In particular, when enforcing the even symmetry, we can restrict the indices of Fourier coefficients from $\mathbb{Z}$ to the reduced set $\mathbb{N}_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{n\in\mathbb{Z}\ |\ 0\leq n\right\}.$ Now, let $(\alpha_{n})_{n\in\mathbb{N}_{0}}$ be defined by
$\alpha_{n}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{cases}1&\text{ if }n=0\\ 2&\text{ if }n>0.\end{cases}$ (23)
In particular, $\alpha_{n}$ is the size of $\mathrm{orb}_{\mathbb{Z}_{2}}(n)$. Next, let $\ell^{p}(\mathbb{N}_{0})$ denote the following Banach space
$\ell^{p}(\mathbb{N}_{0})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{U=(u_{n})_{n\in\mathbb{N}_{0}}:\penalty 10000\ \|U\|_{p}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\sum_{n\in\mathbb{N}_{0}}\alpha_{n}|u_{n}|^{p}\right)^{\frac{1}{p}}<\infty\right\}.$ (24)
In the above, note that $U$ is a sequence of scalars, and not a sequence of vectors. In particular, $\ell^{2}(\mathbb{N}_{0})$ is an Hilbert space associated to its natural inner product $(\cdot,\cdot)_{2}$ given by
$(U,V)_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sum_{n\in\mathbb{N}_{0}}\alpha_{n}u_{n}v_{n}^{*}$
for all $U=(u_{n})_{n\in\mathbb{N}_{0}},V=(v_{n})_{n\in\mathbb{N}_{0}}\in\ell^{2}(\mathbb{N}_{0})$ and * denotes complex conjugation. Then, we define the Banach space $\ell^{p}_{e}$ as
$\ell^{p}_{e}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\ell^{p}(\mathbb{N}_{0})\times\ell^{p}(\mathbb{N}_{0}),\penalty 10000\ \text{ with norm }\penalty 10000\ \|\mathbf{U}\|_{p}=(\|U_{1}\|_{p}^{p}+\|U_{2}\|_{p}^{p})^{\frac{1}{p}}.$ (25)
For a bounded operator $K:\ell^{2}_{e}\to\ell^{2}_{e}$ (resp. $K:\ell^{2}(\mathbb{N}_{0})\to\ell^{2}(\mathbb{N}_{0})$), $K^{*}$ denotes the adjoint of $K$ in $\ell^{2}_{e}$ (resp. $\ell^{2}(\mathbb{N}_{0})$). Now, similarly as what is achieved gs_cadiot_blanco , we define $\gamma\penalty 10000\ :\penalty 10000\ L^{2}_{e}(\mathbb{R})\to\ell^{2}(\mathbb{N}_{0})$ as
$\left(\gamma(u)\right)_{n}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{|{\Omega_{0}}|}\int_{\Omega_{0}}u(x)e^{-2\pi i\tilde{n}\cdot x}dx$ (26)
for all $n\in\mathbb{N}_{0}$ and all $u\in L^{2}_{e}(\mathbb{R})$. Similarly, we define $\gamma^{\dagger}:\ell^{2}(\mathbb{N}_{0})\to L^{2}_{e}(\mathbb{R})$ as
$\gamma^{\dagger}\left(U\right)(x)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{1}_{\Omega_{0}}(x)\sum_{n\in\mathbb{N}_{0}}\alpha_{n}u_{n}\cos(2\pi\tilde{m}x)$ (27)
for all $x\in\mathbb{R}$ and all $U=\left(u_{n}\right)_{n\in\mathbb{N}_{0}}\in\ell^{2}(\mathbb{N}_{0})$, where $\mathbb{1}_{\Omega_{0}}$ is the characteristic function on $\Omega_{0}$}. We now introduce a similar notation for functions in the product space $L^{2}_{e}$. In particular, define $\text{{$\gamma$}}:L^{2}_{e}\to\ell^{2}_{e}$ as
$\text{{$\gamma$}}(\mathbf{u})=(\gamma(u_{1}),\gamma(u_{2}))$ (28)
and $\text{{$\gamma$}}^{\dagger}:\ell^{2}_{e}\to L^{2}_{e}$ as
$\text{{$\gamma$}}^{\dagger}(\mathbf{U})=(\gamma^{\dagger}(U_{1}),\gamma^{\dagger}(U_{2})).$ (29)
More specifically, given $\mathbf{u}\in L^{2}_{e}$, $\text{{$\gamma$}}(\mathbf{u})$ represents the Fourier coefficients indexed on $\mathbb{N}_{0}$ of the restriction of $\mathbf{u}$ on $\Omega_{0}$}. Conversely, given a sequence $\mathbf{U}\in\ell^{2}_{e}$, $\text{{$\gamma$}}^{\dagger}\left(\mathbf{U}\right)$ is the function representation of $\mathbf{U}$ in $L^{2}_{e}.$ In particular, notice that $\text{{$\gamma$}}^{\dagger}\left(\mathbf{U}\right)(x)=0$ for all $x\notin{\Omega_{0}}.$
Then, recall similar notations from unbounded_domain_cadiot
$L^{2}_{e,{\Omega_{0}}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{\mathbf{u}\in L^{2}_{e}:\text{supp}(\mathbf{u})\subset\overline{{\Omega_{0}}}\right\}\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ \mathcal{H}_{e,{\Omega_{0}}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{\mathbf{u}\in\mathcal{H}_{e}:\text{supp}(\mathbf{u})\subset\overline{{\Omega_{0}}}\right\}.$ (30)
Moreover, define $\mathcal{B}(L^{2}_{e})$ (respectively $\mathcal{B}(\ell^{2}_{e})$) as the space of bounded linear operators on $L^{2}_{e}$ (respectively $\ell^{2}_{e}$) and denote by $\mathcal{B}_{\Omega_{0}}(L^{2}_{e})$ the following subspace of $\mathcal{B}(L^{2}_{e})$
$\mathcal{B}_{\Omega_{0}}(L^{2}_{e})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{\mathbb{K}\in\mathcal{B}(L^{2}_{e}):\mathbb{K}=\mathbb{1}_{\Omega_{0}}\mathbb{K}\mathbb{1}_{\Omega_{0}}\}.$ (31)
Finally, define $\text{{$\Gamma$}}:\mathcal{B}(L^{2}_{e})\to\mathcal{B}(\ell^{2}_{e})$ and $\text{{$\Gamma$}}^{\dagger}:\mathcal{B}(\ell^{2}_{e})\to\mathcal{B}(L^{2}_{e})$ as follows
[…]
for all $\mathbb{K}\in\mathcal{B}(L^{2}_{e})$ and all $K\in\mathcal{B}(\ell^{2}_{e}).$
The maps defined above in (28), (29) and (32) are fundamental in our analysis as they allow to pass from the problem on $\mathbb{R}$ to the one in $\ell^{2}_{e}$ and vice-versa. Furthermore, the following lemma, provides that this passage is an isometric isomorphism when restricted to the relevant spaces.
Lemma 2.3.
The map $\sqrt{|{\Omega_{0}}|}\text{{$\gamma$}}:L^{2}_{e,{\Omega_{0}}}\to\ell^{2}_{e}$ (respectively $\text{{$\Gamma$}}:\mathcal{B}_{\Omega_{0}}(L^{2}_{e})\to\mathcal{B}(\ell^{2}_{e})$) is an isometric isomorphism whose inverse is given by $\frac{1}{\sqrt{|{\Omega_{0}}|}}\text{{$\gamma$}}^{\dagger}:\ell^{2}_{e}\to L^{2}_{e,{\Omega_{0}}}$ (respectively $\text{{$\Gamma$}}^{\dagger}:\mathcal{B}(\ell^{2}_{e})\to\mathcal{B}_{\Omega_{0}}(L^{2}_{e})$). In particular,
[…]
for all $\mathbf{u}\in L^{2}_{e,{\Omega_{0}}}$ and $\mathbb{K}\in\mathcal{B}_{\Omega_{0}}(L^{2}_{e})$ where $\mathbf{U}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\text{{$\gamma$}}(\mathbf{u})$ and $K\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\text{{$\Gamma$}}(\mathbb{K})$.
Proof.
The proof is obtained following similar steps as the ones of Lemma 3.2 in unbounded_domain_cadiot . ∎
The above lemma not only provides a one-to-one correspondence between the elements in $L^{2}_{e,{\Omega_{0}}}$ (respectively $\mathcal{B}_{\Omega_{0}}(L^{2}_{e})$) and the ones in $\ell^{2}_{e}$ (respectively $\mathcal{B}(\ell^{2}_{e})$) but it also provides an identity on norms. This property is essential in our construction of an approximate inverse in Section 3.3.
Now, we define the Hilbert space $\mathscr{h}$ as
$\mathscr{h}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{\mathbf{U}\in\ell^{2}_{e}\text{ such that }\|\mathbf{U}\|_{\mathscr{h}}<\infty\right\}$
associated to its inner product $(\cdot,\cdot)_{\mathscr{h}}$ and norm $\|\cdot\|_{\mathscr{h}}$ defined as
$(\mathbf{U},\mathbf{V})_{\mathscr{h}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sum_{n\in\mathbb{N}_{0}}\alpha_{n}\left(l(\tilde{n})\mathbf{u}_{n}\right)\cdot\left(l(\tilde{n})\overline{\mathbf{v}_{n}}\right),\penalty 10000\ \|\mathbf{U}\|_{\mathscr{h}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sqrt{(\mathbf{U},\mathbf{U})_{\mathscr{h}}}$
for all $\mathbf{U}=(\mathbf{u}_{n})_{n\in\mathbb{N}_{0}},\mathbf{V}=(\mathbf{v}_{n})_{n\in\mathbb{N}_{0}}\in\mathscr{h}$ and $\alpha_{n}$ is defined as in (23). Denote by $l:\mathscr{h}\to\ell^{2}_{e}$ and $g:\mathscr{h}\to\ell^{2}_{e}$ the Fourier coefficients representation of $\mathbb{l}$ and $\mathbb{g}$ respectively. More specifically,
$l\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}l_{11}&l_{12}\\ l_{21}&l_{22}\end{bmatrix},\penalty 10000\ g(\mathbf{U})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}U_{1}*U_{1}*U_{2}+\nu_{4}U_{1}*U_{1}+\nu_{5}U_{1}*U_{2}\\ -U_{1}*U_{1}*U_{2}-\nu_{4}U_{1}*U_{1}-\nu_{5}U_{1}*U_{2}\end{bmatrix}$
for all $\mathbf{U}\in\mathscr{h}.$ For all $(i,j)\in\{(1,1),(1,2),(2,1),(2,2)\}$, $l_{ij}$ is an infinite diagonal matrix with coefficients $\left(\mathscr{l}_{ij}(\tilde{n})\right)_{n\in\mathbb{N}_{0}}$ on its diagonal. Moreover $U*V\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma(\gamma^{\dagger}(U)\gamma^{\dagger}(V))$ is defined as the discrete convolution (under even symmetry). In particular, notice that Young’s inequality for convolution is applicable
$\|U*V\|_{2}\leq\|U\|_{2}\|V\|_{1}$ (34)
for all $U\in\ell^{2}(\mathbb{N}_{0}),V\in\ell^{1}(\mathbb{N}_{0})$. Furthermore, using the definition of $l$, notice that
$\|\mathbf{U}\|_{\mathscr{h}}=\|l\mathbf{U}\|_{2}.$
Finally, we define $f(\mathbf{U})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}l\mathbf{U}+g(\mathbf{U})$ and introduce
[…]
as the periodic equivalent on $\Omega_{0}$} of (22).
3 Setting up the Computer-Assisted Approach for Localized Patterns
The goal of this section is to set-up a Newton-Kantorovich approach for proving the existence of solutions to (22) thanks to a fixed-point argument. We state the necessary theorem to perform such an analysis, and construct the related objects to apply it.
3.1 Radii-Polynomial Theorem
Given $\overline{\mathbf{u}}\in\mathcal{H}_{e}$, an approximate solution to (22), and $\mathbb{A}:L^{2}_{e}\to\mathcal{H}_{e}$, an approximate inverse to $D\mathbb{f}(\overline{\mathbf{u}})$, we want to prove that there exists $r>0$ such that $\mathbb{T}:\overline{B_{r}(\overline{\mathbf{u}})}\to\overline{B_{r}(\overline{\mathbf{u}})}$ given by
$\mathbb{T}(\mathbf{u})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbf{u}-\mathbb{A}\mathbb{f}(\mathbf{u})$
is well-defined and a contraction. In order to determine a possible value for $r>0$ that would provide the contraction, we wish to use a Radii-Polynomial theorem. In particular, we build $\mathbb{A}:L_{e}^{2}\to\mathcal{H}_{e}$, $\mathcal{Y}_{0},\mathcal{Z}_{1}>0$ and $\mathcal{Z}_{2}:(0,\infty)\to[0,\infty)$ in such a way that the hypotheses of the following theorem are satisfied.
Theorem 3.1.
Let $\mathbb{A}:L_{e}^{2}\to\mathcal{H}_{e}$ be a bounded linear operator. Moreover, let $\mathcal{Y}_{0},\mathcal{Z}_{1}$ be non-negative constants and let $\mathcal{Z}_{2}:(0,\infty)\to[0,\infty)$ be a non-negative function such that for all $r>0$
$\|\mathbb{A}\mathbb{f}(\overline{\mathbf{u}})\|_{\mathcal{H}}\leq$ $\mathcal{Y}_{0}$ (36)
$\|I-\mathbb{A}D\mathbb{f}(\overline{\mathbf{u}})\|_{\mathcal{H}}\leq$ $\mathcal{Z}_{1}$ (37)
$\|\mathbb{A}\left({D}\mathbb{f}(\overline{\mathbf{u}}+\mathbf{h})-D\mathbb{f}(\overline{\mathbf{u}})\right)\|_{\mathcal{H}}\leq$ $\mathcal{Z}_{2}(r)r,\penalty 10000\ \penalty 10000\ \text{for all }\mathbf{h}\in B_{r}(0)$ (38)
If there exists $r_{0}>0$ such that
$\frac{1}{2}\mathcal{Z}_{2}(r_{0})r_{0}^{2}-(1-\mathcal{Z}_{1})r_{0}+\mathcal{Y}_{0}<0,\ and\ \mathcal{Z}_{1}+\mathcal{Z}_{2}(r_{0})r_{0}<1$ (39)
then there exists a unique $\tilde{\mathbf{u}}\in\overline{B_{r_{0}}(\overline{\mathbf{u}})}\subset\mathcal{H}_{e}$ such that $\mathbb{f}(\tilde{\mathbf{u}})=0$, where $B_{r_{0}}(\overline{\mathbf{u}})$ is the open ball of radius $r_{0}$ in $\mathcal{H}_{e}$ and centered at $\overline{\mathbf{u}}$.
Proof.
The proof can be found in gs_cadiot_blanco . ∎
In order to apply Theorem 3.1, we need to construct explicitly $\overline{\mathbf{u}}\in\mathcal{H}_{e}$ and $\mathbb{A}:\mathcal{H}_{e}\to L^{2}_{e}$. These are the topics of the next sections.
3.2 Construction of $\overline{\mathbf{u}}$
In this section, we discuss the construction of $\overline{\mathbf{u}}$, which is an approximate solution to (22). This is generally a challenging problem. In the specific case of (1), there is a rich variety of work on computing numerical solutions to reaction diffusion systems (e.g alsaadi_spikes ; AlSaadi_unifying_framework ; Champneys_Bistability ). Our approach relies on $\overline{\mathbf{u}}$ being constructed numerically on ${\Omega_{0}}=(-d,d)$ thanks to its Fourier coefficients representation. Fix $N\in\mathbb{N}$ to be the size of our numerical approximation for linear operators (i.e. matrices) and $N_{0}\in\mathbb{N}$ to be the one of our Fourier coefficients approximations (i.e. vectors). Now, given $\mathcal{N}\in\mathbb{N}$, let us introduce the following projection operators
[…]
where $I^{\mathcal{N}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{n\in\mathbb{N}_{0},\penalty 10000\ n\leq\mathcal{N}\}$ for all $V=(v_{n})_{n\in\mathbb{N}_{0}}\in\ell^{2}_{e}.$ Then, we define
$(\text{{$\Pi$}}^{\leq\mathcal{N}}\mathbf{U})_{n}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}((\Pi^{\leq\mathcal{N}}U_{1})_{n},(\Pi^{\leq\mathcal{N}}U_{2})_{n})\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ (\text{{$\Pi$}}^{>\mathcal{N}}\mathbf{U})_{n}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}((\Pi^{>\mathcal{N}}U_{1})_{n},(\Pi^{>\mathcal{N}}U_{2})_{n})$ (40)
for all $\mathbf{U}=\mathbf{u}_{n\in\mathbb{N}_{0}}\in\ell^{2}_{e}$. To obtain a numerical approximation of a pattern, we looked at the numerical solutions found in AlSaadi_unifying_framework ; Champneys_Bistability and tried to recover them as a Fourier series representation using guesses of the form:
$\begin{cases}u_{1,0}(x)=\gamma_{1}\frac{\beta_{1}}{\cosh(\alpha_{1}x)}+\lambda_{1}\\ u_{2,0}(x)=\gamma_{2}\frac{\beta_{2}}{\cosh(\alpha_{2}x)}+\lambda_{2}\end{cases}$
Then we tuned each parameter to replicate the pulse solutions. $\alpha_{i},i=1,2$ controls the stiffness of each spike, and $\beta_{i}$ their amplitude. Once we found a promising candidate, we applied Newton’s method to obtain a better approximate solution of (35). Then we applied the change of variable introduced in Section 1 to get a candidate solution of (2). The multi-pulse solution was obtained through numerical continuation of a candidate in the Brusselator equation, using the AUTO continuation software auto_article . From here, we compute a cosine Fourier sequence approximation of $\mathbf{u}_{0}=(\bar{u}_{1,0},\bar{u}_{2,0})$, $\mathbf{U}_{0}=\gamma(\bar{\mathbf{u}}_{0})\in\ell^{2}_{e}$. We note that $\mathbf{U}_{0}$ satisfies the identity $\mathbf{U}_{0}=\text{{$\Pi$}}^{\leq N_{0}}\mathbf{U}_{0}$ as it is a Fourier sequence with finitely many coefficients.
At this point, we now have a vector representation on $\Omega_{0}$, $\mathbf{u}_{0}=\mathbf{\gamma}^{\dagger}(\mathbf{U}_{0})$ that we extend by 0 to have $\mathbf{u}_{0}\in L^{2}_{e}$; however, such a function is not necessarily in $\mathcal{H}_{e}$. To ensure $\mathbf{u}_{0}\in\mathcal{H}_{e}$, we need to impose that $\mathbf{u}_{0}$ and some of its derivatives to vanish at $\pm d$.
In fact, note that $\mathcal{H}_{e}$ contains the even functions of $H^{2}(\mathbb{R})\times H^{2}(\mathbb{R})$. Consequently, in order to have $\mathbf{u}_{0}\in\mathcal{H}_{e}$, we simply need
$\mathbf{u}_{0}(\pm d)=\partial_{x}\mathbf{u}_{0}(\pm d)=0.$
Note that $\partial_{x}\mathbf{u}_{0}(\pm d)=0$ is automatically satisfied since $\mathbf{u}_{0}(\pm d)$ admits a cosine series representation on $\Omega_{0}$}. Then, using our Fourier representation of $\mathbf{u}_{0}$, we have that for $i=1,2$
$u_{i,0}(d)=u_{i,0}(-d)=\sum_{n=0}^{N_{0}}(U_{i,0})_{n}\cos\left(\frac{n\pi x}{d}\right)=\sum_{n=0}^{N_{0}}(U_{i,0})_{n}(-1)^{n}$
We can ensure that $u_{i,0}(d)=0$ by setting $(U_{i,0})_{0}=\sum_{n=1}^{N_{0}}(U_{i,0})_{n}(-1)^{n+1}$. Hence we can define the operator $\mathcal{P}:\ell^{2}_{e}\rightarrow\ell_{e}^{2}$,
[…]
By construction of $\mathcal{P}$, one can verify that $\overline{\mathbf{u}}(\pm d)=0$, and consequently $\overline{\mathbf{u}}\in\mathcal{H}_{e}$. Moreover, we have that $\overline{\mathbf{U}}=\Pi^{\leq N_{0}}\overline{\mathbf{U}}$ by construction. Overall, we have constructed our approximate solution $\overline{\mathbf{u}}$ as
$\overline{\mathbf{u}}=\gamma^{\dagger}(\overline{\mathbf{U}})\in\mathcal{H}_{e},\penalty 10000\ \text{ and }\overline{\mathbf{U}}=\Pi^{\leq N_{0}}\overline{\mathbf{U}}.$ (41)
3.3 The Operator $\mathbb{A}$
In this section, we focus our attention to the operator $\mathbb{A}:L^{2}_{e}\rightarrow\mathcal{H}_{e}$ required by Theorem 3.1. $\mathbb{A}$ is an approximate inverse of $D\mathbb{f}(\overline{\mathbf{u}})$. We now present its construction following a similar argument as in gs_cadiot_blanco . First we observe that $\mathbb{L}:\mathcal{H}_{e}\rightarrow L^{2}_{e}$ is an isometric isomorphism. Therefore building $\mathbb{A}:L^{2}_{e}\rightarrow\mathcal{H}_{e}$ is equivalent to building $\mathbb{B}:L^{2}_{e}\rightarrow L^{2}_{e}$ approximating the inverse of $D\mathbb{F}(\overline{\mathbf{u}})\mathbb{L}^{-1}$, and setting $\mathbb{A}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{L}^{-1}\mathbb{B}$. We will then construct $\mathbb{B}$ using Lemma 2.3 and a bounded linear operator on Fourier coefficients. First we construct $B^{N}$ a numerical approximate inverse to $\text{{$\Pi$}}^{\leq N}Df(\overline{\mathbf{U}})\mathscr{l}^{-1}\text{{$\Pi$}}^{\leq N}$. We choose $B^{N}$ satisfying $B^{N}=\text{{$\Pi$}}^{\leq N}B^{N}\text{{$\Pi$}}^{\leq N}$, hence it admits a matrix representation. We now describe further its construction. Due to the form of $g$, we obtain
$Df(\overline{\mathbf{U}})l^{-1}=\begin{bmatrix}I_{d}+Dg_{11}(\overline{\mathbf{U}})l_{\text{den}}^{-1}l_{22}-Dg_{12}(\overline{\mathbf{U}})l_{\text{den}}^{-1}l_{21}&-Dg_{11}(\overline{\mathbf{U}})l_{\text{den}}^{-1}l_{12}+Dg_{12}(\overline{\mathbf{U}})l_{\text{den}}^{-1}l_{11}\\ -Dg_{11}(\overline{\mathbf{U}})l_{\text{den}}^{-1}l_{22}+Dg_{12}(\overline{\mathbf{U}})l_{\text{den}}^{-1}l_{21}&I_{d}+Dg_{11}(\overline{\mathbf{U}})l_{\text{den}}^{-1}l_{12}-Dg_{12}(\overline{\mathbf{U}})l_{\text{den}}^{-1}l_{11}\end{bmatrix}$ (42)
where
$l_{\text{den}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}l_{11}l_{22}-l_{12}l_{21}.$ (43)
Notice that since (42) is a full matrix, we will choose $B^{N}$ to be a full matrix as well. More specifically, we construct $B^{N}$ and $B$ as follows
$B^{N}=\begin{bmatrix}B_{11}^{N}&B_{12}^{N}\\ B_{21}^{N}&B_{22}^{N}\end{bmatrix},\penalty 10000\ B=\text{{$\Pi$}}^{>N}+B^{N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}B_{11}&B_{12}\\ B_{21}&B_{22}\end{bmatrix}$ (44)
with $B_{11}^{N},B_{12}^{N},B_{21}^{N},$ and $B_{22}^{N}$ constructed numerically and satisfying $B_{1j}^{N}=\Pi^{\leq N}B_{1j}^{N}\Pi^{\leq N}$. Moreover, we define $B_{11}=B_{11}^{N}+\Pi^{>N}$, $B_{12}=B_{12}^{N},B_{21}=B_{21}^{N},B_{22}=B_{22}^{N}+\Pi^{>N}$ for convenience. Finally, we get $\mathbb{B}:L_{e}^{2}\rightarrow L_{e}^{2}$ and $\mathbb{A}:L_{e}^{2}\rightarrow\mathcal{H}_{e}$ as
$\mathbb{B}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}\mathbb{1}_{\mathbb{R}^{2}\backslash\Omega_{0}}&0\\ 0&\mathbb{1}_{\mathbb{R}^{2}\backslash\Omega_{0}}\end{bmatrix}+\mathbf{\Gamma}^{\dagger}(B),\penalty 10000\ \mathbb{A}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{l}^{-1}\mathbb{B},\penalty 10000\ \text{and}\penalty 10000\ \mathbb{B}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}\mathbb{B}_{11}&\mathbb{B}_{12}\\ \mathbb{B}_{21}&\mathbb{B}_{22}\end{bmatrix}$ (45)
where $\mathbb{B}_{ij}=\Gamma^{\dagger}(B_{ij})$ for $i,j\in\{1,2\}$. We will justify later on such a choice for $\mathbb{A}$ by computing explicitly the defect $\|I-\mathbb{A}D\mathbb{f}(\bar{\mathbf{u}})\|_{\mathcal{H}}$. In fact, we already obtain the following result, allowing to compute the operator norm of $\mathbb{A}$.
Lemma 3.2.
Let $\mathbb{A}:L^{2}_{e}\rightarrow\mathcal{H}_{e}$ be as in (45). Then,
$\|\mathbb{A}\|_{2,\mathcal{H}}=\|\mathbb{B}\|_{2}=\max\{1,\|B^{N}\|_{2}\}.$ (46)
Proof.
The proof can be found in unbounded_domain_cadiot . ∎
With $\mathbb{A}$ now built, we can begin to compute the bounds necessary to apply Theorem 3.1. This will be the goal of the next section.
4 Computing the Bounds for Localized Patterns
In this section, we provide explicit formulas for the bounds $\mathcal{Y}_{0},\mathcal{Z}_{1},$ and $\mathcal{Z}_{2}$ in Theorem 3.1. Before we begin the computation of the bounds, we recall some preliminary notations and results from gs_cadiot_blanco . First, given $u\in L^{\infty}(\mathbb{R}^{2})$, denote by
$\mathbb{u}\colon L^{2}(\mathbb{R}^{2})$ $\to L^{2}(\mathbb{R}^{2})$ (47)
$v$ $\mapsto uv$ (48)
the multiplication operator associated to $u$. Similarly, given $U=(u_{n})_{n\in\mathbb{N}_{0}}\in\ell^{1}(\mathbb{N}_{0})$, denote by
$\mathbb{U}:\ell^{2}(\mathbb{N}_{0})$ $\to\ell^{2}(\mathbb{N}_{0})$ (49)
$V$ $\mapsto U*V$ (50)
the discrete convolution operator associated to $U$. We now define the following map $\varphi:\mathbb{R}^{4}\to\mathbb{R}$ as
$\varphi(x_{1},x_{2},x_{3},x_{4})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\min\left\{\max\{x_{1},x_{4}\}+\max\{x_{2},x_{3}\},\penalty 10000\ \sqrt{x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}}\right\}$ (51)
for a given $(x_{1},x_{2},x_{3},x_{4})\in\mathbb{R}^{4}$. For the purpose of our analysis, we recall Lemma 4.2 from gs_cadiot_blanco which allows us to estimate the norm of an operator by block in terms of the norms of its individual blocks.
Lemma 4.1.
Let $K_{1},K_{2},K_{3},K_{4}\in\mathcal{B}(X)$ for $X\in\left\{L_{e}^{2}(\mathbb{R}^{2}),\ell_{e}^{2}\right\}$ where $\mathcal{B}(X)$ is the set of bounded linear operators on $X$. Then,
$\left\|\begin{bmatrix}K_{1}&K_{2}\\ K_{3}&K_{4}\\ \end{bmatrix}\right\|_{2}\leq\varphi(\left\|K_{1}\right\|_{2},\left\|K_{2}\right\|_{2},\left\|K_{3}\right\|_{2},\left\|K_{4}\right\|_{2}),$
where $\varphi:\mathbb{R}^{4}\to\mathbb{R}$ is defined as in (51).
Proof.
The proof can be found in gs_cadiot_blanco . ∎
Now, we provide explicit formulas for the bounds of Theorem 3.1 in the following list of lemmas. In particular, each formula relies on finite dimensional computations (that is vector or matrix norms), which can be rigorously evaluated thanks to computer-assisted techniques (cf. julia_blanco_fassler ). This justifies how specific choices for the approximate objects $\bar{\mathbf{u}}$ and $\mathbb{A}$.
Lemma 4.2.
Let $\mathcal{Y}_{0}>0$ be defined as
$\mathcal{Y}_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|^{\frac{1}{2}}\left(\|B^{N}f(\overline{\mathbf{U}})\|_{2}^{2}+\|(\text{{$\Pi$}}^{\leq N_{0}}-\text{{$\Pi$}}^{\leq N})l\overline{\mathbf{U}}+(\text{{$\Pi$}}^{\leq 3N_{0}}-\text{{$\Pi$}}^{\leq N})g(\overline{\mathbf{U}})\|_{2}^{2}\right)^{\frac{1}{2}}.$ (52)
Then $\|\mathbb{A}\mathbb{f}(\overline{\mathbf{u}})\|_{\mathcal{H}}\leq\mathcal{Y}_{0}.$
Proof.
We follow the same steps as those of Lemma 4.3 in gs_cadiot_blanco . In particular, we combine (13),the operator $\mathbb{A}$, and the properties of $\gamma$ to get
$\|\mathbb{A}\mathbb{f}(\overline{\mathbf{u}})\|_{\mathcal{H}}$ $=\|\mathbb{B}\mathbb{f}(\overline{\mathbf{u}})\|_{2}=|{\Omega_{0}}|^{\frac{1}{2}}\|\left(Bf(\overline{\mathbf{U}})\right)\|_{2}.$ (53)
We then obtain the result using the steps of Lemma 4.11 in unbounded_domain_cadiot . ∎
Lemma 4.3.
Let $\kappa>0$ be defined as in Lemma 2.2, Moreover, define $q_{1},q_{2}\in L^{\infty}(\mathbb{R})\cap L^{2}(\mathbb{R})$ as
$q_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\overline{u}_{2}+2\nu_{4}\mathbb{1}_{{\Omega_{0}}},\penalty 10000\ q_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\overline{u}_{1}+\nu_{5}\mathbb{1}_{{\Omega_{0}}}$
and their associated Fourier Coefficients $Q_{1}=\gamma(q_{1}),Q_{2}=\gamma(q_{2})$. Let $\mathcal{Z}_{2}:[0,\infty)\rightarrow[0,\infty)$ be defined as
$\mathcal{Z}_{2}(r)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sqrt{5}\kappa\sqrt{\varphi\left(\mathcal{Z}_{2,1},\mathcal{Z}_{2,2},\mathcal{Z}_{2,2},\mathcal{Z}_{2,3}\right)}+3\kappa\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}\max\left\{1,\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\right\|_{2}\right\}r$
for all $r\geq 0$, where
$\mathcal{Z}_{2,1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\|\ \begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\left(\mathbb{Q}_{1}^{2}+\mathbb{Q}_{2}^{2}\right)\begin{bmatrix}(B_{11}^{N}-B_{12}^{N})^{*}&(B_{21}^{N}-B_{22}^{N})^{*}\end{bmatrix}\right\|_{2},$
$\mathcal{Z}_{2,2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sqrt{\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\left(\mathbb{Q}_{1}^{2}+\mathbb{Q}_{2}^{2}\right)\Pi^{>N}\left(\mathbb{Q}_{1}^{2}+\mathbb{Q}_{2}^{2}\right)\begin{bmatrix}(B_{11}^{N}-B_{12}^{N})^{*}&(B_{21}^{N}-B_{22}^{N})^{*}\end{bmatrix}\right|_{2}},$
$\mathcal{Z}_{2,3}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\|\mathbb{Q}_{1}^{2}+\mathbb{Q}_{2}^{2}\|_{1}$
and $\varphi$ is defined in Lemma 4.1. Then $\|\mathbb{A}(D\mathbb{F}(\mathbf{h}+\bar{\mathbf{u}})-D\mathbb{F}(\bar{\mathbf{u}}))\|_{\mathcal{H}}\leq\mathcal{Z}_{2}(r)r$ for all $r>0$ and $\mathbf{h}\in\overline{B_{r}(0)}\subset H_{e}^{l}$
Proof.
Let $\mathbf{h}\in\overline{B_{r}(0)},\mathbf{z}\in\overline{B_{1}(0)}\subset\mathcal{H}_{e}$. Let $g_{3}:\left(\mathcal{H}_{e}\right)^{3}\rightarrow\mathcal{H}_{e}$ be defined as the following multilinear operator
$g_{3}(\mathbf{u},\mathbf{v},\mathbf{w})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}u_{1}v_{1}w_{2}\\ -u_{1}v_{1}w_{2}\end{bmatrix}\text{ for all }\mathbf{u}=(u_{1},u_{2}),\mathbf{v}=(v_{1},v_{2}),\mathbf{w}=(w_{1},w_{2})\in\mathcal{H}_{e}.$ (54)
In fact, notice that $g_{3}(\mathbf{u},\mathbf{u},\mathbf{u})=\mathbb{g}_{3}(\mathbf{u})$. Moreover, we have the symmetry $g_{3}(\mathbf{u},\mathbf{v},\mathbf{w})=g_{3}(\mathbf{v},\mathbf{u},\mathbf{w})$ for all $\mathbf{u},\mathbf{v},\mathbf{w}\in\mathcal{H}_{e}$. Following similar steps to those done in Lemma 4.5 of gs_cadiot_blanco , we obtain
$\|\mathbb{A}(D\mathbb{F}(\mathbf{h}+\bar{\mathbf{u}})-D\mathbb{F}(\bar{\mathbf{u}}))\|_{\mathcal{H}}\leq\|\mathbb{B}D^{2}\mathbb{g}(\overline{\mathbf{u}})(\mathbf{h},\mathbf{z})\|_{2}+\|\mathbb{B}(2g_{3}(\mathbf{z},\mathbf{h},\mathbf{h})+g_{3}(\mathbf{h},\mathbf{h},\mathbf{z}))\|_{2}$ (55)
$\leq\left\|\mathbb{B}\begin{bmatrix}(\mathbb{q}_{1}\mathbb{h}_{1}+\mathbb{q}_{2}\mathbb{h}_{2})z_{1}+\mathbb{q}_{2}\mathbb{h}_{1}z_{2}\\ -(\mathbb{q}_{1}\mathbb{h}_{1}+\mathbb{q}_{2}\mathbb{h}_{2})z_{1}-\mathbb{q}_{2}\mathbb{h}_{1}z_{2}\end{bmatrix}\right\|_{2}+\left\|\begin{bmatrix}\mathbb{B}_{11}-\mathbb{B}_{12}\\ \mathbb{B}_{21}-\mathbb{B}_{22}\end{bmatrix}\begin{bmatrix}2z_{1}h_{1}h_{2}+h_{1}^{2}z_{2}\end{bmatrix}\right\|_{2}$ (56)
$\leq\left\|\mathbb{B}\begin{bmatrix}\mathbb{q}_{1}&\mathbb{q}_{2}\\ -\mathbb{q}_{1}&-\mathbb{q}_{2}\end{bmatrix}\begin{bmatrix}\mathbb{h}_{1}&\mathbb{h}_{2}\\ \mathbb{h}_{2}&\mathbb{h}_{1}\end{bmatrix}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}\right\|_{2}+\left\|\begin{bmatrix}\mathbb{B}_{11}-\mathbb{B}_{12}\\ \mathbb{B}_{21}-\mathbb{B}_{22}\end{bmatrix}\right\|_{2}\|2z_{1}h_{1}h_{2}+h_{1}^{2}z_{2}\|_{2}.$ (57)
Now, observe that
$\|2z_{1}h_{1}h_{2}+h_{1}^{2}z_{2}\|_{2}\leq 2\|z_{1}h_{2}h_{2}\|_{2}+\|h_{1}^{2}z_{2}\|_{2}\leq 3\kappa\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}r^{2}.$ (58)
Hence, we obtain
$\left\|\mathbb{B}\begin{bmatrix}\mathbb{q}_{1}&\mathbb{q}_{2}\\ -\mathbb{q}_{1}&-\mathbb{q}_{2}\end{bmatrix}\begin{bmatrix}\mathbb{h}_{1}&\mathbb{h}_{2}\\ \mathbb{h}_{2}&\mathbb{h}_{1}\end{bmatrix}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}\right\|_{2}+\left\|\begin{bmatrix}\mathbb{B}_{11}-\mathbb{B}_{12}\\ \mathbb{B}_{21}-\mathbb{B}_{22}\end{bmatrix}\right\|_{2}\|2z_{1}h_{1}h_{2}+h_{1}^{2}z_{2}\|_{2}$ (59)
$\leq\left\|\begin{bmatrix}\mathbb{B}_{11}-\mathbb{B}_{12}\\ \mathbb{B}_{21}-\mathbb{B}_{22}\end{bmatrix}\begin{bmatrix}\mathbb{q}_{1}&\mathbb{q}_{2}\end{bmatrix}\right\|_{2}\left\|\begin{bmatrix}\mathbb{h}_{1}&\mathbb{h}_{2}\\ \mathbb{h}_{2}&\mathbb{h}_{1}\end{bmatrix}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}\right\|_{2}+3\kappa\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}\max\left\{1,\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\right\|_{2}\right\}r^{2}$ (60)
$\leq\sqrt{5}\kappa\sqrt{\left\|\begin{bmatrix}\mathbb{B}_{11}-\mathbb{B}_{12}\\ \mathbb{B}_{21}-\mathbb{B}_{22}\end{bmatrix}(\mathbb{q}_{1}^{2}+\mathbb{q}_{2}^{2})\begin{bmatrix}(\mathbb{B}_{11}-\mathbb{B}_{12})^{*}&(\mathbb{B}_{21}-\mathbb{B}_{22})^{*}\end{bmatrix}\right\|_{2}}r+3\kappa\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}\max\left\{1,\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\right\|_{2}\right\}r^{2}$ (62)
where we also used Lemma 2.2 in the last inequality. Then, using the properties of $\Gamma$, we similar steps to those used in Lemma 4.5 of gs_cadiot_blanco to estimate
$\left\|\begin{bmatrix}\mathbb{B}_{11}-\mathbb{B}_{12}\\ \mathbb{B}_{21}-\mathbb{B}_{22}\end{bmatrix}(\mathbb{q}_{1}^{2}+\mathbb{q}_{2}^{2})\begin{bmatrix}(\mathbb{B}_{11}-\mathbb{B}_{12})^{*}&(\mathbb{B}_{21}-\mathbb{B}_{22})^{*}\end{bmatrix}\right\|_{2}$ $\leq\varphi(\mathcal{Z}_{2,1},\mathcal{Z}_{2,2},\mathcal{Z}_{2,2},\mathcal{Z}_{2,3})$ (63)
and conclude the result. ∎
4.1 The Bound $\mathcal{Z}_{1}$
For now, we provided explicit formulas for the bounds $\mathcal{Y}$ and $\mathcal{Z}_{2}$, and it remains to estimate $\|I-\mathbb{A}D\mathbb{f}(\bar{\mathbf{u}})\|_{\mathcal{H}}\leq\mathcal{Z}_{1}$. This bound requires a bit more analysis, and we present its treatment in this section. First let us define $v_{1},v_{2}\in L^{\infty}(\mathbb{R})\cap L^{2}(\mathbb{R})$
$v_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\overline{u}_{1}\overline{u}_{2}+2\nu_{4}\overline{u}_{1}+\nu_{5}\overline{u}_{2}\text{ and }v_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{u}_{1}^{2}+\nu_{5}\overline{u}_{1}.$ (64)
In particular, we can write $D\mathbb{g}(\bar{\mathbf{u}})=\begin{bmatrix}\mathbb{v_{1}}&\mathbb{v_{2}}\\ -\mathbb{v}_{1}&-\mathbb{v}_{2}\end{bmatrix}$, where $\mathbb{v}_{1},\mathbb{v}_{2}$ are defined as in (47). Now we define
$V_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma(v_{1}),V_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma(v_{2}),V_{1}^{N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\Pi^{\leq N}V_{1},V_{2}^{N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\Pi^{\leq N}V_{2}\text{ and }Dg^{N}(\overline{\mathbf{U}})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}\mathbb{V}_{1}^{N}&\mathbb{V}_{2}^{N}\\ -\mathbb{V}_{1}^{N}&-\mathbb{V}_{2}^{N}\end{bmatrix}$
where $\mathbb{V}_{1}^{N},\mathbb{V}_{2}^{N}$ are the discrete convolution operators associated to $V_{1}^{N},V_{2}^{N}$. Finally we define
$D\mathbb{g}^{N}(\overline{\mathbf{U}})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\Gamma^{\dagger}(Dg^{N}(\overline{\mathbf{U}})),\penalty 10000\ \penalty 10000\ v_{1}^{N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma^{\dagger}(V_{1}^{N})\text{ and }v_{2}^{N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\Gamma^{\dagger}(V_{2}^{N}).$
This notation allows us to decompose both $D\mathbb{g}(\overline{\mathbf{U}})$ and $Dg(\overline{\mathbf{U}})$ using a truncation of size $N$. We now state a lemma for the $\mathcal{Z}_{1}$ bound.
Lemma 4.4.
Let $\mathcal{Z}_{u},Z_{1},\mathcal{Z}_{1}>0$ be bounds satisfying
$\mathcal{Z}_{u}\geq\|\mathbb{B}D\mathbb{g}^{N}(\overline{\mathbf{u}})(\mathbb{l}^{-1}-\text{{$\Gamma$}}^{\dagger}(l^{-1})\|_{2},\penalty 10000\ Z_{1}\geq\|I-B(I+Dg^{N}(\overline{\mathbf{U}})l^{-1})\|_{2}$ (65)
$\mathcal{Z}_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}Z_{1}+\mathcal{Z}_{u}+\max\left\{1,\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\right\|_{2}\right\}\|\mathbb{l}^{-1}\|_{2}\sqrt{\|V_{1}^{N}-V_{1}\|_{1}^{2}+\|V_{2}^{N}-V_{2}\|_{1}^{2}}.$ (66)
Then, it follows that $\|I-{\mathbb{A}}D\mathbb{f}(\overline{\mathbf{u}})\|_{\mathcal{H}}\leq\mathcal{Z}_{1}.$
Proof.
First, using the definition of the norm $\|\cdot\|_{\mathcal{H}}$ and the triangle inequality, we get
$\|I-\mathbb{A}D\mathbb{f}(\overline{\mathbf{u}})\|_{\mathcal{H}}=\|I-\mathbb{B}(I+D\mathbb{g}(\overline{\mathbf{u}})\mathbb{l}^{-1})\|_{2},$ $\leq\|I-\mathbb{B}(I+D\mathbb{g}^{N}(\overline{\mathbf{u}})\mathbb{l}^{-1})\|_{2}+\|\mathbb{B}(D\mathbb{g}^{N}(\overline{\mathbf{u}})-D\mathbb{g}(\overline{\mathbf{u}}))\mathbb{l}^{-1}\|_{2}$ (67)
$\leq Z_{1}+\mathcal{Z}_{u}+\left\|\begin{bmatrix}\mathbb{B}_{11}-\mathbb{B}_{12}\\ \mathbb{B}_{21}-\mathbb{B}_{22}\end{bmatrix}\begin{bmatrix}\mathbb{V}_{1}^{N}-\mathbb{V}_{1}&\mathbb{V}_{2}^{N}-\mathbb{V}_{2}\end{bmatrix}\mathbb{l}^{-1}\right\|_{2}$ (68)
where the final step followed from Theorem 3.5 of unbounded_domain_cadiot and the definition of $\mathbb{B}$. For the final term in (68), we use submultiplicativity to obtain
$\left\|\begin{bmatrix}\mathbb{B}_{11}-\mathbb{B}_{12}\\ \mathbb{B}_{21}-\mathbb{B}_{22}\end{bmatrix}\begin{bmatrix}\mathbb{V}_{1}^{N}-\mathbb{V}_{1}&\mathbb{V}_{2}^{N}-\mathbb{V}_{2}\end{bmatrix}\mathbb{l}^{-1}\right\|_{2}$ $\leq\left\|\begin{bmatrix}\mathbb{B}_{11}-\mathbb{B}_{12}\\ \mathbb{B}_{21}-\mathbb{B}_{22}\end{bmatrix}\right\|_{2}\|\mathbb{l}^{-1}\|_{2}\|\begin{bmatrix}\mathbb{V}_{1}^{N}-\mathbb{V}_{1}&\mathbb{V}_{2}^{N}-\mathbb{V}_{2}\end{bmatrix}\|_{2}$ (69)
$\hskip-65.44142pt\leq\max\left\{1,\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\right\|_{2}\right\}\|\mathbb{l}^{-1}\|_{2}\sqrt{\|V_{1}^{N}-V_{1}\|_{1}^{2}+\|V_{2}^{N}-V_{2}\|_{1}^{2}}$ (70)
where the last step followed by Parseval’s Identity and Lemma 4.6 of gs_cadiot_blanco . ∎
Now, we must compute the bounds $Z_{1}$ and $\mathcal{Z}_{u}$. We will begin with $Z_{1}$.
Lemma 4.5.
Let
$Z_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sqrt{Z_{0}^{2}+2Z_{1,1}^{2}+2Z_{1,2}^{2}}$ (71)
where
$Z_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\|\text{{$\Pi$}}^{\leq 2N}-\text{{$\Pi$}}^{\leq 3N}B(I_{d}+Dg^{N}(\overline{\mathbf{U}})l^{-1})\text{{$\Pi$}}^{\leq 2N}\|_{2}$
$Z_{1,1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\frac{l_{22}}{l_{\mathrm{den}}}\right)_{2N}\|V_{1}^{N}\|_{1}+\left(\frac{l_{21}}{l_{\mathrm{den}}}\right)_{2N}\|V_{2}^{N}\|_{1},\penalty 10000\ \penalty 10000\ Z_{1,2}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\frac{l_{12}}{l_{\mathrm{den}}}\right)_{2N}\|V_{1}^{N}\|_{1}+\left(\frac{l_{11}}{l_{\mathrm{den}}}\right)_{2N}\|V_{2}^{N}\|_{1}$
and $\left(\frac{\mathscr{l}_{ij}}{l_{\mathrm{den}}}\right)_{2N}=\max_{n\in\mathbb{Z}\setminus I^{2N}}\frac{\mathscr{l}_{ij}(\tilde{n})}{\mathscr{l}_{den}(\tilde{n})}$. Then, it follows that $\|I_{d}-B(I_{d}+Dg^{N}(\overline{U})l^{-1})\|_{2}\leq Z_{1}$.
Proof.
To begin, we introduce a truncation.
$\|I_{d}-B(I_{d}+Dg^{N}(\overline{\mathbf{U}})l^{-1})\|_{2}^{2}$ $\leq\|\text{{$\Pi$}}^{>2N}-B(I_{d}+Dg^{N}(\overline{\mathbf{U}})l^{-1})\text{{$\Pi$}}^{>2N}\|_{2}^{2}+\|\text{{$\Pi$}}^{\leq 2N}-B(I_{d}+Dg^{N}(\overline{\mathbf{U}})l^{-1})\text{{$\Pi$}}^{\leq 2N}\|_{2}^{2}$ (72)
$=\|\text{{$\Pi$}}^{>2N}-B\text{{$\Pi$}}^{>2N}-BDg^{N}(\overline{\mathbf{U}})\text{{$\Pi$}}^{<2N}l^{-1}\|_{2}^{2}+Z_{0}^{2}.$ (73)
Now, by definition of $B$, we have $B\text{{$\Pi$}}^{>2N}=\text{{$\Pi$}}^{>2N}$ and $B\text{{$\Pi$}}^{>N}=\text{{$\Pi$}}^{>N}$. Additionally, since $v_{1}^{N}$ and $v_{2}^{N}$ are of size $N$, it follows that $Dg^{N}(\overline{U})\text{{$\Pi$}}^{>2N}=\text{{$\Pi$}}^{>N}Dg^{N}(\overline{U})\text{{$\Pi$}}^{>2N}$. Hence, we have
$\|\text{{$\Pi$}}^{>2N}-B\text{{$\Pi$}}^{>2N}-BDg^{N}(\overline{U})\text{{$\Pi$}}^{<2N}l^{-1}\|_{2}$ $=\|\text{{$\Pi$}}^{>2N}-\text{{$\Pi$}}^{>2N}-B\text{{$\Pi$}}^{>N}Dg^{N}(\overline{U})\text{{$\Pi$}}^{<2N}l^{-1}\|_{2}$ (74)
$=\|\text{{$\Pi$}}^{>N}Dg^{N}(\overline{U})l^{-1}\text{{$\Pi$}}^{<2N}\|_{2}$ (75)
$\hskip-28.45274pt=\left\|\begin{bmatrix}\mathbb{V}_{1}^{N}l_{22}l_{\mathrm{den}}^{-1}-\mathbb{V}_{2}^{N}l_{21}l_{\mathrm{den}}^{-1}&-\mathbb{V}_{1}^{N}l_{12}l_{\mathrm{den}}^{-1}+\mathbb{V}_{2}^{N}l_{11}l_{\mathrm{den}}^{-1}\\ -\mathbb{V}_{1}^{N}l_{22}l_{\mathrm{den}}^{-1}+\mathbb{v}_{2}^{N}l_{21}l_{\mathrm{den}}^{-1}&\mathbb{V}_{1}^{N}l_{12}l_{\mathrm{den}}^{-1}-\mathbb{V}_{2}^{N}l_{11}l_{\mathrm{den}}^{-1}\end{bmatrix}\right\|_{2}$ (76)
$\hskip-42.67912pt\leq\sqrt{2}\sqrt{\|\mathbb{V}_{1}^{N}l_{22}l_{\mathrm{den}}^{-1}-\mathbb{V}_{2}^{N}l_{21}l_{\mathrm{den}}^{-1}\|_{2}^{2}+\|-\mathbb{V}_{1}^{N}l_{12}l_{\mathrm{den}}^{-1}+\mathbb{V}_{2}^{N}l_{11}l_{\mathrm{den}}^{-1}\|_{2}^{2}}$ (77)
where we used the Cauchy-Schwarz inequality. We now examine each term
$\|\mathbb{V}_{1}^{N}l_{22}l_{\mathrm{den}}^{-1}-\mathbb{v}_{2}^{N}l_{21}l_{\mathrm{den}}^{-1}\|_{2}\leq\left(\frac{l_{22}}{l_{\mathrm{den}}}\right)_{2N}\|V_{1}^{N}\|_{1}+\left(\frac{l_{21}}{l_{\mathrm{den}}}\right)_{2N}\|V_{2}^{N}\|_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}Z_{1,1}$ (78)
$\|-\mathbb{V}_{1}^{N}l_{12}l_{\mathrm{den}}^{-1}+\mathbb{V}_{2}^{N}l_{11}l_{\mathrm{den}}^{-1}\|_{2}\leq\left(\frac{l_{12}}{l_{\mathrm{den}}}\right)_{2N}\|V_{1}^{N}\|_{1}+\left(\frac{l_{11}}{l_{\mathrm{den}}}\right)_{2N}\|V_{2}^{N}\|_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}Z_{1,2}.$ (79)
Therefore, we have
$\|I_{d}-B(I_{d}+Dg^{N}(\overline{U})l^{-1})\|_{2}\leq\sqrt{Z_{0}^{2}+2Z_{1,1}^{2}+2Z_{1,2}^{2}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}Z_{1}$ (80)
as desired. ∎
We now work on the bound $\mathcal{Z}_{u}$. We first need some preliminary results.
Lemma 4.6.
Assume that Assumption 1 holds, then we have the following disjunction :
  1. 1.
    If $(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})<0$, then define $z_{1},z_{2}$ as
    $z_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\pi iy\text{ and }z_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}-2\pi i\bar{y},\text{ where }y\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2\pi}\left(\frac{\nu_{1}-\nu_{2}\nu+i\sqrt{-(\nu\nu_{2}-\nu_{1})^{2}-4\nu\nu_{2}(\nu_{1}+\nu_{3})}}{2\nu}\right)^{\frac{1}{2}}.$ (81)
  2. 2.
    If $(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})\geq 0$ and $\nu\nu_{2}-\nu_{1}>\sqrt{(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})}$, then define $z_{1},z_{2}$ as
    $z_{1}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\frac{\nu\nu_{2}-\nu_{1}+\sqrt{(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})}}{2\nu}\right)^{\frac{1}{2}}$ (82)
    $z_{2}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\frac{\nu\nu_{2}-\nu_{1}-\sqrt{(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})}}{2\nu}\right)^{\frac{1}{2}}.$ (83)
In both cases, for $d_{1},d_{2}\in\mathbb{R}$, we obtain that
$\left|\mathcal{F}^{-1}\left(\frac{d_{1}|2\pi\xi|^{2}+d_{2}}{\mathscr{l}_{den}(\xi)}\right)(x)\right|\leq C_{0}e^{-a|x|},$
with
$a=\min\{\mathrm{Re}(z_{1}),\mathrm{Re}(z_{2})\}\text{ and }C_{0}(d_{1},d_{2})=\frac{1}{|2\nu(z_{1}^{2}-z_{2}^{2})|}\left(|d_{1}z_{2}|+\frac{|d_{2}|}{|z_{2}|}\right)$ (84)
Proof.
Case 1: $(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})<0$. In this case, $\ell_{\text{den}}$ possesses 4 complex roots $\pm y$ and $\pm\overline{y}$, where $y$ is given in (81).
Case 2: $(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})\geq 0$ and $\nu\nu_{2}-\nu_{1}>\sqrt{(\nu\nu_{2}-\nu_{1})^{2}+4\nu\nu_{2}(\nu_{1}+\nu_{3})}$. In this case, $\ell_{\text{den}}$ possesses 4 imaginary roots $\pm y_{1}$ and $\pm y_{2}$ where $y_{1}=\frac{z_{1}}{2\pi i}$ and $y_{2}=\frac{z_{2}}{2\pi i}$.
In any case, we define
$f(x)=\mathcal{F}^{-1}\left(\frac{1}{\ell_{\text{den}}(\xi)}\right)(x)=\mathcal{F}^{-1}\left(\frac{1}{((2\pi\xi)^{2}+z_{1}^{2})((2\pi\xi)^{2}+z_{2}^{2})}\right)(x)$
and using that
$\mathcal{F}^{-1}\left(\frac{1}{|2\pi\xi^{2}|+{z}^{2}}\right)(t)=\frac{e^{-z|t|}}{2z},$ (85)
we obtain
$f(t)=\frac{1}{2\nu({z}_{2}^{2}-z_{1}^{2})}\left(\frac{e^{-z_{1}|t|}}{z_{1}}-\frac{e^{-{z}_{2}|t|}}{{z}_{2}}\right).$ (86)
In the case 1., we have that $z_{2}=\bar{z_{1}}$, this implies that
$|f(t)|\leq\frac{e^{-a|t|}}{|2\nu z_{2}\nu({z}_{1}^{2}-z_{2}^{2})|}$
where $a=\min\{\mathrm{Re}(z_{1}),\mathrm{Re}(z_{2})\}$. Now, since $\frac{(2\pi\xi)^{2}}{\mathscr{l}_{den}(\xi)}\in L^{1}(\mathbb{R})$, we have that $f$ is twice continuously differentiable. In particular, we have
$f^{\prime\prime}(t)=\mathcal{F}^{-1}\left(\frac{(2\pi\xi)^{2}}{\nu(|2\pi\xi^{2}|+{z}_{1}^{2})(|2\pi\xi^{2}|+{z}_{2}^{2})}\right)(t).$
In fact, using the above, we readily have that
$|f^{\prime\prime}(t)|\leq\frac{|z_{2}|e^{-a|t|}}{|\sqrt{2}\nu({z}_{1}^{2}-z_{2}^{2})|}.$
For the case 2., notice that $z_{1}>z_{2}>0$. This implies that
$0<f(t)=\frac{1}{2\nu({z}_{1}^{2}-z_{2}^{2})}\left(\frac{e^{-z_{2}|t|}}{z_{2}}-\frac{e^{-{z}_{1}|t|}}{{z}_{1}}\right)\leq\frac{e^{-z_{2}|t|}}{2\nu z_{2}({z}_{1}^{2}-z_{2}^{2})}.$ (87)
Similarly, we get
$|f^{\prime\prime}(t)|\leq\frac{z_{2}e^{-z_{2}|t|}}{2\nu({z}_{1}^{2}-z_{2}^{2})}.$ (88)
We obtain the desired result by observing that:
$\left|\mathcal{F}^{-1}\left(\frac{d_{1}|2\pi\xi|^{2}+d_{2}}{\mathscr{l}_{den}(\xi)}\right)(t)\right|\leq|d_{1}f^{\prime\prime}(t)|+|d_{2}f(t)|\leq C_{0}e^{-a|t|}.$
Now, we break down the $\mathcal{Z}_{u}$ bound as follows.
Lemma 4.7.
Let $\left(\mathcal{Z}_{u,k,j}\right)_{k\in\{1,2\},j\in\{1,2,3,4\}}$ be bounds satisfying
$\mathcal{Z}_{u,1,1}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{22}\mathbb{l}_{den}^{-1}\mathbb{v}_{1}^{N}\|_{2},\penalty 10000\ \mathcal{Z}_{u,2,1}\geq\|\mathbb{1}_{{\Omega_{0}}}(\mathbb{l}_{22}\mathbb{l}_{den}^{-1}-\Gamma^{\dagger}(l_{22}l_{\mathrm{den}}^{-1}))\mathbb{v}_{1}^{N}\|_{2}$ (89)
$\mathcal{Z}_{u,1,2}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{21}\mathbb{l}_{den}^{-1}\mathbb{v}_{2}^{N}\|_{2},\penalty 10000\ \mathcal{Z}_{u,2,2}\geq\|\mathbb{1}_{{\Omega_{0}}}(\mathbb{l}_{21}\mathbb{l}_{den}^{-1}-\Gamma^{\dagger}(l_{21}l_{\mathrm{den}}^{-1}))\mathbb{v}_{2}^{N}\|_{2}$ (90)
$\mathcal{Z}_{u,1,3}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{12}\mathbb{l}_{den}^{-1}\mathbb{v}_{1}^{N}\|_{2},\penalty 10000\ \mathcal{Z}_{u,2,3}\geq\|\mathbb{1}_{{\Omega_{0}}}(\mathbb{l}_{12}\mathbb{l}_{den}^{-1}-\Gamma^{\dagger}(l_{12}l_{\mathrm{den}}^{-1}))\mathbb{v}_{1}^{N}\|_{2}$ (91)
$\mathcal{Z}_{u,1,4}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{22}\mathbb{l}_{den}^{-1}\mathbb{v}_{2}^{N}\|_{2},\penalty 10000\ \mathcal{Z}_{u,2,4}\geq\|\mathbb{1}_{{\Omega_{0}}}(\mathbb{l}_{11}\mathbb{l}_{den}^{-1}-\Gamma^{\dagger}(l_{11}l_{\mathrm{den}}^{-1}))\mathbb{v}_{2}^{N}\|_{2}$ (92)
Moreover, given $k\in\{1,2\}$, define $\mathcal{Z}_{u,k}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sqrt{2}\sqrt{(\mathcal{Z}_{u,k,1}+\mathcal{Z}_{u,k,2})^{2}+(\mathcal{Z}_{u,k,3}+\mathcal{Z}_{u,k,4})^{2}}$. Then, it follows that $\mathcal{Z}_{u,1}$ and $\mathcal{Z}_{u,2}$ satisfy
$\mathcal{Z}_{u,1}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}D\mathbb{g}(\overline{u})^{N}\mathbb{l}^{-1}\|_{2},\penalty 10000\ \mathcal{Z}_{u,2}\geq\|\mathbb{1}_{{\Omega_{0}}}D\mathbb{g}(\overline{u})^{N}(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})\|_{2}.$ (93)
Finally, defining
$\mathcal{Z}_{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sqrt{\mathcal{Z}_{u,1}^{2}+\mathcal{Z}_{u,2}^{2}},$ (94)
it follows that $\|\mathbb{B}D\mathbb{g}(\overline{u})^{N}(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})\|_{2}\leq\max\left\{1,\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\right\|_{2}\right\}\mathcal{Z}_{u}$.
Proof.
To begin, we use the definitions of (64) and $\mathbb{B}$ to see that
$\left\|\mathbb{B}D\mathbb{g}(\overline{u})^{N}(\mathbf{u}_{0})\left(\text{{$\Gamma$}}^{\dagger}\left(l^{-1}\right)-\mathbb{l}^{-1}\right)\right\|_{2}$ $\leq\max\left\{1,\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\right\|_{2}\right\}\mathcal{Z}_{u}$ (95)
$=\max\left\{1,\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\right\|_{2}\right\}\sqrt{\mathcal{Z}_{u,1}^{2}+\mathcal{Z}_{u,2}^{2}}.$ (96)
Now, we examine $\mathcal{Z}_{u,1}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{m}^{N}(\overline{\mathbf{x}})\mathbb{l}^{-1}\|_{2}.$ To do so, let $\mathbf{u}=(u_{1},u_{2})\in L^{2}$ such that $\|\mathbf{u}\|_{2}=1$. Then, observe that
$\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}^{-*}(D\mathbb{g}(\overline{u})^{N})^{*}\mathbf{u}\|_{2}$ $=\left\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\begin{bmatrix}\mathbb{l}_{22}\mathbb{l}_{den}^{-1}&-\mathbb{l}_{21}\mathbb{l}_{den}^{-1}\\ -\mathbb{l}_{12}\mathbb{l}_{den}^{-1}&\mathbb{l}_{11}\mathbb{l}_{den}^{-1}\end{bmatrix}\begin{bmatrix}\mathbb{v}_{1}^{N}&-\mathbb{v}_{1}^{N}\\ \mathbb{v}_{2}^{N}&-\mathbb{v}_{2}^{N}\end{bmatrix}\begin{bmatrix}u_{1}\\ u_{2}\end{bmatrix}\right\|_{2}$
$=\left\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\begin{bmatrix}\mathbb{l}_{22}\mathbb{l}_{den}^{-1}&-\mathbb{l}_{21}\mathbb{l}_{den}^{-1}\\ -\mathbb{l}_{12}\mathbb{l}_{den}^{-1}&\mathbb{l}_{11}\mathbb{l}_{den}^{-1}\end{bmatrix}\begin{bmatrix}\mathbb{v}_{1}^{N}(u_{1}-u_{2})\\ \mathbb{v}_{2}^{N}(u_{1}-u_{2})\end{bmatrix}\right\|_{2}$
$=\left\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\begin{bmatrix}\mathbb{l}_{22}\mathbb{l}_{den}^{-1}\mathbb{v}_{1}^{N}(u_{1}-u_{2})-\mathbb{l}_{21}\mathbb{l}_{den}^{-1}\mathbb{v}_{2}^{N}(u_{1}-u_{2})\\ -\mathbb{l}_{12}\mathbb{l}_{den}^{-1}\mathbb{v}_{1}^{N}(u_{1}-u_{2})+\mathbb{l}_{11}\mathbb{l}_{den}^{-1}\mathbb{v}_{2}^{N}(u_{1}-u_{2})\end{bmatrix}\right\|_{2}$
$\leq\biggl(\left(\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{22}\mathbb{l}_{den}^{-1}\mathbb{v}_{1}^{N}(u_{1}-u_{2})\|_{2}+\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{21}\mathbb{l}_{den}^{-1}\mathbb{v}_{2}^{N}(u_{1}-u_{2})\|_{2}\right)^{2}$ (97)
$\hskip 56.9055pt+(\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{12}\mathbb{l}_{den}^{-1}\mathbb{v}_{1}^{N}(u_{1}-u_{2})\|_{2}+\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{11}\mathbb{l}_{den}^{-1}\mathbb{v}_{2}^{N}(u_{1}-u_{2})\|_{2})^{2}\biggr)^{\frac{1}{2}}.$ (98)
Now, observe that
$\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{ij}\mathbb{l}_{den}^{-1}\mathbb{v}_{k}^{N}(u_{1}-u_{2})\|_{2}\leq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{ij}\mathbb{l}_{den}^{-1}\mathbb{v}_{k}^{N}\|_{2}\|u_{1}-u_{2}\|_{2}$ $\leq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{ij}\mathbb{l}_{den}^{-1}\mathbb{v}_{k}^{N}\|_{2}\left\|\begin{bmatrix}1&-1\\ 0&0\end{bmatrix}\begin{bmatrix}u_{1}\\ u_{2}\end{bmatrix}\right\|_{2}$ (99)
$\leq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{ij}\mathbb{l}_{den}^{-1}\mathbb{v}_{k}^{N}\|_{2}\left\|\begin{bmatrix}1&-1\\ 0&0\end{bmatrix}\right\|_{2}\|\mathbf{u}\|_{2}$ (100)
$=\sqrt{2}\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{ij}\mathbb{l}_{den}^{-1}\mathbb{v}_{k}^{N}\|_{2}.$ (101)
Therefore, return to the previous step and obtain
$\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}^{-*}(D\mathbb{g}(\overline{\mathbf{u}})^{N})^{*}\mathbf{u}\|_{2}$ $\leq\biggl(\left(\sqrt{2}\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{22}\mathbb{l}_{den}^{-1}\mathbb{v}_{1}^{N}\|_{2}+\sqrt{2}\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{21}\mathbb{l}_{den}^{-1}\mathbb{v}_{2}^{N}\|_{2}\right)^{2}$ (102)
$\hskip 85.35826pt+(\sqrt{2}\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{12}\mathbb{l}_{den}^{-1}\mathbb{v}_{1}^{N}\|_{2}+\sqrt{2}\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{11}\mathbb{l}_{den}^{-1}\mathbb{v}_{2}^{N}\|_{2})^{2}\biggr)^{\frac{1}{2}}$ (103)
$\leq\biggl(2\left(\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{22}\mathbb{l}_{den}^{-1}\mathbb{v}_{1}^{N}\|_{2}+\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{21}\mathbb{l}_{den}^{-1}\mathbb{v}_{2}^{N}\|_{2}\right)^{2}$ (104)
$\hskip 85.35826pt+2(\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{12}\mathbb{l}_{den}^{-1}\mathbb{v}_{1}^{N}\|_{2}+\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{11}\mathbb{l}_{den}^{-1}\mathbb{v}_{2}^{N}\|_{2})^{2}\biggr)^{\frac{1}{2}}$ (105)
$\leq\sqrt{2}\sqrt{(\mathcal{Z}_{u,1,1}+\mathcal{Z}_{u,1,2})^{2}+(\mathcal{Z}_{u,1,3}+\mathcal{Z}_{u,1,4})^{2}}.$ (106)
By performing the same steps for $\mathcal{Z}_{u,2}$, we similarly obtain
$\|\mathbb{1}_{{\Omega_{0}}}D\mathbb{g}(\overline{u})^{N}(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})\|_{2}\leq 2\sqrt{(\mathcal{Z}_{u,2,1}+\mathcal{Z}_{u,2,2})^{2}+(\mathcal{Z}_{u,2,3}+\mathcal{Z}_{u,2,4})^{2}}.$ (107)
We now estimate $\mathcal{Z}_{u,1}$ and $\mathcal{Z}_{u,2}$.
Lemma 4.8.
Let $a$ be defined as in (84). Moreover, let $E\in\ell^{2}$ and $C(d),C_{1},C_{2},C_{3},C_{4}>0$ be defined as
$E\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma(\mathbb{1}_{{\Omega_{0}}}\cosh(2ax)),\penalty 10000\ C(d)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}4d+\frac{4e^{-ad}}{a(1-e^{-\frac{3ad}{2}})}+\frac{2}{a(1-e^{-2ad})},$ (108)
$C_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}C_{0}(-1,-\nu_{2}),\penalty 10000\ C_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}C_{0}(0,\nu_{2}),\penalty 10000\ C_{3}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}C_{0}(0,\nu_{3}),\penalty 10000\ \text{and}\penalty 10000\ C_{4}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}C_{0}(-\nu,\nu_{1}).$ (109)
Then, let $(\mathcal{Z}_{u,k,j})_{k\in\{1,2\},j\in\{1,2,3,4\}}>0$ be defined as
$\mathcal{Z}_{u,1,1}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\frac{C_{1}^{2}e^{-2ad}}{a}(V_{1}^{N},V_{1}^{N}*E)_{2},\penalty 10000\ \mathcal{Z}_{u,2,1}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{u,1,1}^{2}+e^{-4ad}C(d)C_{1}^{2}|{\Omega_{0}}|(V_{1}^{N},V_{1}^{N}*E)_{2}$ (110)
$\mathcal{Z}_{u,1,2}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\frac{C_{2}^{2}e^{-2ad}}{a}(V_{2}^{N},V_{2}^{N}*E)_{2},\penalty 10000\ \mathcal{Z}_{u,2,2}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{u,1,2}^{2}+e^{-4ad}C(d)C_{2}^{2}|{\Omega_{0}}|(V_{2}^{N},V_{2}^{N}*E)_{2}$ (111)
$\mathcal{Z}_{u,1,3}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\frac{C_{3}^{2}e^{-2ad}}{a}(V_{1}^{N},V_{1}^{N}*E)_{2},\penalty 10000\ \mathcal{Z}_{u,2,3}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{u,1,3}^{2}+e^{-4ad}C(d)C_{3}^{2}|{\Omega_{0}}|(V_{1}^{N},V_{1}^{N}*E)_{2}$ (112)
$\mathcal{Z}_{u,1,4}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\frac{C_{4}^{2}e^{-2ad}}{a}(V_{2}^{N},V_{2}^{N}*E)_{2},\penalty 10000\ \mathcal{Z}_{u,2,4}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{u,1,4}^{2}+e^{-4ad}C(d)C_{4}^{2}|{\Omega_{0}}|(V_{2}^{N},V_{2}^{N}*E)_{2}.$ (113)
If $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2},$ and $\mathcal{Z}_{u}$ are defined as in Lemma 4.7, then it follows that $\|\mathbb{B}D\mathbb{g}(\overline{u})^{N}(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})\|_{2}\leq\max\left\{1,\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\right\|_{2}\right\}\mathcal{Z}_{u}$.
Proof.
The proof follows from Lemma 6.5 of unbounded_domain_cadiot . In particular, each $\mathcal{Z}_{u,k,j}$ can be computed using the results of the aforementioned lemma. ∎
4.2 Constructive existence proofs of localized solutions
(a) Approximation of a “pulse” solution in the Glycolysis equation
(b) Approximation of a “pulse” solution in the Schnakenberg equation
(c) Approximation of a “pulse” solution in the Selkov-Schnakenberg equation
(d) Approximation of a “pulse” solution in the root-hair
equation
(e) Approximation of a “pulse” solution in the Brusselator equation
(f) Approximation of a “multi-pulse” solution in the Brusselator equation
Figure 2: Approximations of “pulse” solutions for general activator-inhibitor models
In this section, we apply the Newton-Kantorovich approach presented in Section 3 thanks to the explicit formulas derived in Section 4. In fact, the bounds are evaluated rigorously thanks to the Julia packages IntervalArithmetic.jl julia_interval and RadiiPolynomial.jl julia_olivier . More specifically, the computer-assisted proofs details can be found on julia_blanco_fassler . Now, we present our obtained constructive proofs of existence for localized patterns in (1).
Theorem 4.9 (Pulse in Glycolsis).
Let $\delta=0.45,c=a,b=0.44,j=0.5,a=0,h=0$. Moreover, let $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\times 10^{-7}$. Then there exists a unique solution $\tilde{\mathbf{u}}$ to (2) in $\overline{B_{r_{0}}(\mathbf{u}_{gyl})}\subset\mathcal{H}_{e}$ and we have that $\|\tilde{\mathbf{u}}-\mathbf{u}_{gyl}\|_{\mathcal{H}}\leq r_{0}$.
Proof.
Choose $N_{0}=1400,N=1100,d=30$. Then, we perform the full construction to build $\overline{\mathbf{u}}=\text{{$\gamma$}}^{\dagger}(\overline{\mathbf{U}})$. Then, we define $\mathbf{u}_{gly}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{\mathbf{u}}$. Next, we construct $B^{N}$, and use Lemma 2.2 to find
$\left\|\begin{bmatrix}B_{11}^{N}-B_{12}^{N}\\ B_{21}^{N}-B_{22}^{N}\end{bmatrix}\right\|_{2}\leq 14.36,\penalty 10000\ \kappa\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}14.2.$ (114)
This allows us to compute the $\mathcal{Z}_{2}(r)$ bound defined in Section 4. Finally, using julia_blanco_fassler , we choose $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\times 10^{-7}$ and define
$\mathcal{Y}_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}5.07\times 10^{-8}\text{,}\penalty 10000\ \mathcal{Z}_{2}(r_{0})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}952.07\text{,}\penalty 10000\ Z_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}0.142\text{,}\penalty 10000\ \mathcal{Z}_{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}3.911\times 10^{-5}$ (115)
and prove that these values satisfy Theorem 3.1. ∎
Theorem 4.10 (Pulse in Schnakenberg).
Let $\delta=0.011,c=1,b=3.6,j=0,a=2.8,h=0$. Moreover, let $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}3\times 10^{-10}$. Then there exists a unique solution $\tilde{\mathbf{u}}$ to (2) in $\overline{B_{r_{0}}(\mathbf{u}_{sch})}\subset\mathcal{H}_{e}$ and we have that $\|\tilde{\mathbf{u}}-\mathbf{u}_{sch}\|_{\mathcal{H}}\leq r_{0}$.
Proof.
Choose $N_{0}=1300,N=850,d=8$. The proof is obtained similarly to that of Theorem 4.9. In particular, we define
$\mathcal{Y}_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}3.54\times 10^{-11},\penalty 10000\ \mathcal{Z}_{2}(r_{0})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}8.71\times 10^{6},\penalty 10000\ Z_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}0.099,\penalty 10000\ \mathcal{Z}_{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}1.32\times 10^{-3}$ (116)
and prove that these values satisfy Theorem 3.1. ∎
Theorem 4.11 (Pulse in Brusselator).
Let $\delta=0.05,c=1.51,b=0,j=0,a=0.3,h=0.51$. Moreover, let $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}8\times 10^{-5}$. Then there exists a unique solution $\tilde{\mathbf{u}}$ to (2) in $\overline{B_{r_{0}}(\mathbf{u}_{bruss})}\subset\mathcal{H}_{e}$ and we have that $\|\tilde{\mathbf{u}}-\mathbf{u}_{bruss}\|_{\mathcal{H}}\leq r_{0}$.
Proof.
Choose $N_{0}=3200,N=1750,d=80$. The proof is obtained similarly to that of Theorem 4.9. In particular, we define
$\mathcal{Y}_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}1.62\times 10^{-8},\penalty 10000\ \mathcal{Z}_{2}(r_{0})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}1912.32,\penalty 10000\ Z_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}0.2621,\penalty 10000\ \mathcal{Z}_{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2.23\times 10^{-5}$ (117)
and prove that these values satisfy Theorem 3.1. ∎
Theorem 4.12 (Multi-Pulse in Brusselator).
Let $\delta=0.0095742405424,c=2.3299999237,b=0,j=0,a=16.421100616,h=1.3299999237$. Moreover, let $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}9\times 10^{-11}$. Then there exists a unique solution $\tilde{\mathbf{u}}$ to (2) in $\overline{B_{r_{0}}(\mathbf{u}_{bruss})}\subset\mathcal{H}_{e}$ and we have that $\|\tilde{\mathbf{u}}-\mathbf{u}_{bruss}\|_{\mathcal{H}}\leq r_{0}$.
Proof.
Choose $N_{0}=2000,N=1300,d=10$. The proof is obtained similarly to that of Theorem 4.9. In particular, we define
$\mathcal{Y}_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}7.51\times 10^{-12},\penalty 10000\ \mathcal{Z}_{2}(r_{0})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}5.44\times 10^{7},\penalty 10000\ Z_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}0.066,\penalty 10000\ \mathcal{Z}_{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}3.71\times 10^{-3}$ (118)
and prove that these values satisfy Theorem 3.1. ∎
Theorem 4.13 (Pulse in Selkov-Schnakenberg).
Let $\delta=0.01,c=1,b=0.589818,j=0.005,a=1.095,h=0$. Moreover, let $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}4\times 10^{-6}$. Then there exists a unique solution $\tilde{\mathbf{u}}$ to (2) in $\overline{B_{r_{0}}(\mathbf{u}_{sel})}\subset\mathcal{H}_{e}$ and we have that $\|\tilde{\mathbf{u}}-\mathbf{u}_{sel}\|_{\mathcal{H}}\leq r_{0}$.
Proof.
Choose $N_{0}=4000,N=3300,d=25$. The proof is obtained similarly to that of Theorem 4.9. In particular, we define
$\mathcal{Y}_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}1.32\times 10^{-6},\penalty 10000\ \mathcal{Z}_{2}(r_{0})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}40821.1,\penalty 10000\ Z_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}0.1603,\penalty 10000\ \mathcal{Z}_{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}8.66\times 10^{-4}$ (119)
and prove that these values satisfy Theorem 3.1. ∎
Theorem 4.14 (Pulse in Root-Hair).
Let $\delta=0.01,c=2,b=1.8,j=3.5,a=0,h=1$. Moreover, let $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}6\times 10^{-7}$. Then there exists a unique solution $\tilde{\mathbf{u}}$ to (2) in $\overline{B_{r_{0}}(\mathbf{u}_{rh})}\subset\mathcal{H}_{e}$ and we have that $\|\tilde{\mathbf{u}}-\mathbf{u}_{rh}\|_{\mathcal{H}}\leq r_{0}$.
Proof.
Choose $N_{0}=300,N=200,d=3$. The proof is obtained similarly to that of Theorem 4.9. In particular, we define
$\mathcal{Y}_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2.62\times 10^{-8},\penalty 10000\ \mathcal{Z}_{2}(r_{0})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}6417.5,\penalty 10000\ Z_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}0.1875,\penalty 10000\ \mathcal{Z}_{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}1.34\times 10^{-5}$ (120)
and prove that these values satisfy Theorem 3.1. ∎
5 Saddle-Node Bifurcations : zeros of $\mathbb{F}$
In this manuscript, we present a novel methodology for proving the existence of saddle-node bifurcations on unbounded domains. In fact, our approach relies on the setup described in Section 5 of unbounded_domain_cadiot , which deals with PDEs augmented with extra equations and extra parameters. In fact, in this section we focus on a computer-assisted methodology for proving constructively the existence of non-degenerate zeros of $\mathbb{F}$ given in (7).
Let $\bar{\nu}$ be fixed and be such that $\mathbb{l}_{\overline{\nu}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{l}(\overline{\nu})$ is invertible and define the $\mathcal{H}$-norm as
$\|\mathbf{u}\|_{\mathcal{H}}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\int_{\mathbb{R}^{2}}|\hat{\mathbf{u}}(\xi)||\mathscr{l}_{\overline{\nu}}(\xi)|d\xi$ (121)
where $\mathscr{l}_{\overline{\nu}}(\xi)$ is the Fourier transform of $\mathbb{l}_{\overline{\nu}}$. We now define the Hilbert spaces $H_{1}=\mathbb{R}\times(\mathcal{H}_{e})^{2}$ and $H_{2}=\mathbb{R}\times(L_{e}^{2})^{2}$ endowed with their norms
$\|(\nu,\mathbf{u},\mathbf{w})\|_{H_{1}}=\left(|\nu|^{2}+\|\mathbf{u}\|_{\mathcal{H}}^{2}+\|\mathbf{w}\|_{\mathcal{H}}^{2}\right)^{\frac{1}{2}}\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \|(\nu,\mathbf{u})\|_{H_{2}}=\left(|\nu|^{2}+\|\mathbf{u}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}\right)^{\frac{1}{2}}.$ (122)
We now define
$\mathbb{L}_{\nu}=\mathbb{L}(\nu)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}1&0&0\\ 0&\mathbb{l}(\nu)&0\\ 0&0&\mathbb{l}(\nu)\end{bmatrix}\penalty 10000\ \penalty 10000\ \mathbb{G}(\mathbf{u},\mathbf{w})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}0\\ \mathbb{g}(\mathbf{u})\\ D_{\mathbf{u}}\mathbb{g}(\mathbf{u})\mathbf{w}\end{bmatrix}$ (123)
and specify the definition of the Saddle-Node Map $\mathbb{F}:H_{1}\to H_{2}$ (7) as follows
$\mathbb{F}(\mathbf{x})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}(\mathbb{l}_{\overline{\nu}}\boldsymbol{\xi},\mathbf{w})_{2}-\alpha\\ \mathbb{f}(\nu,\mathbf{u})\\ D_{\mathbf{u}}\mathbb{f}(\nu,\mathbf{u})\mathbf{w}\end{bmatrix}$ (124)
where $\boldsymbol{\xi}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(\xi_{1},\xi_{2})\in\mathcal{H}_{e}$ is a non-trivial fixed function and $\alpha\in\mathbb{R}$ is a fixed constant.
As provided in Definition 1.1, non-degenerate zeros of $\mathbb{F}$ provide candidates for a saddle-node bifurcation. Moreover, unbounded_domain_cadiot provides that, given $\tilde{\mathbf{u}}\in\mathcal{H}_{e}$, $D\mathbb{f}(\tilde{\nu},\tilde{\mathbf{u}}):\mathbb{R}\times\mathcal{H}_{e}\to\mathbb{R}\times L^{2}_{e}$ is a Fredholm operator of index zero. The spectral condition of Definition 1.1 will then be verified in the next section. For now, we perform rigorous existence proofs of zeros of (7) in $H_{1}$ using the following theorem.
Theorem 5.1.
Let $\mathbb{A}:H_{2}\to H_{1}$ be a bounded linear operator. Moreover, let $\mathcal{Y}_{0},\mathcal{Z}_{1}$ be non-negative constants and let $\mathcal{Z}_{2}:(0,\infty)\to[0,\infty)$ be a non-negative function such that for all $r>0$
$\|\mathbb{A}\mathbb{F}(\overline{\mathbf{x}})\|_{H_{1}}\leq$ $\mathcal{Y}_{0}$ (125)
$\|I-\mathbb{A}D\mathbb{F}(\overline{\mathbf{x}})\|_{H_{1}}\leq$ $\mathcal{Z}_{1}$ (126)
$\|\mathbb{A}\left({D}\mathbb{F}(\overline{\mathbf{x}}+\mathbf{h})-D\mathbb{F}(\overline{\mathbf{x}})\right)\|_{H_{1}}\leq$ $\mathcal{Z}_{2}(r)r,\penalty 10000\ \penalty 10000\ \text{for all }\mathbf{h}\in B_{r}(0)$ (127)
If there exists $r>0$ such that (39) is satisfied, then there exists a unique $\tilde{\mathbf{x}}\in\overline{B_{r}(\overline{\mathbf{x}})}\subset H_{1}$ such that $\mathbb{F}(\tilde{\mathbf{x}})=0$, where $B_{r}(\overline{\mathbf{x}})$ is the open ball of radius $r$ in $H_{1}$ and centered at $\overline{\mathbf{x}}$.
Proof.
The same proof of Theorem 3.1 applies with the changing of the map and spaces. ∎
The augmented zero finding problem (7) has a periodic boundary value problem correspondence. In fact, as $\boldsymbol{\xi}\in L^{2}$, we can define $\boldsymbol{\Xi}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\text{{$\gamma$}}(\boldsymbol{\xi})\in\ell^{2}_{e}$. Then we define $X_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{R}\times(X^{l}_{e})^{2}$ and $X_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{R}\times(\ell^{2}_{e})^{2}$ the corresponding Hilbert spaces associated to their natural norms
$\|(\nu,\mathbf{U},\mathbf{W})\|_{X_{1}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(|\nu|^{2}+{|{\Omega_{0}}|}\|\mathbf{U}\|_{l}^{2}+|{\Omega_{0}}|\|\mathbf{W}\|_{\mathcal{H}}^{2}\right)^{\frac{1}{2}},\penalty 10000\ \|(\nu,\mathbf{U},\mathbf{W})\|_{X_{2}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(|\nu|^{2}+{|{\Omega_{0}}|}\|\mathbf{U}\|_{2}^{2}+|{\Omega_{0}}|\|\mathbf{W}\|_{2}^{2}\right)^{\frac{1}{2}}.$ (128)
Note that the normalization by $|{\Omega_{0}}|$} is useful to obtain an isometry (cf. Lemma 2.3). Next, we introduce $\mathbf{X}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(\nu,\text{{$\gamma$}}(\mathbf{u}),\text{{$\gamma$}}(\mathbf{w}))=(\nu,\mathbf{U},\mathbf{W})\in X_{1}$. Finally, we define a corresponding operator $F:X_{1}\to X_{2}$ along with $L$ and $G$ as
$F(\mathbf{X})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}|{\Omega_{0}}|((l_{\overline{\nu}}\boldsymbol{\Xi},\mathbf{W})_{2}-\alpha\\ f(\nu,\mathbf{U})\\ D_{\mathbf{U}}f(\nu,\mathbf{U})\mathbf{W}\end{bmatrix},\penalty 10000\ L=L(\nu)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}1&0&0\\ 0&l&0\\ 0&0&l\end{bmatrix},\penalty 10000\ \penalty 10000\ G(\mathbf{U},\mathbf{W})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}0\\ g(\mathbf{U})\\ D_{\mathbf{U}}g(\mathbf{U})\mathbf{W}\end{bmatrix}$ (129)
where $f(\nu,\mathbf{U})=f(\mathbf{U})$ is defined as in (35).
Note that the construction of $\overline{\mathbf{u}}$ can be done the same way as in Section 3.2. Furthermore, one can then construct $\overline{\mathbf{w}}$ by computing the approximate eigenvector corresponding to the 0 eigenvalue of $DF(\overline{U})$. Let us now discuss the approximate inverse, $\mathbb{A}$.
5.1 Approximate Inverse
We now briefly recall the setup of Section 5.2 of unbounded_domain_cadiot to build the approximate inverse. Denote $H_{2,{\Omega_{0}}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{(\nu,\mathbf{u},\mathbf{w})\in H_{2},\penalty 10000\ \penalty 10000\ \text{supp}(\mathbf{u})\subset\overline{{\Omega_{0}}},\text{supp}(\mathbf{w})\subset\overline{{\Omega_{0}}}\}$ and $H_{1,{\Omega_{0}}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{(\nu,\mathbf{u},\mathbf{w})\in H_{1},\penalty 10000\ \penalty 10000\ \text{supp}(\mathbf{u})\subset\overline{{\Omega_{0}}},\text{supp}(\mathbf{w})\subset\overline{{\Omega_{0}}}\}$. By construction, we have $\overline{\mathbf{x}}=(\overline{\nu},\overline{\mathbf{u}},\overline{\mathbf{w}})\in H_{1,{\Omega_{0}}}$. Recall that $\mathcal{B}(Y)$ is the set of bounded linear operators on some space $Y$. Denote by $\mathcal{B}_{\Omega_{0}}(H_{2})$ the following subset of $\mathcal{B}(H_{2})$
$\mathcal{B}_{\Omega_{0}}(H_{2})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{\mathbb{K}\in\mathcal{B}(H_{2}),\penalty 10000\ \penalty 10000\ \mathbb{K}=\begin{bmatrix}1&0&0\\ 0&\mathbb{1}_{\Omega_{0}}&0\\ 0&0&\mathbb{1}_{\Omega_{0}}\end{bmatrix}\mathbb{K}\begin{bmatrix}1&0&0\\ 0&\mathbb{1}_{\Omega_{0}}&0\\ 0&0&\mathbb{1}_{\Omega_{0}}\end{bmatrix}\right\}$.
Now, let $\overline{\text{{$\gamma$}}}:H_{2}\to X_{2}$, $\overline{\text{{$\gamma$}}}^{\dagger}:X_{2}\to H_{2}$$\overline{\text{{$\Gamma$}}}:\mathcal{B}(H_{2})\to\mathcal{B}(X_{2})$, and $\overline{\text{{$\Gamma$}}}^{\dagger}:\mathcal{B}(X_{2})\to\mathcal{B}(H_{2})$ be defined as follows
$\overline{\text{{$\gamma$}}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}1&0&0\\ 0&\text{{$\gamma$}}&0\\ 0&0&\text{{$\gamma$}}\end{bmatrix},\penalty 10000\ \overline{\text{{$\gamma$}}}^{\dagger}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}1&0&0\\ 0&\text{{$\gamma$}}^{\dagger}&0\\ 0&0&\text{{$\gamma$}}^{\dagger}\end{bmatrix},\penalty 10000\ \overline{\text{{$\Gamma$}}}\left(\mathbb{K}\right)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{\text{{$\gamma$}}}\mathbb{K}\overline{\text{{$\gamma$}}}^{\dagger},\penalty 10000\ \text{ and }\penalty 10000\ \overline{\text{{$\Gamma$}}}^{\dagger}\left(K\right)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{\text{{$\gamma$}}}^{\dagger}K\overline{\text{{$\gamma$}}}$ (130)
for all $\mathbb{K}\in\mathcal{B}(H_{2})$ and all $K\in\mathcal{B}(X_{2}).$ We now recall the equivalent to Lemma 2.3 for this case.
Lemma 5.2.
The map $\overline{\text{{$\gamma$}}}:H_{2,{\Omega_{0}}}\to X_{2}$ (respectively $\overline{\text{{$\Gamma$}}}:\mathcal{B}_{\Omega_{0}}(H_{2})\to\mathcal{B}(X_{2})$) is an isometric isomorphism whose inverse is $\overline{\text{{$\gamma$}}}^{\dagger}:X_{2}\to H_{2,{\Omega_{0}}}$ (respectively $\overline{\text{{$\Gamma$}}}^{\dagger}:\mathcal{B}(X_{2})\to\mathcal{B}_{\Omega_{0}}(H_{2})$). In particular,
$\|(\nu,\mathbf{u},\mathbf{w})\|_{H_{2}}=\|(\nu,\mathbf{U},\mathbf{W})\|_{X_{2}}\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ \|\mathbb{K}\|_{H_{2}}=\|K\|_{X_{2}}$
for all $(\nu,\mathbf{u},\mathbf{w})\in H_{2,{\Omega_{0}}}$ and all $\mathbb{K}\in\mathcal{B}_{\Omega_{0}}(H_{2})$, where $(\nu,\mathbf{U},\mathbf{W})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{\text{{$\gamma$}}}\left(\nu,\mathbf{u},\mathbf{w}\right)$ and $K\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{\text{{$\Gamma$}}}\left(\mathbb{K}\right)$.
Finally, let $N\in\mathbb{N}$ (NN is the size of the numerical approximation) and define $\overline{\text{{$\Pi$}}}^{N}$ and $\overline{\text{{$\Pi$}}}_{N}$ as
$\overline{\text{{$\Pi$}}}^{\leq N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}1&0&0\\ 0&\text{{$\Pi$}}^{\leq N}&0\\ 0&0&\text{{$\Pi$}}^{\leq N}\end{bmatrix}\penalty 10000\ \penalty 10000\ \text{ and }\overline{\text{{$\Pi$}}}^{>N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}0&0&0\\ 0&\text{{$\Pi$}}^{>N}&0\\ 0&0&\text{{$\Pi$}}^{>N}\end{bmatrix}$, (131)
where $\text{{$\Pi$}}^{\leq N}$ and $\text{{$\Pi$}}^{>N}$ are given in (3.2).
As before, we want to approximate the inverse of $D\mathbb{F}(\overline{\mathbf{x}})$ by some operator $\mathbb{A}:H_{2}\to H_{1}$ through the construction of an operator $\mathbb{B}:H_{2}\to H_{2}$, where $\mathbb{A}=\mathbb{L}_{\overline{\nu}}^{-1}\mathbb{B}.$ As described in Section 5 of unbounded_domain_cadiot , we compute a numerical approximate inverse $B^{N}:X_{2}\to X_{2}$ for $\overline{\text{{$\Pi$}}}^{\leq N}DF(\overline{\mathbf{X}})L_{\overline{\nu}}^{-1}\overline{\text{{$\Pi$}}}^{\leq N}$ such that $B^{N}=\overline{\text{{$\Pi$}}}^{\leq N}B^{N}\overline{\text{{$\Pi$}}}^{\leq N}$. Then, we define $\mathbb{A}:H_{2}\to H_{1}$ and $A:X_{2}\to X_{1}$ as
$\mathbb{A}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{L}_{\overline{\nu}}^{-1}\left(\begin{bmatrix}0&0&0\\ 0&\mathbb{1}_{\mathbb{R}\setminus\Omega_{0}}&0\\ 0&0&\mathbb{1}_{\mathbb{R}\setminus\Omega_{0}}\end{bmatrix}+\overline{\text{{$\Gamma$}}}^{\dagger}\left(\overline{\text{{$\Pi$}}}^{>N}+B^{N}\right)\right)\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ A\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}L_{\overline{\nu}}^{-1}\left(\overline{\text{{$\Pi$}}}^{>N}+B^{N}\right).$ (132)
With $\mathbb{A}$ now constructed, we are able to apply Theorem 3.1 to (7). What remains is to compute the needed bounds, $\mathcal{Y}_{0},\mathcal{Z}_{1},$ and $\mathcal{Z}_{2}$ in this case. This will be the focus of the next section.
5.2 Bounds for Saddle Nodes
In this section, we compute the bounds of Theorem 5.1 in order to prove existence and uniqueness of zeros to (7). As a result, we obtain a rigorous enclosure of a saddle node bifurcation to (2). We will begin with the $\mathcal{Y}_{0}$ bound. This can be done by combining the approach in Section 6 of unbounded_domain_cadiot to treat a scalar PDE with an extra equation and that of Section 4.1 of gs_cadiot_blanco where a system of PDEs was considered.
Lemma 5.3.
Let $\mathcal{Y}_{0}>0$ be such that
$\mathcal{Y}_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|^{\frac{1}{2}}\sqrt{\left\|B^{N}\begin{bmatrix}((l_{\overline{\nu}}\boldsymbol{\Xi},\mathbf{W}_{0})_{2}-\alpha\\ f(\overline{\nu},\overline{\mathbf{U}})\\ D_{\mathbf{U}}f(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}\end{bmatrix}\right\|_{2}^{2}+\left\|\begin{bmatrix}0\\ (\text{{$\Pi$}}^{\leq N_{0}}-\text{{$\Pi$}}^{\leq N})l_{\overline{\nu}}\overline{\mathbf{U}}+(\text{{$\Pi$}}^{3N_{0}}-\text{{$\Pi$}}^{\leq N})g(\overline{\mathbf{U}})\\ (\text{{$\Pi$}}^{\leq N_{0}}-\text{{$\Pi$}}^{\leq N})l_{\overline{\nu}}\overline{\mathbf{W}}+(\text{{$\Pi$}}^{3N_{0}}-\text{{$\Pi$}}^{\leq N})D_{\mathbf{U}}g(\overline{\mathbf{U}})\overline{\mathbf{W}}\end{bmatrix}\right\|_{2}^{2}}.$ (133)
Then $\|\mathbb{A}\mathbb{F}(\overline{\mathbf{x}})\|_{H_{1}}\leq\mathcal{Y}_{0}.$
Proof.
Following similar to steps to those from Lemma 4.2, we have
$\|\mathbb{A}\mathbb{F}(\overline{\mathbf{x}})\|_{H_{1}}=\|\mathbb{B}\mathbb{F}(\overline{\mathbf{x}})\|_{H_{2}}=\|BF(\overline{\mathbf{X}})\|_{X_{2}}$ $\leq\left(\left\|B^{N}F(\overline{\mathbf{X}})\right\|_{X_{2}}^{2}+\left\|\begin{bmatrix}0\\ \text{{$\Pi$}}^{>N}f(\overline{\nu},\overline{\mathbf{U}})\\ \text{{$\Pi$}}^{>N}D_{\mathbf{U}}f(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}\end{bmatrix}\right\|_{X_{2}}^{2}\right)^{\frac{1}{2}}.$ (134)
Following the same steps as Lemma 4.2 but on the larger system with more components will yield the result. ∎
Next, we compute the $\mathcal{Z}_{2}$ bound. Due to the presence of the extra equation, the computation of $\mathcal{Z}_{2}$ changes significantly. We summarize the results in the following lemma.
Lemma 5.4.
Let
$q_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\overline{u}_{2}+2\nu_{4}\mathbb{1}_{{\Omega_{0}}},\penalty 10000\ q_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\overline{u}_{1}+\nu_{5}\mathbb{1}_{\Omega_{0}},\penalty 10000\ q_{3}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\overline{w}_{2},\penalty 10000\ q_{4}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\overline{w}_{1}$ (135)
$\mathbb{B}_{1\to 5,24}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}0&\mathbb{B}_{12}&0&\mathbb{B}_{14}&0\\ 0&\mathbb{B}_{22}&0&\mathbb{B}_{24}&0\\ 0&\mathbb{B}_{32}&0&\mathbb{B}_{34}&0\\ 0&\mathbb{B}_{42}&0&\mathbb{B}_{44}&0\\ 0&\mathbb{B}_{52}&0&\mathbb{B}_{54}&0\end{bmatrix},\penalty 10000\ B_{1\to 5,24}^{N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{\text{{$\Pi$}}}^{\leq N}\overline{\text{{$\Gamma$}}}^{\dagger}(\mathbb{B}_{1\to 5,24})$ (136)
$M_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}B_{12}^{N}-B_{13}^{N}&B_{14}^{N}-B_{15}^{N}\\ B_{22}^{N}-B_{23}^{N}&B_{24}^{N}-B_{25}^{N}\\ B_{32}^{N}-B_{33}^{N}&B_{34}^{N}-B_{35}^{N}\\ B_{42}^{N}-B_{43}^{N}&B_{44}^{N}-B_{45}\\ B_{52}^{N}-B_{53}^{N}&B_{54}^{N}-B_{55}^{N}\end{bmatrix},\penalty 10000\ M_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}\mathbb{Q}_{1}^{2}+\mathbb{Q}_{2}^{2}&\mathbb{Q}_{1}\mathbb{Q}_{3}+\mathbb{Q}_{2}\mathbb{Q}_{4}\\ \mathbb{Q}_{3}\mathbb{Q}_{1}+\mathbb{Q}_{4}\mathbb{Q}_{2}&\mathbb{Q}_{3}^{2}+\mathbb{Q}_{4}^{2}+\mathbb{Q}_{1}^{2}+\mathbb{Q}_{2}^{2}\end{bmatrix}$. (137)
Additionally, we introduce the non-negative function $\mathcal{Z}_{2,2}:[0,\infty)\to(0,\infty)$ for and the non-negative constants $\mathcal{Z}_{2,j}>0$ for $j=1,3,4,5$ as
$\mathcal{Z}_{2,1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\max(1,\|B^{N}_{1\to 5,24}\|_{X_{2}})\left\|\begin{bmatrix}-|2\pi\xi|^{2}&0\\ 0&0\end{bmatrix}\mathscr{l}^{-1}(\xi)\right\|_{\mathcal{M}_{1}},\penalty 10000\ \mathcal{Z}_{2,2}(r)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\|M_{1}\|_{2,X_{2}}2\sqrt{10}\kappa\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}r$ (138)
$\mathcal{Z}_{2,3}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\|\overline{\text{{$\Pi$}}}^{\leq N}M_{1}M_{2}M_{1}^{*}\overline{\text{{$\Pi$}}}^{\leq N}\|_{X_{2}},\penalty 10000\ \mathcal{Z}_{2,4}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sqrt{\|\overline{\text{{$\Pi$}}}^{\leq N}M_{1}M_{2}\text{{$\Pi$}}^{>N}M_{2}^{*}M_{1}^{*}\overline{\text{{$\Pi$}}}^{\leq N}\|_{X_{2}}},$ (139)
$\mathcal{Z}_{2,5}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\varphi(\|Q_{1}^{2}+Q_{2}^{2}\|_{1},\|Q_{1}Q_{3}+Q_{2}Q_{4}\|_{1},\|Q_{3}Q_{1}+Q_{4}Q_{2}\|_{1},\|Q_{3}^{2}+Q_{4}^{2}+Q_{1}^{2}+Q_{2}^{2}\|_{1}).$ (140)
Finally, we let $\mathcal{Z}_{2}(r):[0,\infty)\to(0,\infty)$ be a non-negative function defined as
$\mathcal{Z}_{2}(r)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(1+\sqrt{3})\mathcal{Z}_{2,1}+\mathcal{Z}_{2,2}(r)+5\kappa\sqrt{\varphi(\mathcal{Z}_{2,3},\mathcal{Z}_{2,4},\mathcal{Z}_{2,4},\mathcal{Z}_{2,5})}.$ (141)
Then, it follows that $\|\mathbb{A}(D\mathbb{F}(\mathbf{h}+\overline{\mathbf{x}})-D\mathbb{F}(\overline{\mathbf{x}}))\|_{H_{1}}\leq\mathcal{Z}_{2}(r)r$
Proof.
The proof can be found in Appendix A. ∎
Next, we examine the $\mathcal{Z}_{1}$ bound for (7). To begin, we introduce some new notations. First, recall the definitions of $v_{1},v_{2}$ from (64). Then, we introduce
$v_{3}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}q_{1}w_{1}+q_{2}w_{2},\penalty 10000\ \text{and}\penalty 10000\ v_{4}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}q_{2}w_{1}$ (142)
and define
$\mathbb{m}(\overline{\mathbf{x}})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}D_{\mathbf{u}}\mathbb{g}(\overline{\mathbf{u}})&0\\ D^{2}_{\mathbf{u}}\mathbb{g}(\overline{\mathbf{u}})\overline{\mathbf{w}}&D_{\mathbf{u}}\mathbb{g}(\overline{\mathbf{u}})\end{bmatrix}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}\mathbb{v}_{1}&\mathbb{v}_{2}&0&0\\ -\mathbb{v}_{1}&-\mathbb{v}_{2}&0&0\\ \mathbb{v}_{3}&\mathbb{v}_{4}&\mathbb{v}_{1}&\mathbb{v}_{2}\\ -\mathbb{v}_{3}&-\mathbb{v}_{4}&-\mathbb{v}_{1}&-\mathbb{v}_{2}\end{bmatrix}$, $\mathbb{m}^{N}(\overline{\mathbf{x}})=\begin{bmatrix}\mathbb{v}_{1}^{N}&\mathbb{v}_{2}^{N}&0&0\\ -\mathbb{v}_{1}^{N}&-\mathbb{v}_{2}^{N}&0&0\\ \mathbb{v}_{3}^{N}&\mathbb{v}_{4}^{N}&\mathbb{v}_{1}^{N}&\mathbb{v}_{2}^{N}\\ -\mathbb{v}_{3}^{N}&-\mathbb{v}_{4}^{N}&-\mathbb{v}_{1}^{N}&-\mathbb{v}_{2}^{N}\end{bmatrix}$ (143)
where $\mathbb{v}_{j}$ is the multiplication operator associated to $v_{j}$ for $j=1,2,3,4$ and
$V_{j}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma(v_{j}),V_{j}^{N}=\Pi^{\leq N}V_{j},v_{j}^{N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma^{\dagger}(V_{j}^{N}).$ (144)
For $k=1,2$, we also define
$\Xi_{k}^{N}=\Pi^{\leq N}\Xi_{k},\xi_{k}^{N}=\gamma^{\dagger}(\Xi_{k}^{N})\penalty 10000\ \penalty 10000\ U_{1}^{N}=\Pi^{\leq N}U_{1},u_{1}^{N}=\gamma^{\dagger}(U_{1}^{N})\penalty 10000\ \penalty 10000\ W_{1}^{N}=\Pi^{\leq N}W_{1},w_{1}^{N}=\gamma^{\dagger}(W_{1}^{N}).$ (145)
Now, given $u\in L^{2}(\mathbb{R}^{2})$ and $U\in\ell^{2}(\mathbb{Z}^{2})$, define $u^{*}$ and $U^{*}$ as the duals in $L^{2}(\mathbb{R}^{2})$ and $\ell^{2}(\mathbb{Z}^{2})$ respectively of $u$ and $U$. More specifically, we have
$u^{*}(v)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(u,v)_{2},\penalty 10000\ \penalty 10000\ U^{*}(V)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(U,V)_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{ for all }v\in L^{2}(\mathbb{R}^{2})\text{ and }V\in\ell^{2}(\mathbb{Z}^{2}).$ (146)
Moreover, we slightly abuse notation and, given $u\in L^{2}$ and $U\in\ell^{2}$, we denote $u^{*}\mathbb{l}:\mathcal{H}\to\mathbb{R}$ and $U^{*}l:\mathscr{h}\to\mathbb{R}$ the following linear operators
$u^{*}\mathbb{l}v\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(u,\mathbb{l}v)_{2},\penalty 10000\ \penalty 10000\ U^{*}lV\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(U,lV)_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{ for all }v\in L^{2}(\mathbb{R}^{2})\text{ and }V\in\ell^{2}(\mathbb{Z}^{2}).$ (147)
Using the above notation, we introduce
$\mathbb{M}(\overline{\mathbf{x}})$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}-1&0&\boldsymbol{\xi}^{*}\mathbb{l}_{\overline{\nu}}^{*}\\ D_{\nu}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})&D_{\mathbf{u}}\mathbb{g}(\overline{\mathbf{u}})&0\\ D_{\nu,\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}}&D^{2}_{\mathbf{u}}\mathbb{g}(\overline{\mathbf{u}})\overline{\mathbf{w}}&D_{\mathbf{u}}\mathbb{g}(\overline{\mathbf{u}})\end{bmatrix}$, (148)
$\mathbb{M}^{N}(\overline{\mathbf{x}})$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}-1&0&(\boldsymbol{\xi}^{N})^{*}\mathbb{l}_{\overline{\nu}}^{*}\\ D_{\nu}\mathbb{f}^{N}(\overline{\nu},\overline{\mathbf{u}})&D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})&0\\ D_{\nu,\mathbf{u}}\mathbb{f}^{N}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}}&D^{2}_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})\overline{\mathbf{w}}&D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})\end{bmatrix},M^{N}(\overline{\mathbf{X}})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{\text{{$\Gamma$}}}^{\dagger}(\mathbb{M}^{N}(\overline{\mathbf{x}}))$ (149)
We further introduce notation for clarify
$\Delta U\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(\Delta U)_{n\in\mathbb{Z}}=(-\tilde{n}^{2}u_{n})_{n\in\mathbb{Z}}.$ (150)
Now, we state the following lemma for the $\mathcal{Z}_{1}$ bound.
Lemma 5.5.
Let $\mathcal{Z}_{\infty,1},\mathcal{Z}_{\infty,2},\mathcal{Z}_{\infty,3},\mathcal{Z}_{\infty},\mathcal{Z}_{u},Z_{1},$ and $\mathcal{Z}_{1}>0$ be constants satisfying
$\mathcal{Z}_{\infty,1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\sqrt{\|\Xi_{1}-\Xi_{1}^{N}\|_{2}^{2}+\|\Xi_{2}-\Xi_{2}^{N}\|_{2}^{2}},\penalty 10000\ \mathcal{Z}_{\infty,2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\sqrt{\|\Delta(\overline{U}_{1}-\overline{U}_{1}^{N})\|_{2}^{2}+\|\Delta(\overline{W}_{1}-\overline{W}_{1}^{N})\|_{2}^{2}},$ (151)
$\mathcal{Z}_{\infty,3}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\|\mathbb{l}_{\overline{\nu}}^{-1}\|_{2}\varphi\left(\left(\sum_{j=1}^{2}\|V_{j}-V_{j}^{N}\|_{1}^{2}\right)^{\frac{1}{2}},0,\left(\sum_{j=3}^{4}\|V_{j}-V_{j}^{N}\|_{1}^{2}\right)^{\frac{1}{2}},\left(\sum_{j=1}^{2}\|V_{j}-V_{j}^{N}\|_{1}^{2}\right)^{\frac{1}{2}}\right),$ (152)
$\mathcal{Z}_{\infty}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\varphi(0,\mathcal{Z}_{\infty,1},\mathcal{Z}_{\infty,2},\mathcal{Z}_{\infty,3}),\penalty 10000\ \mathcal{Z}_{u}\geq\left\|\mathbb{B}\begin{bmatrix}0&0\\ 0&\mathbb{m}^{N}(\overline{\mathbf{x}})\end{bmatrix}\left(\mathbb{L}_{\overline{\nu}}^{-1}-\overline{\text{{$\Gamma$}}}^{\dagger}(L_{\overline{\nu}}^{-1})\right)\right\|_{H_{2}},$ (153)
$Z_{1}\geq\left\|I-B(I+M^{N}(\overline{\mathbf{X}})L_{\overline{\nu}}^{-1})\right\|_{X_{2}},\penalty 10000\ \mathcal{Z}_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}Z_{1}+\mathcal{Z}_{u}+\|M_{1}\|_{2,X_{2}}\mathcal{Z}_{\infty}.$ (154)
Then, it follows that $\|I-{\mathbb{A}}D\mathbb{F}(\overline{\mathbf{x}})\|_{H_{1}}\leq\mathcal{Z}_{1}.$
Proof.
Following similar steps to those performed in (4.4), we obtain
$\|I-\mathbb{A}D\mathbb{F}(\overline{\mathbf{x}})\|_{H_{1}}$ $\leq Z_{1}+\mathcal{Z}_{u}+\|M_{1}\|_{2,X_{2}}\|(\mathbb{M}(\overline{\mathbf{x}})-\mathbb{M}^{N}(\overline{\mathbf{x}}))\mathbb{L}_{\overline{\nu}}^{-1}\|_{H_{2}}.$ (155)
For the final term, we can use $\varphi$ defined in Lemma 4.1 to get
$\|(\mathbb{M}(\overline{\mathbf{x}})-\mathbb{M}^{N}(\overline{\mathbf{x}}))\mathbb{L}_{\overline{\nu}}^{-1}\|_{H_{2}}\leq\varphi\left(0,\left\|\boldsymbol{\xi}-\boldsymbol{\xi}^{N}\right\|_{2},\left\|\begin{bmatrix}D_{\nu}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})-D_{\nu}\mathbb{f}^{N}(\overline{\nu},\overline{\mathbf{u}})\\ D_{\nu,\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}}-D_{\nu,\mathbf{u}}\mathbb{f}^{N}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}}^{N}\end{bmatrix}\right\|_{2},\|\mathbb{l}_{\overline{\nu}}^{-1}\|_{2}\left\|\mathbb{m}(\overline{\mathbf{x}})-\mathbb{m}^{N}(\overline{\mathbf{x}})\right\|_{2}\right).$ (156)
Now, we will examine each term required to evaluate this expression. To begin, observe that
$\|\boldsymbol{\xi}-\boldsymbol{\xi}^{N}\|_{2}=|{\Omega_{0}}|\|\boldsymbol{\Xi}-\boldsymbol{\Xi}^{N}\|_{2}=|{\Omega_{0}}|\sqrt{\|\Xi_{1}-\Xi_{1}^{N}\|_{2}^{2}+\|\Xi_{2}-\Xi_{2}^{N}\|_{2}^{2}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{\infty,1}.$ (157)
This is the second term required. For the third term, notice we can write
$\left\|\begin{bmatrix}D_{\nu}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})-D_{\nu}\mathbb{f}^{N}(\overline{\nu},\overline{\mathbf{u}})\\ D_{\nu,\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}}-D_{\nu,\mathbf{u}}\mathbb{f}^{N}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}}^{N}\end{bmatrix}\right\|_{2}=|{\Omega_{0}}|\left\|\begin{bmatrix}D_{\nu}f(\overline{\nu},\overline{\mathbf{U}})-D_{\nu}f^{N}(\overline{\nu},\overline{\mathbf{U}})\\ D_{\nu,\mathbf{U}}f(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}-D_{\nu,\mathbf{U}}f^{N}(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}^{N}\end{bmatrix}\right\|_{2}$ (158)
$\hskip 42.67912pt=|{\Omega_{0}}|\sqrt{\|D_{\nu}f(\overline{\nu},\overline{\mathbf{U}})-D_{\nu}f^{N}(\overline{\nu},\overline{\mathbf{U}})\|_{2}^{2}+\|D_{\nu,\mathbf{U}}f(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}-D_{\nu,\mathbf{U}}f^{N}(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}^{N}\|_{2}^{2}}$ (159)
$\hskip 42.67912pt=|{\Omega_{0}}|\sqrt{\|\Delta(U_{1}-U_{1}^{N})\|_{2}^{2}+\|\Delta(W_{1}-W_{1}^{N})\|_{2}^{2}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{\infty,2}.$ (160)
Finally, we estimate the fourth term.
$\|\mathbb{l}_{\overline{\nu}}^{-1}\|_{2}\left\|\mathbb{m}(\overline{\mathbf{x}})-\mathbb{m}^{N}(\overline{\mathbf{x}})\right\|_{2}\leq\|\mathbb{l}_{\overline{\nu}}^{-1}\|_{2}\varphi\left(\left(\sum_{j=1}^{2}\|V_{j}\|_{1}^{2}\right)^{\frac{1}{2}},0,\left(\sum_{j=3}^{4}\|V_{j}\|_{1}^{2}\right)^{\frac{1}{2}},\left(\sum_{j=1}^{2}\|V_{j}\|_{1}^{2}\right)^{\frac{1}{2}}\right)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{\infty,3}.$ (161)
This concludes the proof. ∎
Next, we must compute the $Z_{1}$ and $\mathcal{Z}_{u}$ bounds. We begin with the $Z_{1}$ bound.
Lemma 5.6.
Let
$Z_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sqrt{Z_{0}^{2}+2\left(\sqrt{Z_{1,1}^{2}+Z_{1,2}^{2}}+\sqrt{Z_{1,3}^{2}+Z_{1,4}^{2}}\right)^{2}}$ (162)
where $Z_{1,1}$ and $Z_{1,2},\left(\frac{\mathscr{l}_{ij}}{l_{\mathrm{den}}}\right)_{2N}$ are defined as in Lemma 4.5 and
$Z_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\|\overline{\text{{$\Pi$}}}^{\leq 2N}-\overline{\text{{$\Pi$}}}^{\leq 3N}B(I_{d}+M^{N}(\overline{X})L^{-1})\overline{\text{{$\Pi$}}}^{\leq 2N}\|_{X_{2}}$
$Z_{1,3}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\frac{l_{22}}{l_{\mathrm{den}}}\right)_{2N}\|V_{3}^{N}\|_{1}+\left(\frac{l_{21}}{l_{\mathrm{den}}}\right)_{2N}\|V_{4}^{N}\|_{1},\penalty 10000\ \penalty 10000\ Z_{1,4}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\frac{l_{12}}{l_{\mathrm{den}}}\right)_{2N}\|V_{3}^{N}\|_{1}+\left(\frac{l_{11}}{l_{\mathrm{den}}}\right)_{2N}\|V_{4}^{N}\|_{1}.$
Then, it follows that $\|I_{d}-B(I_{d}+Mg^{N}(\overline{X})L_{\overline{\nu}}^{-1})\|_{X_{2}}\leq Z_{1}$.
Proof.
To begin, we introduce a truncation.
$\|I_{d}-B(I_{d}+M^{N}(\overline{X})L_{\overline{\nu}}^{-1})\|_{X_{2}}^{2}$ (163)
$\leq\|\overline{\text{{$\Pi$}}}^{>2N}-B(I_{d}+M^{N}(\overline{X})L_{\overline{\nu}}^{-1})\overline{\text{{$\Pi$}}}^{>2N}\|_{X_{2}}^{2}+\|\overline{\text{{$\Pi$}}}^{\leq 2N}-B(I_{d}+M^{N}(\overline{X})L_{\overline{\nu}}^{-1})\text{{$\Pi$}}^{\leq 2N}\|_{X_{2}}^{2}$ (164)
$=\|\overline{\text{{$\Pi$}}}^{>2N}-B\overline{\text{{$\Pi$}}}^{>2N}-BM^{N}(\overline{X})\overline{\text{{$\Pi$}}}^{<2N}L_{\overline{\nu}}^{-1}\|_{X_{2}}^{2}+Z_{0}^{2}$ (165)
$=\|\overline{\text{{$\Pi$}}}^{>N}M^{N}(\overline{X})L_{\overline{\nu}}^{-1}\overline{\text{{$\Pi$}}}^{<2N}\|_{X_{2}}^{2}+Z_{0}^{2}$ (166)
$=\left\|\begin{bmatrix}\text{{$\Pi$}}^{>N}&0\\ 0&\text{{$\Pi$}}^{>N}\end{bmatrix}\begin{bmatrix}Dg^{N}(\overline{\nu},\overline{\mathbf{U}})&0\\ D^{2}g^{N}(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}&Dg^{N}(\overline{\nu},\overline{\mathbf{U}})\end{bmatrix}\begin{bmatrix}l_{\overline{\nu}}^{-1}&0\\ 0&l_{\overline{\nu}}^{-1}\end{bmatrix}\begin{bmatrix}\text{{$\Pi$}}^{>2N}&0\\ 0&\text{{$\Pi$}}^{>2N}\end{bmatrix}\right\|_{2}^{2}+Z_{0}^{2}.$ (167)
Now, we use the structure of $\begin{bmatrix}Dg^{N}(\overline{\nu},\overline{\mathbf{U}})&0\\ D^{2}g^{N}(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}&Dg^{N}(\overline{\nu},\overline{\mathbf{U}})\end{bmatrix}$ to get
$\left\|\begin{bmatrix}\text{{$\Pi$}}^{>N}&0\\ 0&\text{{$\Pi$}}^{>N}\end{bmatrix}\begin{bmatrix}Dg^{N}(\overline{\nu},\overline{\mathbf{U}})&0\\ D^{2}g^{N}(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}&Dg^{N}(\overline{\nu},\overline{\mathbf{U}})\end{bmatrix}\begin{bmatrix}l_{\overline{\nu}}^{-1}&0\\ 0&l_{\overline{\nu}}^{-1}\end{bmatrix}\begin{bmatrix}\text{{$\Pi$}}^{>2N}&0\\ 0&\text{{$\Pi$}}^{>2N}\end{bmatrix}\right\|_{2}$ (168)
$=\left\|\begin{bmatrix}\text{{$\Pi$}}^{>N}Dg^{N}(\overline{\nu},\overline{\mathbf{U}})l_{\overline{\nu}}^{-1}\text{{$\Pi$}}^{>2N}&0\\ \text{{$\Pi$}}^{>N}D^{2}g^{N}(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}l_{\overline{\nu}}^{-1}\text{{$\Pi$}}^{>2N}&\text{{$\Pi$}}^{>N}Dg^{N}(\overline{\nu},\overline{\mathbf{U}})l_{\overline{\nu}}^{-1}\text{{$\Pi$}}^{>2N}\end{bmatrix}\right\|_{2}$ (169)
$\leq\left\|\begin{bmatrix}\text{{$\Pi$}}^{>N}Dg^{N}(\overline{\nu},\overline{\mathbf{U}})l_{\overline{\nu}}^{-1}\text{{$\Pi$}}^{>2N}&0\\ 0&\text{{$\Pi$}}^{>N}Dg^{N}(\overline{\nu},\overline{\mathbf{U}})l_{\overline{\nu}}^{-1}\text{{$\Pi$}}^{>2N}\end{bmatrix}\right\|_{2}+\left\|\begin{bmatrix}0&0\\ \text{{$\Pi$}}^{>N}D^{2}g^{N}(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}l_{\overline{\nu}}^{-1}\text{{$\Pi$}}^{>2N}&0\end{bmatrix}\right\|_{2}$ (170)
$=\left\|\text{{$\Pi$}}^{>N}Dg^{N}(\overline{\nu},\overline{\mathbf{U}})l_{\overline{\nu}}^{-1}\text{{$\Pi$}}^{>2N}\right\|_{2}+\|\text{{$\Pi$}}^{>N}D^{2}g^{N}(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}l_{\overline{\nu}}^{-1}\text{{$\Pi$}}^{>2N}\|_{2}$ (171)
$\leq\sqrt{2Z_{1,1}^{2}+2Z_{1,2}^{2}}+\|\text{{$\Pi$}}^{>N}D^{2}g^{N}(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}l_{\overline{\nu}}^{-1}\text{{$\Pi$}}^{>2N}\|_{2}.$ (172)
The final term above can be estimated in a very similar way to $Z_{1,1}$ and $Z_{1,2}$. We obtain
$\|\text{{$\Pi$}}^{>N}D^{2}g^{N}(\overline{\nu},\overline{\mathbf{U}})\overline{\mathbf{W}}l_{\overline{\nu}}^{-1}\text{{$\Pi$}}^{>2N}\|_{2}\leq\sqrt{2Z_{1,3}^{2}+2Z_{1,4}^{2}}.$ (173)
Therefore, we have
$\|I_{d}-B(I_{d}+M^{N}(\overline{X})L_{\overline{\nu}}^{-1})\|_{2}\leq\sqrt{Z_{0}^{2}+2\left(\sqrt{Z_{1,1}^{2}+Z_{1,2}^{2}}+\sqrt{Z_{1,3}^{2}+Z_{1,4}^{2}}\right)^{2}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}Z_{1}$ (174)
as desired. ∎
Finally, we compute the bound $\mathcal{Z}_{u}$ for the saddle node map (7).
Lemma 5.7.
Let $\left(\mathcal{Z}_{u,k,j}\right)_{k\in\{1,2\},j\in\{1,2,3,4\}}$ be defined as Lemma 4.7. Let $\left(\mathcal{Z}_{u,k,j}\right)_{k\in\{3,4\},j\in\{1,2,3,4\}}$ bounds satisfying
$\mathcal{Z}_{u,3,1}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{22}\mathbb{l}_{den}^{-1}\mathbb{v}_{3}^{N}\|_{2},\penalty 10000\ \mathcal{Z}_{u,4,1}\geq\|\mathbb{1}_{{\Omega_{0}}}(\mathbb{l}_{22}\mathbb{l}_{den}^{-1}-\Gamma^{\dagger}(l_{22}l_{\mathrm{den}}^{-1}))\mathbb{v}_{3}^{N}\|_{2}$ (175)
$\mathcal{Z}_{u,3,2}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{21}\mathbb{l}_{den}^{-1}\mathbb{v}_{4}^{N}\|_{2},\penalty 10000\ \mathcal{Z}_{u,4,2}\geq\|\mathbb{1}_{{\Omega_{0}}}(\mathbb{l}_{21}\mathbb{l}_{den}^{-1}-\Gamma^{\dagger}(l_{21}l_{\mathrm{den}}^{-1}))\mathbb{v}_{4}^{N}\|_{2}$ (176)
$\mathcal{Z}_{u,3,3}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{12}\mathbb{l}_{den}^{-1}\mathbb{v}_{3}^{N}\|_{2},\penalty 10000\ \mathcal{Z}_{u,4,3}\geq\|\mathbb{1}_{{\Omega_{0}}}(\mathbb{l}_{12}\mathbb{l}_{den}^{-1}-\Gamma^{\dagger}(l_{12}l_{\mathrm{den}}^{-1}))\mathbb{v}_{3}^{N}\|_{2}$ (177)
$\mathcal{Z}_{u,3,4}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}\mathbb{l}_{22}\mathbb{l}_{den}^{-1}\mathbb{v}_{4}^{N}\|_{2},\penalty 10000\ \mathcal{Z}_{u,4,4}\geq\|\mathbb{1}_{{\Omega_{0}}}(\mathbb{l}_{11}\mathbb{l}_{den}^{-1}-\Gamma^{\dagger}(l_{11}l_{\mathrm{den}}^{-1}))\mathbb{v}_{4}^{N}\|_{2}$ (178)
Moreover, given $k\in\{1,2,3,4\}$, define $\mathcal{Z}_{u,k}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sqrt{2}\sqrt{(\mathcal{Z}_{u,k,1}+\mathcal{Z}_{u,k,2})^{2}+(\mathcal{Z}_{u,k,3}+\mathcal{Z}_{u,k,4})^{2}}$. Then, it follows that $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2},\mathcal{Z}_{u,3},$ and $\mathcal{Z}_{u,4}$ satisfy
$\mathcal{Z}_{u,1}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})\mathbb{l}^{-1}\|_{2},\penalty 10000\ \mathcal{Z}_{u,2}\geq\|\mathbb{1}_{{\Omega_{0}}}D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})\|_{2}$ (179)
$\mathcal{Z}_{u,3}\geq\|\mathbb{1}_{\mathbb{R}\setminus{\Omega_{0}}}D^{2}_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})\overline{\mathbf{w}}\mathbb{l}^{-1}\|_{2},\penalty 10000\ \mathcal{Z}_{u,4}\geq\|\mathbb{1}_{{\Omega_{0}}}D^{2}_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})\overline{\mathbf{w}}(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})\|_{2}.$ (180)
Finally, defining
$\mathcal{Z}_{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sqrt{\mathcal{Z}_{u,1}^{2}+\mathcal{Z}_{u,2}^{2}}+\sqrt{\mathcal{Z}_{u,3}^{2}+\mathcal{Z}_{u,4}^{2}},$ (181)
it follows that $\left\|\mathbb{B}\begin{bmatrix}0&0\\ 0&\mathbb{m}^{N}(\overline{\mathbf{x}})\end{bmatrix}\left(\overline{\text{{$\Gamma$}}}^{\dagger}(L^{-1})-\mathbb{L}^{-1}\right)\right\|_{X_{2}}\leq\|M_{1}\|_{2,X_{2}}\mathcal{Z}_{u}$.
Proof.
To begin, observe that
$\left\|\mathbb{B}\begin{bmatrix}0&0\\ 0&\mathbb{m}^{N}(\overline{\mathbf{x}})\end{bmatrix}\left(\overline{\text{{$\Gamma$}}}^{\dagger}(L^{-1})-\mathbb{L}^{-1}\right)\right\|_{2}$ (182)
$\leq\|M_{1}\|_{2,X_{2}}\left\|\begin{bmatrix}D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})&0\\ D^{2}_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})\overline{\mathbf{w}}&D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})\end{bmatrix}\begin{bmatrix}\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1}&0\\ 0&\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1}\end{bmatrix}\right\|_{2}$ (183)
$=\|M_{1}\|_{2,X_{2}}\left\|\begin{bmatrix}D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})&0\\ D^{2}_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})\overline{\mathbf{w}}(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})&D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})\end{bmatrix}\right\|_{2}$ (184)
$\leq\|M_{1}\|_{2,X_{2}}\biggl(\left\|\begin{bmatrix}D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})&0\\ 0&D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})\end{bmatrix}\right\|_{2}$ (185)
$\hskip 56.9055pt+\left\|\begin{bmatrix}0&0\\ D^{2}_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})\overline{\mathbf{w}}(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})&0\end{bmatrix}\right\|_{2}\biggr)$ (186)
$=\|M_{1}\|_{2,X_{2}}\left(\|D_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})\|_{2}+\|D^{2}_{\mathbf{u}}\mathbb{g}^{N}(\overline{\mathbf{u}})\overline{\mathbf{w}}(\text{{$\Gamma$}}^{\dagger}(l^{-1})-\mathbb{l}^{-1})\|_{2}\right)$ (187)
$\leq\|M_{1}\|_{2,X_{2}}\mathcal{Z}_{u}$ (188)
as desired. ∎
We now estimate $\mathcal{Z}_{u,3}$ and $\mathcal{Z}_{u,4}$.
Lemma 5.8.
Let $a,E,C(d),$ and $C_{j}$ for $j=1,2,3,4$ be defined as in Lemma 4.8. Then, let $(\mathcal{Z}_{u,k,j})_{k\in\{3,4\},j\in\{1,2,3,4\}}>0$ be defined as
$\mathcal{Z}_{u,3,1}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\frac{C_{1}^{2}e^{-2ad}}{a}(V_{3}^{N},V_{3}^{N}*E)_{2},\penalty 10000\ \mathcal{Z}_{u,4,1}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{u,3,1}^{2}+e^{-4ad}C(d)C_{1}^{2}|{\Omega_{0}}|(V_{3}^{N},V_{3}^{N}*E)_{2}$ (189)
$\mathcal{Z}_{u,3,2}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\frac{C_{2}^{2}e^{-2ad}}{a}(V_{4}^{N},V_{4}^{N}*E)_{2},\penalty 10000\ \mathcal{Z}_{u,4,2}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{u,3,2}^{2}+e^{-4ad}C(d)C_{2}^{2}|{\Omega_{0}}|(V_{4}^{N},V_{4}^{N}*E)_{2}$ (190)
$\mathcal{Z}_{u,3,3}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\frac{C_{3}^{2}e^{-2ad}}{a}(V_{3}^{N},V_{3}^{N}*E)_{2},\penalty 10000\ \mathcal{Z}_{u,4,3}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{u,3,3}^{2}+e^{-4ad}C(d)C_{3}^{2}|{\Omega_{0}}|(V_{3}^{N},V_{3}^{N}*E)_{2}$ (191)
$\mathcal{Z}_{u,3,4}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\frac{C_{4}^{2}e^{-2ad}}{a}(V_{4}^{N},V_{4}^{N}*E)_{2},\penalty 10000\ \mathcal{Z}_{u,4,4}^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{u,3,4}^{2}+e^{-4ad}C(d)C_{4}^{2}|{\Omega_{0}}|(V_{4}^{N},V_{4}^{N}*E)_{2}.$ (192)
If $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2},\mathcal{Z}_{u,3},\mathcal{Z}_{u,4},$ and $\mathcal{Z}_{u}$ are defined as in Lemma 5.7, then it follows that
$\left\|\mathbb{B}\begin{bmatrix}0&0\\ 0&\mathbb{m}^{N}(\overline{\mathbf{x}})\end{bmatrix}\left(\overline{\text{{$\Gamma$}}}^{\dagger}(L^{-1})-\mathbb{L}^{-1}\right)\right\|_{X_{2}}\leq\|M_{1}\|_{2,X_{2}}\mathcal{Z}_{u}$.
Proof.
The proof follows from Lemma 6.5 of unbounded_domain_cadiot . In particular, each $\mathcal{Z}_{u,k,j}$ can be computed using the results of the aforementioned lemma. ∎
5.3 Proofs of Saddle Nodes
In this section, we present the computer assisted proofs of Saddle Node bifurcations in (2).
Figure 3: Approximation of a saddle-node bifurcation “pulse” solution for the Glycolysis equation.
Theorem 5.9.
[…]
Proof.
Choose $N_{0}=800,N=620,d=40$. Then, we perform the full construction to build $\overline{\mathbf{x}}=\overline{\text{{$\gamma$}}}^{\dagger}(\nu,\overline{\mathbf{U}},\overline{\mathbf{W}})$. Then, we define $\mathbf{x}_{glys}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\overline{\mathbf{x}}$. Next, we construct $B^{N}$ and use Lemma 2.2 to compute $\kappa$. This allows us to compute the $\mathcal{Z}_{2}(r)$ bound defined in Section 5.2. Finally, using julia_blanco_fassler , we choose $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}4\times 10^{-8}$ and define
$\mathcal{Y}_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}3.31\times 10^{-9}\text{,}\penalty 10000\ \mathcal{Z}_{2}(r_{0})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2.51\times 10^{4}\text{,}\penalty 10000\ Z_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}0.112\text{,}\penalty 10000\ \mathcal{Z}_{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}1.56\times 10^{-4}.$ (193)
and prove that these values satisfy Theorem 5.1. ∎
[…]
$),\displaystyle\sigma(D\mathbb{f}(\tilde{\mathbf{u}}))=\sigma_{ess}(D\mathbb{f}(\tilde{\mathbf{u}}))\cup Eig(D\mathbb{f}(\tilde{\mathbf{u}}$)), (194)
[…]
$\sigma_{ess}(D\mathbb{f}(\tilde{\mathbf{u}}))=\sigma_{ess}(\mathbb{l})=\{\lambda\in\sigma(\mathscr{l}(\xi)),\penalty 10000\ \text{ for }\xi\in\mathbb{R}\}.$ (195)
In other words, $)\sigma_{ess}(D\mathbb{f}(\tilde{\mathbf{u}}$)) corresponds to the union of spectrum of the symbol $\mathscr{l}(\xi)$, for all $\xi\in\mathbb{R}$. The next lemma provides an explicit enclosure of such a set
Lemma 6.1.
Assume that $\nu_{1}+\nu_{2}\geq 0$ and that $(\nu_{1}+\nu_{2})^{2}+4\nu_{2}\nu_{3}\geq 0$ . Moreover, let $\lambda_{sup}$ be defined as
$\lambda_{sup}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sup_{\xi\in\mathbb{R}}\frac{\mathscr{l}_{11}(\xi)+\mathscr{l}_{22}(\xi)+\sqrt{(\mathscr{l}_{11}(\xi)-\mathscr{l}_{22}(\xi))^{2}+4\nu_{2}\nu_{3}}}{2}.$ (196)
Then, we have that $\lambda_{sup}<0$ and that $\sigma_{ess}(D\mathbb{f}(\tilde{\mathbf{u}}))=(-\infty,\lambda_{sup}]$.
Proof.
The proof relies on the fact that, given $\xi\in\mathbb{R}$, the eigenvalues of $\mathscr{l}(\xi)$ are given by
$\lambda_{\pm}(\xi)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\mathscr{l}_{11}(\xi)+\mathscr{l}_{22}(\xi)\pm\sqrt{(\mathscr{l}_{11}(\xi)-\mathscr{l}_{22}(\xi))^{2}+4\nu_{2}\nu_{3}}}{2}.$ (197)
In particular, assuming that $\nu_{1}+\nu_{2}\geq 0$ and that $(\nu_{1}+\nu_{2})^{2}+4\nu_{2}\nu_{3}\geq 0$, we obtain that $(\mathscr{l}_{11}(\xi)-\mathscr{l}_{22}(\xi))^{2}+4\nu_{2}\nu_{3}\geq 0$ for all $\xi\in\mathbb{R}$. Moreover, an alternative form for the eigenvalues is $\lambda_{\pm}(\xi)=\frac{\mathscr{l}_{11}(\xi)+\mathscr{l}_{22}(\xi)\pm\sqrt{(\mathscr{l}_{11}(\xi)+\mathscr{l}_{22}(\xi))^{2}-4\mathscr{l}_{den}(\xi)}}{2}$. But using that $\nu_{2}\geq-\nu_{1}$, we obtain that
$\mathscr{l}_{11}(\xi)+\mathscr{l}_{22}(\xi)\leq\nu_{1}-\nu_{2}<0$
since $\nu_{2}>0$ by definition. This implies that $\lambda_{sup}$ must be strictly negative. ∎
[…]
6.1 Control of the eigenvalues
In this section, given $\delta>0$, we define $\sigma_{\delta}$ be the following set
$\sigma_{\delta}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{\lambda\in\mathbb{C},\penalty 10000\ |\lambda|\leq\delta,\penalty 10000\ Re(\lambda)\geq 0\}.$ (198)
[…]
Lemma 6.2.
There exists $C>0$ such that
$\|(\mathscr{l}-\lambda)^{-1}\|_{\mathcal{M}_{1}}\leq\frac{C}{\sqrt{\lambda_{sup}^{2}+|\lambda|^{2}}}$
for all $\lambda\in\mathbb{C}$ such that $Re(\lambda)\geq 0$.
Proof.
First, recall that $\mathscr{l}(\xi)$ possesses two negative eigenvalues $\lambda_{\pm}(\xi)$, given in (197). Given $\lambda$ with $Re(\lambda)\geq 0$, this implies that $\mathscr{l}-\lambda$ is invertible, and we have
$(\mathscr{l}-\lambda)^{-1}=\frac{1}{(\lambda-\lambda_{+})(\lambda-\lambda_{-})}\begin{bmatrix}\mathscr{l}_{22}-\lambda&-\mathscr{l}_{12}\\ -\mathscr{l}_{21}&\mathscr{l}_{11}-\lambda\end{bmatrix}$.
In particular, since $Re(\lambda)\geq 0$ and $\lambda_{\pm}$ are real valued, one easily notice that there exists $C>0$ such that
$\left\|\frac{1}{(\lambda-\lambda_{-})}\begin{bmatrix}\mathscr{l}_{22}-\lambda&-\mathscr{l}_{12}\\ -\mathscr{l}_{21}&\mathscr{l}_{11}-\lambda\end{bmatrix}\right\|_{\mathcal{M}_{1}}\leq C$ (199)
for all $\lambda\in\mathbb{C}$ such that $Re(\lambda)\geq 0$. This concludes the proof. ∎
The above lemma allows us to derive an upper bound for the value of $\delta_{0}$.
Lemma 6.3.
Let $r_{0}$ be given in Theorem 5.9, and let $)\lambda\in Eig(D\mathbb{f}(\tilde{\mathbf{u}}$)) with $Re(\lambda)\geq 0$, then
$|\lambda|\leq C\left(\sqrt{2(\|\bar{V}_{1}\|_{1}^{2}+\|\bar{V}_{2}\|_{1}^{2})}+\sqrt{5}\kappa\mathcal{Z}_{2,3}r_{0}+3\kappa\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}r_{0}^{2}\right)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\delta_{0}$ (200)
where $C$ is given in Lemma 6.2.
Proof.
Let $\mathbf{u}\in L^{2}_{e}$ be an associated eigenfunction to $\lambda$ such that $\|\mathbf{u}\|_{2}=1$. In fact, we have that
$\mathbb{l}\mathbf{u}-\lambda\mathbf{u}+D\mathbb{g}(\tilde{\mathbf{u}})\mathbf{u}=0.$
In fact, this implies that
$\frac{1}{\|(\mathbb{L}-\lambda I)^{-1}\|_{2}}\leq\|(\mathbb{L}-\lambda I)\mathbf{u}\|_{2}\leq\|D\mathbb{g}(\tilde{\mathbf{u}})\mathbf{u}\|_{2}\leq\|D\mathbb{g}(\tilde{\mathbf{u}})\|_{2}$
since $\|\mathbf{u}\|_{2}=1.$ But using Lemma 6.2, we obtain that
$\frac{|\lambda|}{C}\leq\|D\mathbb{g}(\tilde{\mathbf{u}})\|_{2}.$
Now, using the proof of Lemma 4.3, we notice that
$\|D\mathbb{g}(\tilde{\mathbf{u}})\|_{2}$ $\leq\|D\mathbb{g}(\bar{\mathbf{u}})\|_{2}+\|D\mathbb{g}(\tilde{\mathbf{u}})-D\mathbb{g}(\bar{\mathbf{u}})\|_{2}$
$\leq\sqrt{2(\|\bar{V}_{1}\|_{1}^{2}+\|\bar{V}_{2}\|_{1}^{2})}+\sqrt{5}\kappa\mathcal{Z}_{2,3}r_{0}+3\kappa\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}r_{0}^{2}.$
This concludes the proof. ∎
[…]
Now in order to obtain a precise control on the set $)\sigma_{\delta_{0}}\cap\sigma(D\mathbb{f}(\tilde{\mathbf{u}}$)), we use the spectrum of $Df(\bar{\mathbf{U}})$, as presented in cadiot2025stabilityanalysislocalizedsolutions . For this purpose, we first compute a pseudo-diagonalization of $Df(\bar{\mathbf{U}})$, that is $P$ and $P^{-1}$ such that
$\mathcal{D}=P^{-1}Df(\bar{\mathbf{U}})P$ (201)
is close to being diagonal. In fact, we make the following choice for the construction of $P$
$P=P^{N}+\Pi^{>N}$ (202)
where $P^{N}=\Pi^{\leq N}P^{N}\Pi^{\leq N}$, is essentially a square matrix of size $N+1$. Then, assuming that $P^{N}:\Pi^{\leq N}\ell^{2}\to\Pi^{\leq N}\ell^{2}$ is invertible, we denote $(P^{N})^{-1}:\Pi^{\leq N}\ell^{2}\to\Pi^{\leq N}\ell^{2}$ its inverse and get that $P^{-1}$ is given as follows
$P^{-1}=(P^{N})^{-1}+\Pi^{>N}.$ (203)
Then, using similar notations as in cadiot2025stabilityanalysislocalizedsolutions , we define $S$ as the diagonal matrix having the same diagonal as $\mathcal{D}$, that is
$(SU)_{n}=\lambda_{n}U_{n},\text{ where }\lambda_{n}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{D}_{n,n}\text{ for all }n\in\mathbb{N}_{0}.$
Moreover, we define $R$ as
$R\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{D}-S.$
[…]
Lemma 6.4.
Let $z_{1},z_{2}$ be defined as in Lemma 4.6 and let $\mu\in\sigma_{\delta_{0}}$. Then, define $\mathscr{l}_{den,\mu}$ as
$\mathscr{l}_{den,\mu}(\xi)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(-\nu(2\pi\xi)^{2}+\nu_{1}-\mu)(-(2\pi\xi)^{2}-\nu_{2}-\mu)-\nu_{2}\nu_{3}\text{ for all }\xi\in\mathbb{R}.$ (204)
In fact, $\mathscr{l}_{den,\mu}(\xi)$ is the determinant of $\mathscr{l}(\xi)-\mu I$. Then, we have that
$\left|\mathcal{F}^{-1}\left(\frac{d_{1}|2\pi\xi|^{2}+d_{2}}{\mathscr{l}_{den,\mu}(\xi)}\right)(x)\right|\leq C_{0}e^{-a|x|},$
for all $\mu\in\sigma_{\delta_{0}}$, with $a$ and $C_{0}$ given as in (84).
Proof.
The proof follows from the proof of Lemma 4.6. In particular, we prove that there exists $a(\mu)$ and $C_{0}(\mu)$ such that
$\left|\mathcal{F}^{-1}\left(\frac{d_{1}|2\pi\xi|^{2}+d_{2}}{\mathscr{l}_{den,\mu}(\xi)}\right)(x)\right|\leq C_{0}(\mu)e^{-a(\mu)|x|}.$
Then, following tedious but straightforward computations, we prove that $a(\mu)$ is minimized at 0 and that $C_{0}(\mu)$ is maximized at 0, concluding the proof. ∎
[…]
Lemma 6.5.
Let $\delta_{0}$ be defined as in (200) and let $\mathcal{J}=\sigma_{\delta_{0}}$. Furthermore, let $t\in\mathbb{C}$ be such that $S+tI$ is invertible. Moreover, let $r_{0}$ and $\tilde{\mathbf{u}}$ be given as in Theorem 5.9. Now, let us introduce various constants satisfying the following
$\mathcal{Z}_{u,1}$ $\geq\sup_{\mu\in\mathcal{J}}\left\|\mathbb{1}_{\mathbb{R}\setminus\Omega_{0}}\left(\mathbb{l}-\mu I\right)^{-1}D\mathbb{g}(\bar{\mathbf{u}})\right\|_{2}$
$\mathcal{Z}_{u,2}$ $\geq\sup_{\mu\in\mathcal{J}}\|\mathbb{1}_{\Omega_{0}}\left(\Gamma^{\dagger}\left((l-\mu I)^{-1}\right)-(\mathbb{l}-\mu I)^{-1}\right)D\mathbb{g}(\bar{\mathbf{u}})\|_{2}$
$\mathcal{Z}_{u,3}$ $\geq\sup_{\mu\in\mathcal{J}}\|\pi^{N}(S+tI)^{-1}P^{-1}(l-\mu I)\|_{2}\|\mathbb{1}_{\Omega_{0}}\left(\Gamma^{\dagger}\left((l-\mu I)^{-1}\right)-(\mathbb{l}-\mu I)^{-1}\right)D\mathbb{g}(\bar{\mathbf{u}})\|_{2}$
$\mathcal{C}_{1}$ $\geq\sup_{\mu\in\mathcal{J}}\frac{1}{r_{0}}\left\|\left(\mathbb{l}-\mu I\right)^{-1}\left(D\mathbb{g}(\bar{\mathbf{u}})-D\mathbb{g}(\tilde{\mathbf{u}})\right)\right\|_{2}$
$\mathcal{C}_{2}$ $\geq\sup_{\mu\in\mathcal{J}}\frac{1}{r_{0}}\|\pi^{N}(S+tI)^{-1}P^{-1}(l-\mu I)\|_{2}\left\|\mathbb{1}_{\Omega_{0}}\left(\mathbb{l}-\mu I\right)^{-1}\left(D\mathbb{g}(\bar{\mathbf{u}})-D\mathbb{g}(\tilde{\mathbf{u}})\right)\right\|_{2}$
$Z_{1,1}$ $\geq\sup_{\mu\in\mathcal{J}}\|\pi_{N}(l-\mu I)^{-1}R\pi^{N}\|_{2},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ Z_{1,2}\geq\sup_{\mu\in\mathcal{J}}\|\pi_{N}(l-\mu I)^{-1}R\pi_{N}\|_{2}$
$Z_{1,3}$ $\geq\|\pi^{N}(S+tI)^{-1}R\pi^{N}\|_{2},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ Z_{1,4}\geq\|\pi^{N}(S+tI)^{-1}R\pi_{N}\|_{2}.$
If
$\mathcal{C}_{1}r_{0}<1,$ (205)
we define $\beta_{1}$ as $\beta_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\mathcal{Z}_{u,1}+\mathcal{C}_{1}r_{0}}{1-\mathcal{C}_{1}r_{0}}.$ If in addition
$1-Z_{1,2}-\mathcal{Z}_{u,2}-(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0}>0,$ (206)
then we define
$\beta_{2}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{Z_{1,1}+(\mathcal{Z}_{u,2}+(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0})\|P^{N}\|_{2}}{1-Z_{1,2}-\mathcal{Z}_{u,2}-(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0}}$ (207)
$\epsilon_{n}^{(\infty)}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|\lambda_{n}+t|\left(Z_{1,3}+Z_{1,4}\beta_{2}+\left(\mathcal{Z}_{u,3}+\mathcal{C}_{2}r_{0}(1+\beta_{1}^{2})^{\frac{1}{2}}\right)\left(\|P^{N}\|_{2}+\beta_{2}\right)\right)$
$\epsilon_{n}^{(q)}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|\lambda_{n}+t|\left(Z_{1,3}+Z_{1,4}\frac{Z_{1,1}+2\mathcal{Z}_{u,2}\|P^{N}\|_{2}}{1-Z_{1,2}-2\mathcal{Z}_{u,2}}+2\mathcal{Z}_{u,3}\left(\|P^{N}\|_{2}+\frac{Z_{1,1}+2\mathcal{Z}_{u,2}\|P^{N}\|_{2}}{1-Z_{1,2}-2\mathcal{Z}_{u,2}}\right)\right)$
$\epsilon_{n}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\max\left\{\epsilon_{n}^{(\infty)},\epsilon_{n}^{(q)}\right\}$
for all $n\in\mathbb{N}_{0}$.
[…]
Remark 6.1.
The bounds $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2},\mathcal{Z}_{u,3}$ introduced in the previous lemma are not the same as the ones introduced in Lemma 4.7. However, they are closely related and can be computed in a very similar fashion. Indeed, their computation is based on the proof of Lemma 4.7, combined with the estimates of Lemma 6.4. Note that the additional factor $2$ in the inequalities for $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2},\mathcal{Z}_{u,3}$ allows to provide upper bounds for $\mathcal{Z}_{u,1}^{(q)},\mathcal{Z}_{u,2}^{(q)},\mathcal{Z}_{u,3}^{(q)}$ given in Lemma 4.2 of cadiot2025stabilityanalysislocalizedsolutions (see the proof of Theorem 5.2 in the aforementioned paper). Details for the computation of the bounds of Lemma 6.5 are given in cadiot2025stabilityanalysislocalizedsolutions , and are implemented explicitly in the code julia_blanco_fassler . In fact, such computations heavily relies on the analysis presented in Section 4, and we prefer not to present them in details to avoid intensive overlapping.
6.2 Existence proof of a saddle-node
Using the above analysis, we are able to rigorously enclose the spectrum $D\mathbb{f}(\tilde{\nu},\tilde{\mathbf{u}}), fo$r a given zero $(\tilde{\nu},\tilde{\mathbf{u}})\in H_{1}$ of $\mathbb{f}$. This is particularly useful for verifying that the zeros of (7) are indeed saddle-nodes, following Definition 1.1. In fact, we obtain the following result.
Theorem 6.6.
[…]
Proof.
[…]
6.3 Stability analysis
In this section, we focus on the stability of some localized stationary solutions of Section 3. In particular, given $\tilde{\mathbf{u}}$ proven in Section 3, that is there exists $r_{0}$ such that $\|\tilde{\mathbf{u}}-\bar{\mathbf{u}}\|_{\mathcal{H}}\leq r_{0}$, then we use Lemma 6.5 to control the spectrum of $D\mathbb{f}(\tilde{\mathbb{u}}):L^{2}_{e}\to L^{2}_{e}$. In fact, this allows us to conclude about the (nonlinear) stability of $\tilde{\mathbf{u}}$ with respect to even perturbations in $L^{2}_{e}$. In the cases for which $\tilde{\mathbf{u}}$ is unstable, we count the number of unstable directions. More specifically, we obtain the following results for which computational details are given in julia_blanco_fassler .
Theorem 6.7.
The solution $\tilde{u}\in\mathcal{H}_{e}$ obtained in Theorem 4.10 is stable. In particular, the elements of the spectrum of $D\mathbb{f}(\tilde{\mathbb{u}}):L^{2}_{e}\to L^{2}_{e}$ possess a strictly negative real part.
Proof.
We prove that all the disks constructed thanks to Lemma 6.5 are strictly contained on the left-part of the complex plane. ∎
Theorem 6.8.
The solution $\tilde{u}\in\mathcal{H}_{e}$ obtained in Theorem 4.14 is unstable. In particular, $D\mathbb{f}(\tilde{\mathbb{u}}):L^{2}_{e}\to L^{2}_{e}$ possesses exactly one positive eigenvalue.
Proof.
We construct the disks of Lemma 6.5 and prove that one of them, say $\overline{B_{\epsilon_{0}}(\lambda_{0})}$ is on strictly included on the right-part of the complex plane. More specifically, we get $\epsilon_{0}=1.8\times 10^{-4}$ and $\lambda_{0}=0.18$. Then, we prove that $\overline{B_{\epsilon_{0}}(\lambda_{0})}$ is disjoint from the rest of the disks. Finally, we conclude the proof by proving the rest of the disks is strictly contained on the left-part of the complex plane. ∎
7 Conclusion
In this manuscript, we further investigated the results of the authors of AlSaadi_unifying_framework . More specifically, the authors of the aforementioned work computed solutions to (2) on long 1D domains along with various bifurcations. The authors asked if their results could be rigorously justified, and we provided the justification for some of their results. Indeed, we provided proofs of existence and local uniqueness of some of their approximate solutions on the unbounded domain $\mathbb{R}$. Moreover, we provide the general framework needed to attempt to prove any of the approximate localized patterns in AlSaadi_unifying_framework . Following this, we proved the existence of a saddle-node bifurcation on the unbounded domain $\mathbb{R}$. This not only provides an answer to one type of bifurcation studied in AlSaadi_unifying_framework , but also provides the framework needed to prove saddle-node bifurcations on unbounded domains using the theory of unbounded_domain_cadiot . Our results have provided partial answers to some of the questions asked by the authors of AlSaadi_unifying_framework as well as an additional application of the method developed in unbounded_domain_cadiot to saddle-node bifurcations.
While we have answered some of the questions asked by the authors of AlSaadi_unifying_framework , there are many more still present. In particular, the authors computed various branches of solutions to (2) by performing numerical continuation. Some of these branches lead to saddle-node bifurcations, which could now be proven using our approach described in Section 5; however, this does not allow for the proof of the branch itself. The necessary theory to obtain proofs of branches of solutions on unbounded domains has already been developed and applied in whitham_cadiot . In particular, the approach of Section 4.6 in whitham_cadiot allowed for one to obtain the proof of the branch in the Whitham equation. This approach relies on the method demonstrated in breden_polynomial_chaos and marschal . The approach would need to be modified slightly as whitham_cadiot used parameter continuation. This approach works well assuming no bifurcations are encountered, which is not the case here as we have saddle-node bifurcations. Combining the approach used in marschal along with the method for unbounded domains used in whitham_cadiot would allow for a rigorous proof of the branch. We consider this a potential future work.
Another question asked by the authors of AlSaadi_unifying_framework was regarding Hopf bifurcations. In particular, the authors describe the location of various approximate Hopf bifurcations in many of the models that fit into (2). Furthermore, they ask about whether or not these Hopf bifurcations are sub/super critical. The approach to prove Hopf bifurcations on unbounded domains would be similar to that for saddle-node bifrucations except with a more complicated augmented system. This work is in-progress.
To further the last point, associated to each Hopf bifurcation is a branch of time-periodic solutions. This branch was not computed by the authors of AlSaadi_unifying_framework . We consider it of interest to investigate such a branch of solutions. Furthermore, could we obtain a rigorous proof of a solution along this branch? In the framework of computer assisted proofs, this would be similar to proving a solution that is localized in its spatial dimensions but periodic in the time dimension. Proving such a solution has not been considered using the approach of unbounded_domain_cadiot . This work is also in-progress.
To conclude, while our work here has provided a partial answer to some of the questions asked by the authors of AlSaadi_unifying_framework , we believe many more of their results could be rigorously verified using similar methods to those described in this manuscript. We believe our methods could be of use to those studying models of similar type to (2). For instance, proofs for the 2D Gray-Scott model were already provided in gs_cadiot_blanco . We believe this manuscript opens many possibilities for rigorous investigation of these models.
8 Acknowledgments
The authors wish to thank Professor Nicolas Verschueren van Rees for his insights and fruitful discussions.
Appendix A Computing the Bound $\mathcal{Z}_{2}$ for Saddle Nodes
We prove Lemma 5.4. In particular, we compute the $\mathcal{Z}_{2}$ bound.
Proof.
Let $\mathbf{h}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(\eta,\mathbf{h}_{1},\mathbf{h}_{2})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(\eta,h_{1,1},h_{1,2},h_{2,1},h_{2,2})\in B_{r}(\overline{\mathbf{x}})$. Now, observe that
$\|\mathbb{A}(D\mathbb{F}(\mathbf{h}+\overline{\mathbf{x}})-D\mathbb{F}(\overline{\mathbf{x}}))\|_{H_{1}}$ (A.1)
$\leq\left\|\mathbb{B}\begin{bmatrix}0&0&0\\ D_{\nu}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})-D_{\nu}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})&0&0\\ D_{\nu,\mathbf{u}}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})(\mathbf{h}_{2}+\overline{\mathbf{w}})-D_{\nu,\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}})&0&0\end{bmatrix}\right\|_{H_{1},H_{2}}$ (A.2)
$\hskip 14.22636pt+\left\|\mathbb{B}\begin{bmatrix}0&0&0\\ 0&D_{\mathbf{u}}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})-D_{\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})&0\\ 0&D^{2}_{\mathbf{u}}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})(\mathbf{h}_{2}+\overline{\mathbf{w}})-D^{2}_{\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}}&D_{\mathbf{u}}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})-D_{\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\end{bmatrix}\right\|_{H_{1},H_{2}}$ (A.3)
$\leq\left\|\mathbb{B}_{1\to 5,24}\right\|_{H_{2}}\left(\left\|\begin{bmatrix}D_{\nu}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})-D_{\nu}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\\ D_{\nu,\mathbf{u}}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})(\mathbf{h}_{2}+\overline{\mathbf{w}})-D_{\nu,\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}})\end{bmatrix}\right\|_{2}\right)$ (A.4)
$\hskip 56.9055pt+\left\|\mathbb{B}\left(\begin{bmatrix}0&0&0&0&0\\ 0&\eta\Delta&0&0&0\\ 0&0&0&0&0\\ 0&\eta\Delta&0&\eta\Delta&0\\ 0&0&0&0&0\end{bmatrix}+D\mathbb{G}(\mathbf{h}_{1}+\overline{\mathbf{u}},\mathbf{h}_{2}+\overline{\mathbf{w}})-D\mathbb{G}(\overline{\mathbf{u}},\overline{\mathbf{w}})\right)\right\|_{H_{1},H_{2}}$ (A.5)
$\leq\left\|\mathbb{B}_{1\to 5,2\to 5}\right\|_{H_{2}}\left(\left\|\begin{bmatrix}D_{\nu}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})-D_{\nu}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\\ D_{\nu,\mathbf{u}}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})(\mathbf{h}_{2}+\overline{\mathbf{w}})-D_{\nu,\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}})\end{bmatrix}\right\|_{2}\right)$ (A.6)
$\hskip 56.9055pt+\sqrt{3}\|\mathbb{B}_{1\to 5,24}\|_{H_{2}}|\eta|\left\|\begin{bmatrix}\Delta&0\\ 0&0\end{bmatrix}\right\|_{\mathcal{H},2}+\left\|\mathbb{B}\left(D\mathbb{G}(\mathbf{h}_{1}+\overline{\mathbf{u}},\mathbf{h}_{2}+\overline{\mathbf{w}})-D\mathbb{G}(\overline{\mathbf{u}},\overline{\mathbf{w}})\right)\right\|_{H_{1},H_{2}}$ (A.7)
where $\mathbb{B}_{1\to 5,24}$ is defined in (135) and we used the facts that
$D_{\mathbf{u}}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})-D_{\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})=\begin{bmatrix}\eta\Delta&0\\ 0&0\end{bmatrix}+D_{\mathbf{u}}\mathbb{g}(\mathbf{h}_{1}+\overline{\mathbf{u}})-D_{\mathbf{u}}\mathbb{g}(\overline{\mathbf{u}})$ (A.8)
$D^{2}_{\mathbf{u}}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})(\mathbf{h}_{2}+\overline{\mathbf{w}})-D^{2}_{\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}}=\begin{bmatrix}\eta\Delta&0\\ 0&0\end{bmatrix}+D^{2}_{\mathbf{u}}\mathbb{g}(\mathbf{h}_{1}+\overline{\mathbf{u}})(\mathbf{h}_{2}+\overline{\mathbf{w}})-D^{2}_{\mathbf{u}}\mathbb{g}(\overline{\mathbf{u}})\overline{\mathbf{w}}.$ (A.9)
For the first term in (A.7), we use the Fourier transform to get
$\left\|\begin{bmatrix}D_{\nu}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})-D_{\nu}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\\ D_{\nu,\mathbf{u}}\mathbb{f}(\eta+\overline{\nu},\mathbf{h}_{1}+\overline{\mathbf{u}})(\mathbf{h}_{2}+\overline{\mathbf{w}})-D_{\nu,\mathbf{u}}\mathbb{f}(\overline{\nu},\overline{\mathbf{u}})\overline{\mathbf{w}})\end{bmatrix}\right\|_{2}=\left\|\begin{bmatrix}\Delta h_{1,1}\\ 0\\ \Delta h_{2,1}\\ 0\end{bmatrix}\right\|_{2}$ $=\left\|\begin{bmatrix}\Delta&0&0&0\\ 0&0&0&0\\ 0&0&\Delta&0\\ 0&0&0&0\end{bmatrix}\begin{bmatrix}h_{1,1}\\ h_{1,2}\\ h_{2,1}\\ h_{2,2}\end{bmatrix}\right\|_{2}$ (A.10)
$\hskip-56.9055pt=\left\|\begin{bmatrix}-|2\pi\xi|^{2}&0&0&0\\ 0&0&0&0\\ 0&0&-|2\pi\xi|^{2}&0\\ 0&0&0&0\end{bmatrix}\begin{bmatrix}\hat{h}_{1,1}\\ \hat{h}_{1,2}\\ \hat{h}_{2,1}\\ \hat{h}_{2,2}\end{bmatrix}\right\|_{2}$ (A.11)
$\hskip-56.9055pt\leq\left\|\begin{bmatrix}-|2\pi\xi|^{2}&0\\ 0&0\end{bmatrix}\mathscr{l}^{-1}(\xi)\right\|_{\mathcal{M}_{1}}\left\|\mathscr{l}(\xi)\hat{\mathbf{h}}\right\|_{2}$ (A.12)
$\hskip-56.9055pt\leq\left\|\begin{bmatrix}-|2\pi\xi|^{2}&0\\ 0&0\end{bmatrix}\mathscr{l}^{-1}(\xi)\right\|_{\mathcal{M}_{1}}r\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\mathcal{Z}_{2,1}r}{\|\mathbb{B}_{1\to 5,24}\|_{H_{2}}}.$ (A.13)
The quantity $\left\|\begin{bmatrix}-|2\pi\xi|^{2}&0\\ 0&0\end{bmatrix}\mathscr{l}^{-1}(\xi)\right\|_{\mathcal{M}_{1}}$ can be computed using standard Calculus techniques. Next, we examine the second term in (A.7).
$\sqrt{3}\|\mathbb{B}_{1\to 5,24}\|_{H_{2}}|\eta|\left\|\begin{bmatrix}\Delta&0\\ 0&0\end{bmatrix}\right\|_{\mathcal{H},2}\leq\sqrt{3}\|\mathbb{B}_{1\to 5,24}\|_{H_{2}}\left\|\begin{bmatrix}\Delta&0\\ 0&0\end{bmatrix}\right\|_{\mathcal{H},2}r.$ (A.14)
Now, let $\mathbf{z}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(z_{1},z_{2})\in\mathcal{H}$ such that $\|\mathbf{z}\|_{\mathcal{H}}=1$. Then,
$\left\|\begin{bmatrix}\Delta&0\\ 0&0\end{bmatrix}\right\|_{\mathcal{H},2}$ $=\left\|\begin{bmatrix}\Delta&0\\ 0&0\end{bmatrix}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}\right\|_{2}\leq\left\|\begin{bmatrix}-|2\pi\xi|^{2}&0\\ 0&0\end{bmatrix}\mathscr{l}^{-1}(\xi)\right\|_{\mathcal{M}_{1}}=\frac{\mathcal{Z}_{2,1}}{\|\mathbb{B}_{1\to 5,24}\|_{H_{2}}}.$ (A.15)
Finally, we examine the third term in (A.7). To do so, we first introduce some notation. Let
$\mathbb{G}_{3}(\mathbf{x})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}0\\ u_{1}^{2}u_{2}\\ -u_{1}^{2}u_{2}\\ u_{1}(2u_{2}w_{1}+u_{1}w_{2})\\ -u_{1}(2u_{2}w_{1}+u_{1}w_{2})\end{bmatrix}$. (A.16)
Then, observe that we can write
$\mathbb{G}_{3}(\mathbf{u},\mathbf{w})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{G}_{3,1}\mathbf{x}(\mathbb{G}_{3,2}\mathbf{x}\mathbb{G}_{3,3}\mathbf{x}+\mathbb{G}_{3,4}\mathbf{x}\mathbb{G}_{3,5}\mathbf{x})$ (A.17)
where
$\mathbb{G}_{3,1}=\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}0&0&0&0&0\\ 0&1&0&0&0\\ 0&-1&0&0&0\\ 0&1&0&0&0\\ 0&-1&0&0&0\end{bmatrix},\penalty 10000\ \mathbb{G}_{3,2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}0&0&0&0&0\\ 0&1&0&0&0\\ 0&1&0&0&0\\ 0&0&2&0&0\\ 0&0&2&0&0\end{bmatrix},\penalty 10000\ \mathbb{G}_{3,3}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}0&0&0&0&0\\ 0&0&1&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&1&0\end{bmatrix}$ (A.18)
$\mathbb{G}_{3,4}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&1&0&0&0\\ 0&1&0&0&0\end{bmatrix},\mathbb{G}_{3,5}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&1\\ 0&0&0&0&1\end{bmatrix}$. (A.19)
Note that the dependency on $\mathbf{x}$ is not needed since the above term does not depend on $\nu$; however, we write it for convenience. Now, let $\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{y}_{3}\in H_{1}$. Then, we write $G_{3}:(H_{1})^{3}\to(\mathcal{H}_{e})^{2}$ as
$G_{3}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{y}_{3})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{G}_{3,1}\mathbf{y}_{1}(\mathbb{G}_{3,2}\mathbf{y}_{2}\mathbb{G}_{3,3}\mathbf{y}_{3}+\mathbb{G}_{3,4}\mathbf{y}_{2}\mathbb{G}_{3,5}\mathbf{y}_{3})$ (A.20)
for all $\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{y}_{3}\in H_{1}$. Now, let $\mathbf{z}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(\alpha,\mathbf{z}_{1},\mathbf{z}_{2})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(\alpha,z_{1,1},z_{1,2},z_{2,1},z_{2,2})\in H_{1},\|\mathbf{z}\|_{H_{1}}=1$. Similar to the proof of Lemma 4.3, observe that
$(D\mathbb{G}(\mathbf{h}_{1}+\overline{\mathbf{u}},\mathbf{h}_{2}+\overline{\mathbf{w}})-D\mathbb{G}(\overline{\mathbf{u}},\overline{\mathbf{w}}))\mathbf{z}=D^{2}\mathbb{G}(\overline{\mathbf{u}},\overline{\mathbf{w}})(\mathbf{h},\mathbf{z})+G_{3}(\mathbf{z},\mathbf{h},\mathbf{h})+G_{3}(\mathbf{h},\mathbf{z},\mathbf{h})+G_{3}(\mathbf{h},\mathbf{h},\mathbf{z}).$ (A.21)
Generally, by definition of $G_{3}$, we have $\|\mathbb{B}G_{3}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{y}_{3})\|_{H_{2}}\leq\left\|M_{1}\right\|_{2,X_{2}}\|G_{3}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{y}_{3})\|_{H_{2}}$ where $M_{1}$ is defined in (135). Now, we examine the specific cases to obtain
$\|G_{3}(\mathbf{z},\mathbf{h},\mathbf{h})\|_{H_{2}}=\sqrt{2}\left(\|z_{1,1}h_{1,1}h_{1,2}\|_{2}^{2}+\|z_{1,1}(2h_{1,2}h_{2,1}+h_{1,1}h_{2,2})\|_{2}^{2}\right)^{\frac{1}{2}}$ (A.22)
$\leq 2\sqrt{10}\kappa\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}\|\mathbf{z}\|_{H_{1}}\|\mathbf{h}\|_{H_{1}}^{2}\leq 2\sqrt{10}\kappa\|\mathscr{l}^{-1}\|_{\mathcal{M}_{2}}r^{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\mathcal{Z}_{2,2}(r)r}{\|M_{1}\|_{2,X_{2}}}.$ (A.23)
The other estimates proceed similarly, and have the same result. That is,
$\|G_{3}(\mathbf{h},\mathbf{z},\mathbf{h})\|_{H_{2}}\leq\frac{\mathcal{Z}_{2,2}(r)}{\|M_{1}\|_{2,X_{2}}},\penalty 10000\ \|G_{3}(\mathbf{h},\mathbf{h},\mathbf{z})\|_{H_{2}}\leq\frac{\mathcal{Z}_{2,2}(r)r}{\|M_{1}\|_{2,X_{2}}}$ (A.24)
where $\mathcal{Z}_{2,2}(r)$ is defined as in (139). Notice that due to our choice of estimate for $\|z_{i,j}h_{k,l}h_{m,n}\|_{2},$ we obtain the same result for each quantity. This significantly simplifies our computation compared to the approach used in gs_cadiot_blanco . It also allow us to use the same bound for the saddle node bounds as we did for patterns. This is one of the reasons we chose this approach in this work.
It now remains to study $\left\|\mathbb{B}D^{2}\mathbb{G}(\overline{\mathbf{u}},\overline{\mathbf{w}})(\mathbf{h},\mathbf{z})\right\|_{H_{2}}$. Observe first the following
$\|\mathbb{B}D^{2}\mathbb{G}(\overline{\mathbf{u}},\overline{\mathbf{w}})(\mathbf{h},\mathbf{z})\|_{H_{2}}$ $\leq\left\|\mathbb{B}\begin{bmatrix}0&0&0&0&0\\ 0&\mathbb{q}_{1}&\mathbb{q}_{2}&0&0\\ 0&-\mathbb{q}_{1}&-\mathbb{q}_{2}&0&0\\ 0&\mathbb{q}_{3}&\mathbb{q}_{4}&\mathbb{q}_{1}&\mathbb{q}_{2}\\ 0&-\mathbb{q}_{3}&-\mathbb{q}_{4}&-\mathbb{q}_{1}&-\mathbb{q}_{2}\end{bmatrix}\begin{bmatrix}0&0&0&0&0\\ 0&\mathbb{h}_{1,1}&0&0&0\\ 0&\mathbb{h}_{1,2}&\mathbb{h}_{1,1}&0&0\\ 0&\mathbb{h}_{2,1}&0&\mathbb{h}_{1,1}&0\\ 0&\mathbb{h}_{2,2}&\mathbb{h}_{2,1}&\mathbb{h}_{1,2}&\mathbb{h}_{1,1}\end{bmatrix}\begin{bmatrix}0\\ z_{1,1}\\ z_{1,2}\\ z_{2,1}\\ z_{2,2}\end{bmatrix}\right\|_{H_{2}}$ (A.25)
$\leq\left\|M_{1}\begin{bmatrix}\mathbb{Q}_{1}&\mathbb{Q}_{2}&0&0\\ \mathbb{Q}_{4}&\mathbb{Q}_{5}&\mathbb{Q}_{1}&\mathbb{Q}_{2}\end{bmatrix}\right\|_{2,X_{2}}\left\|\begin{bmatrix}\mathbb{h}_{1,1}&0&0&0\\ \mathbb{h}_{1,2}&\mathbb{h}_{1,1}&0&0\\ \mathbb{h}_{2,1}&0&\mathbb{h}_{1,1}&0\\ \mathbb{h}_{2,2}&\mathbb{h}_{2,1}&\mathbb{h}_{1,2}&\mathbb{h}_{1,1}\end{bmatrix}\begin{bmatrix}z_{1,1}\\ z_{1,2}\\ z_{2,1}\\ z_{2,2}\end{bmatrix}\right\|_{2}$ (A.26)
$\leq 5\kappa r\sqrt{\left\|M_{1}M_{2}M_{1}^{*}\right\|_{X_{2}}}$ (A.27)
where the last step followed from the definition of $\mathbb{B}$ and properties of $\Gamma^{\dagger}$, similar algebraic steps to those performed in Lemma 4.3, and $M_{1},M_{2},M_{3}$ are defined in (135). Hence, by Parseval’s Identity, we need to analyze the following norms
$\|\overline{\text{{$\Pi$}}}^{\leq N}M_{1}M_{2}M_{1}^{*}\overline{\text{{$\Pi$}}}^{\leq N}\|_{X_{2}},\penalty 10000\ \|\overline{\text{{$\Pi$}}}^{\leq N}M_{1}M_{2}\text{{$\Pi$}}^{>N}\|_{2,X_{2}},\penalty 10000\ \|\text{{$\Pi$}}^{>N}M_{2}M_{1}^{*}\overline{\text{{$\Pi$}}}^{\leq N}\|_{X_{2},2},\penalty 10000\ \text{and}\penalty 10000\ \|\text{{$\Pi$}}^{>N}M_{2}\text{{$\Pi$}}^{>N}\|_{2}.$ (A.28)
The first term can be computed directly and is defined as $\mathcal{Z}_{2,3}$ in (139). Each of the remaining terms can be computed using similar steps to those used in Lemma 4.3. Indeed,
$\|\overline{\text{{$\Pi$}}}^{\leq N}M_{1}M_{2}\text{{$\Pi$}}^{>N}\|_{2,X_{2}}=\|\text{{$\Pi$}}^{>N}M_{2}M_{1}^{*}\overline{\text{{$\Pi$}}}^{\leq N}\|_{X_{2},2}=\sqrt{\|\overline{\text{{$\Pi$}}}^{\leq N}M_{1}M_{2}\text{{$\Pi$}}^{>N}M_{2}^{*}M_{1}^{*}\overline{\text{{$\Pi$}}}^{\leq N}\|_{X_{2}}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{2,4}$ (A.29)
$\|\text{{$\Pi$}}^{>N}M_{2}\text{{$\Pi$}}^{>N}\|_{2}\leq\varphi(\|Q_{1}^{2}+Q_{2}^{2}\|_{1},\|Q_{1}Q_{3}+Q_{2}Q_{4}\|_{1},\|Q_{3}Q_{1}+Q_{4}Q_{2}\|_{1},\|Q_{3}^{2}+Q_{4}^{2}+Q_{1}^{2}+Q_{2}^{2}\|_{1})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{Z}_{2,5}$ (A.30)
where we used the properties of the adjoint and Young’s inequality. Therefore, we obtain
$\left\|\mathbb{B}D^{2}\mathbb{G}(\overline{\mathbf{u}},\overline{\mathbf{w}})(\mathbf{h},\mathbf{z})\right\|_{H_{2}}\leq 5\kappa r\sqrt{\varphi(\mathcal{Z}_{2,3},\mathcal{Z}_{2,4},\mathcal{Z}_{2,4},\mathcal{Z}_{2,5})}.$ (A.31)
This concludes the computation of the $\mathcal{Z}_{2}$ bound. ∎
Appendix B Integral Computation using residues
The goal is to compute
$\int_{\mathbb{R}}\frac{(a_{1}\xi^{2}+a_{2})^{2}}{\ell_{\text{den}}(\xi)^{2}}d\xi$
To do this, we observe that we can write $\ell_{\text{den}}(\xi)=\nu|2\pi|^{4}(\xi^{2}-y_{1}^{2})(\xi^{2}-y_{2}^{2})$ where $y_{1}^{2},y_{2}^{2}$ are the roots of $x\mapsto 16\pi^{4}\nu x^{2}+4\pi^{2}(\nu\nu_{2}-\nu_{1})x-\nu_{2}(\nu_{1}+\nu_{3})$.
Now observe that the function $f(\xi)=\frac{(a_{1}\xi^{2}+a_{2})^{2}}{\nu|2\pi|^{4}(\xi^{2}-y_{1}^{2})^{2}(\xi^{2}-y_{2}^{2})^{2}}$ posses 4 poles of order 2 at $\pm y_{1},\pm y_{2}$. Without loss of generality, assume $\mathrm{Im}(y_{1}),\mathrm{Im}(y_{2})>0$ (if not, take $y_{1}$ or $y_{2}$, Assumption 1 also guarantees that $\mathrm{Im}(y_{1}),\mathrm{Im}(y_{1})\neq 0$). Using Residue Theory, we have that
$\int_{\mathbb{R}}f(\xi)d\xi=2\pi i\left(\mathrm{Res}(f,y_{1})+\mathrm{Res}(f,y_{2})\right)$
It remains to compute $\mathrm{Res}(f,y_{i})$, for $i=1,2$. Direct computations lead to
$\mathrm{Res}(f,y_{i})=\lim_{\xi\rightarrow y_{i}}\frac{d}{d\xi}\left((\xi-y_{i})^{2}f(\xi)\right)$ $=\lim_{\xi\rightarrow y_{i}}\frac{d}{d\xi}\left[\frac{(a_{1}\xi^{2}+a_{2})^{2}}{(2\pi)^{8}\nu^{2}(\xi+y_{i})^{2}(\xi^{2}-y_{j}^{2})^{2}}\right]$
$=\frac{1}{\nu^{2}|2\pi|^{8}}\lim_{\xi\rightarrow y_{i}}\frac{d}{d\xi}\left[\frac{(a_{1}\xi^{2}+a_{2})^{2}}{(\xi+y_{i})^{2}(\xi^{2}-y_{j}^{2})^{2}}\right],$
where $j=2$, if $i=1$ and $j=1$, if $i=2$. It just remains to take compute the derivative. In fact, we obtain that $\frac{d}{d\xi}(\xi+y_{i})^{2}(\xi^{2}-y_{j}^{2})^{2}=2(\xi+y_{i})(\xi^{2}-y_{j}^{2})\left(3\xi^{2}+2y_{i}\xi-y_{2}^{2}\right),\frac{d}{d\xi}(a_{1}\xi^{2}+a_{2})^{2}=4a_{1}\xi(a_{1}\xi^{2}+a_{2})$. Applying the quotient rule, we get:
$\mathrm{Res}(f,y_{i})$ $=\frac{1}{\nu^{2}|2\pi|^{8}}\lim_{\xi\rightarrow y_{i}}\frac{4a_{1}\xi(a_{1}\xi^{2}+a_{2})(\xi+y_{1})^{2}(\xi^{2}-y_{j}^{2})^{2}-(a_{1}\xi^{2}+a_{2})^{2}\cdot 2(\xi+y_{i})(\xi^{2}-y_{j}^{2})(3\xi^{2}+2y_{i}\xi-y_{j}^{2})}{(\xi+y_{i})^{4}(\xi^{2}-y_{j}^{2})^{2}}$
$=\frac{1}{\nu^{2}|2\pi|^{8}}\lim_{\xi\rightarrow y_{i}}2(a_{1}\xi^{2}+a_{2})\frac{2a_{1}\xi(\xi+y_{i})(\xi^{2}-y_{2}^{2})^{2}-(a_{1}\xi^{2}+a_{2})(3\xi^{2}-2y_{i}\xi-y_{2}^{2})}{(\xi+y_{i})^{3}(\xi^{2}-y_{j}^{2})^{3}}$
$=\frac{2(a_{1}y_{i}^{2}+a_{2})(4a_{1}y_{i}^{2}(y_{i}^{2}-y_{j}^{2})-(a_{1}y_{1}^{2}+a_{2})(5y_{1}^{2}-y_{2}^{2}))}{\nu^{2}|2\pi|^{8}\cdot 8y_{i}^{3}(y_{i}^{2}-y_{j}^{2})^{2}}.$
These formulas are then implemented and evaluated using interval arithmetic in julia_blanco_fassler .